### v1

# SAMBO-RL: Shifts-aware Model-based Offline Reinforcement Learning

Abstract
Model-based Offline Reinforcement Learning trains policies based on offline datasets and model dynamics, without direct real-world environment interactions. However, this method is inherently challenged by distribution shift. Previous approaches have primarily focused on tackling this issue directly leveraging off-policy mechanisms and heuristic uncertainty in model dynamics, but they resulted in inconsistent objectives and lacked a unified theoretical foundation. This paper offers a comprehensive analysis that disentangles the problem into two key components: model bias and policy shift. We provide both theoretical insights and empirical evidence to demonstrate how these factors lead to inaccuracies in value function estimation and impose implicit restrictions on policy learning. To address these challenges, we derive adjustment terms for model bias and policy shift within a unified probabilistic inference framework. These adjustments are seamlessly integrated into the vanilla reward function to create a novel Shifts-aware Reward (SAR), aiming at refining value learning and facilitating policy training. Furthermore, we introduce Shifts-aware Model-based Offline Reinforcement Learning (SAMBO-RL), a practical framework that efficiently trains classifiers to approximate the SAR for policy optimization. Empirically, we show that SAR effectively mitigates distribution shift, and SAMBO-RL demonstrates superior performance across various benchmarks, underscoring its practical effectiveness and validating our theoretical analysis.
footnotetext: footnotetext: Preprint. Under Review.
Introduction
Offline reinforcement learning (RL) (Lange, Gabel, and Riedmiller 2012; Levine et al. 2020) learns policies from a pre-collected dataset generated by a behavior policy, avoiding additional online interaction with the environment. This approach showcases remarkable potential in data-driven decision scenarios (Emerson, Guy, and McConville 2023; Sinha, Mandlekar, and Garg 2022) where the exploration is costly or hazardous. The model-based framework (Luo et al. 2018) is particularly effective for offline RL (Yu et al. 2020, 2021; Sun et al. 2023; Li et al. 2024). It involves learning models of the environment from the given dataset and utilizing these models to generate data for policy training. Nevertheless, directly learning a policy using offline and model-generated data introduces the distribution shift challenge. This shift causes the training objective to deviate from the true objective of RL, leading to poor test performance.
Prior methods in model-based offline RL focus on directly leveraging off-policy methods in online RL and heuristic model uncertainty for conservative learning to mitigate the impact of distribution shift. For instance, MOPO (Yu et al. 2020) uses the predicted variance of the learned model as an uncertainty penalty on rewards. MOBILE (Sun et al. 2023) penalizes value learning of model data according to the uncertainty of off-policy methods in online RL and the Bellman Q-function estimated by ensemble models. RAMBO (Rigter, Lacerda, and Hawes 2022) considers the trajectory distribution in the worst case and adversarially trains the policy and transition model while ensuring accurate predictions. However, these approaches have resulted in a biased training objective due to the model uncertainty and remain the distribution shift problem to be solved.
In this paper, we demystify the distribution shift problem in model-based RL and study its fundamental properties. We first show that this problem arises from both model bias and policy shift, and further analyze their impacts on policy optimization. Unlike the heuristic descriptions of model uncertainty, model bias fundamentally characterizes the inconsistency between the learned model and actual environment dynamics in predicting subsequent states. This bias leads to inaccurate estimations of trajectory returns, resulting in distorted objectives. If the learned model overestimates or underestimates the probability of generating a trajectory, it consequently causes the overestimation or underestimation of trajectory returns. Additionally, policy shift depicts the changes in state-action visitation distribution caused by the deviation between the learned policy $\pi$ and behavior policy $\pi_{b}$. This shift implicitly forces $\pi$ towards $\pi_{b}$, hindering the learning process and slowing down policy convergence. In contrast, previous studies heuristically mitigate these two issues from off-policy algorithms and model uncertainty.
Figure 1: The negative impacts of model bias and policy shift. (a) In the 1D grid world, the agent can take two actions: move left (L) or move right (R) to the adjacent grid, with movement obstructed at the boundaries. The initial state is uniformly distributed, and the agent receives rewards upon reaching specific states. (b) The left panel illustrates value estimates during training and the final policy convergence with and without model bias. Even slight model bias can significantly distort value function estimates, leading the learned policies to deviate from the true optimal policy in the actual environment. (c) The right panel showcases expected returns and the KL divergence between the learned policy $\pi$ and the behavior policy $\pi_{b}$ during training with and without policy shift. We find that policy shift implicitly restricts $\pi$ to remain close to $\pi_{b}$, thereby slowing policy convergence.
To address these challenges, we theoretically derive adjustments for model bias and policy shift by employing a unified probabilistic inference framework, and introduce a shifts-aware reward, which integrates the vanilla reward with adjustments for model bias and policy shift. The model bias adjustment penalizes the reward to mitigate issues of overestimation or underestimation, thereby refining value learning and preventing its convergence to local optima. Meanwhile, the policy shift modification revises the reward to motivate $\pi$ to escape the limitation of $\pi_{b}$, accelerating policy convergence. Building upon our theoretical framework, we develop Shifts-aware Model-based Offline Reinforcement Learning (SAMBO-RL). This approach incorporates model bias adjustment for model-generated data and policy shift modification for offline data. Furthermore, it learns a transition classifier and an action classifier to approximate the shifts-aware reward and then performs policy optimization based on this adjusted reward. Our contributions include:
  • We demystify the distribution shift problem in model-based offline RL, i.e., model bias distorts the estimation of value function, while policy shift imposes an implicit policy limitation to hinder policy convergence.
  • We upgrade the vanilla reward into a shifts-aware reward with model bias adjustment and policy shift modification. This shifts-aware reward is used for policy learning to mitigate distribution shift in model-based offline RL.
  • We devise SAMBO-RL, which trains transition and action classifiers to approximate the shifts-aware reward, effectively addressing both model bias and policy shift.
  • We show that shifts-aware reward effectively mitigates distribution shift, and SAMBO-RL demonstrates superior performance across various benchmarks.
Related Work
Model-free Offline RL encompasses two primary approaches: policy constraints and value regularization, both incorporating conservatism (Jin, Yang, and Wang 2021). Policy constraint methods aim to keep the learned policy close to the behavioral policy, mitigating distribution shift. For instance, BEAR (Kumar et al. 2019) and BRAC (Wu, Tucker, and Nachum 2019) constrain the learned policy by minimizing different divergences. BCQ (Fujimoto, Meger, and Precup 2019) perturbs actions during learning to ensure they remain within the action space of the offline dataset by learning a generative model of the behavior policy. On the other hand, value regularization methods add conservative regularization terms to the value optimization objective, avoiding out-of-distribution (OOD) actions. CQL (Kumar et al. 2020) penalizes the Q-values of OOD samples by incorporating regularization into the value loss function. TD3+BC (Fujimoto and Gu 2021) introduces a behavior cloning regularization into the TD3 (Fujimoto, Hoof, and Meger 2018) objective. Similarly, EDAC (An et al. 2021) and PBRL (Bai et al. 2021) penalize the Q-function based on uncertainty derived from an ensemble of Q-networks. MCQ (Lyu et al. 2022) actively trains OOD actions by constructing pseudo-target values to alleviate over-pessimism.
Model-based Offline RL approximates the environment using learned models and performs conservative policy optimization (Lu et al. 2021), achieving high data efficiency (Li et al. 2024). MOPO (Yu et al. 2020) incorporates conservatism by penalizing rewards based on the uncertainty of model predictions. MOReL (Kidambi et al. 2020) introduces a pessimistic MDP to penalize rewards of state-action pairs in unexplored regions. COMBO (Yu et al. 2021) extends CQL to the model-based setting, regularizing the value function on OOD samples. RAMBO (Rigter, Lacerda, and Hawes 2022) adopts an adversarial approach, training the policy and model within a robust framework while ensuring accurate transition predictions of the model. CBOP (Jeong et al. 2022) adaptively weights multi-step returns in a model-based value expansion framework (Feinberg et al. 2018). Lastly, MOBILE (Sun et al. 2023) introduces penalties during value learning by leveraging the uncertainty of the Bellman Q-function estimates derived from ensemble models.
Preliminaries
MDP.
We focus on a Markov decision process (MDP) specified by the tuple $M=(\mathcal{S},\mathcal{A},p,r,\mu_{0},\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $p(s^{\prime}|s,a)$ is the environment transition dynamics, positive $r(s,a)>0$ is the reward function, $\mu_{0}(s)$ is the initial state distribution, and $\gamma\in(0,1)$ is the discount factor. Given an MDP, we define the state value function $V^{\pi}(s)=\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right]$ and the $Q$ function $Q^{\pi}(s,a)=\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right]$.
MBPO.
Similar to previous works (Yu et al. 2020, 2021; Rigter, Lacerda, and Hawes 2022; Sun et al. 2023), we focus on model-based policy optimization (MBPO) (Janner et al. 2019), which employs an actor-critic RL algorithm. MBPO first performs $k$-step rollouts using the model $m(s^{\prime}|s,a)$ starting from state $s\in\mathcal{D}_{env}$ and adds the generated data to a replay buffer $\mathcal{D}_{m}$. Then it optimizes the policy using mini-batches of data sampled from $\mathcal{D}_{env}\cup\mathcal{D}_{m}$, where each datapoint is sampled from the real dataset $\mathcal{D}_{env}$ with probability $f$ and the model dataset $\mathcal{D}_{m}$ with probability $1-f$.
Trajectory Formulation.
To demystify the distribution shift, We consider the trajectory formulation of optimization objectives. Let $p^{\pi}(\cdot)$ represent the probability distribution over trajectories =$\mathcal{H}=\{\tau=(s_{0},a_{0},r_{0},s_{1},...)\}$ generated by executing policy $\pi$ under dynamics $p$. This is given by:
$p^{\pi}(\tau)=\mu_{0}(s_{0})\prod\limits_{t=0}^{\infty}p(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t}).$ (1)
The training data $\mathcal{D}_{train}$ can be interpreted as being generated by a data collection policy $\pi_{c}$ through rollouts in dynamics $q$. Consequently, the probability of observing trajectory $\tau$ in the training data can be formulated as $q^{\pi_{c}}(\tau)$. The reinforcement learning objective and the practical training objective are then formalized as the following equations:
$\mathcal{J}_{M}(\pi)=\mathbb{E}_{\tau\sim p^{\pi}}\left[R(\tau)\right]\text{ and }\mathcal{J}(\pi)=\mathbb{E}_{\tau\sim q^{\pi_{c}}}[R(\tau)],$ (2)
where $R(\tau)=\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})$ represents the discounted return of the whole trajectory $\tau=(s_{0},a_{0},r_{0},s_{1},...)$.
Shifts-aware Policy Optimization
We first formalize the distribution shift problem as shift weighting and further delve into its negative effects on policy optimization from theory and experiments. To address these challenges, we employ probabilistic inference to derive a surrogate objective and formulate a shifts-aware reward. Finally, we implement this theoretical framework within model-based offline RL and introduce SAMBO-RL, a practical algorithm designed to mitigate distribution shift.
Demystifying Distribution Shift
We examine the distribution shift problem from an optimization perspective and identify its origins as model bias and policy shift. Furthermore, we explore how these factors affect accuracy in value estimation and impose implicit policy restrictions, resulting in a biased objective and slow convergence. The supplementary analysis is in Appendix A1.
Shift Weighting.
The training objective $\mathcal{J}(\pi)$ in (2) can be expressed as the following trajectory formulation:
$\mathcal{J}(\pi)=\mathbb{E}_{\tau\sim p^{\pi}(\tau)}\left[\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}R(\tau)\right].$ (3)
Compared to (2), it indicates that for a given trajectory $\tau$, the true objective under practical sampling distributions differs in shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$ from the training objective, which formalizes the distribution shift in trajectory formulation.
Substituting (1) into the shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$, we can attribute this weighting to two primary components: model bias and policy shift, reformulated as the following equation:
$\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}=\underbrace{\prod_{t}\frac{q(s_{t+1}|s_{t},a_{t})}{p(s_{t+1}|s_{t},a_{t})}}_{\text{model bias}}\underbrace{\prod_{t}\frac{\pi_{c}(a_{t}|s_{t})}{\pi(a_{t}|s_{t})}}_{\text{policy shift}}.$ (4)
Model bias illustrates the discrepancy between the dynamics used for training and those of the actual environment, while policy shift represents the divergence between the policies used for sampling training data and those being learned.
Negative Impact of Shift Weighting.
We analyze the impact of shift weighting from model bias and policy shift.
(i) To illustrate the effect of model bias, we examine the scenario where the training policy $\pi$ operates within model dynamics $q$, which diverges from actual environment dynamics $p$. In this context, the training objective $\mathcal{J}(\pi)$ is:
$\mathcal{J}(\pi)$ $=\mathbb{E}_{\tau\sim p^{\pi}}\left[R(\tau)\prod\limits_{t=0}^{\infty}\frac{q(s_{t+1}|s_{t},a_{t})}{p(s_{t+1}|s_{t},a_{t})}\right].$
Compared to the true objective (3), the compounded prediction bias of the model significantly affects value estimation and leads to a distorted objective. Specifically, when the model $q$ overestimates or underestimates the likelihood of generating trajectory $\tau$, it consequently overestimates or underestimates the expected return of this trajectory in the real environment. This essentially introduces a bias for the optimization objective, resulting in wrong training directions.
(ii) To delve deeper into the policy optimization process under policy shift, we consider the setting where a sampling policy $\pi_{c}$ is executed within the true environment transition dynamics $p$. In this scenario, the training objective $\mathcal{J}(\pi)$ is:
$\mathcal{J}(\pi)$ $=\mathbb{E}_{(s,a)\sim d^{p,\pi}}\left[r(s,a)\frac{d^{p,\pi_{c}}(s,a)}{d^{p,\pi}(s,a)}\right].$
Here, $d^{p,\pi}(s,a)$ is the state-action visitation distribution under dynamics $p$ and policy $\pi$, i.e. $d^{p,\pi}(s,a)=(1-\gamma)\mathbb{E}_{s_{0}\sim\mu_{0},s_{t}\sim p,a_{t}\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{I}(s_{t}=s,a_{t}=a)\right]$, where $\mathbb{I}$ represents the indicator function. Note that $\pi_{c}$ is related to $\pi$. Compared to the true objective, the policy shift implicitly forces policy $\pi$ to approach $\pi_{c}$, hindering the learning process and slowing down policy convergence. Specifically, for a state-action pair $(s,a)$ associated with a high reward, ideal policy updates should increase $\pi(a|s)$. However, if the currently learned policy $\pi$ adequately approximates the optimal policy and satisfies $d^{p,\pi}(s,a)>d^{p,\pi_{c}}(s,a)$, the reward for $(s,a)$ can be underestimated. This situation could even prompt policy updates to decrease $\pi(a|s)$, contrary to the optimal direction. Similarly, for the state-action pair $(s,a)$ with a low reward, if the currently $\pi$ is reasonably good and leads to $d^{p,\pi}(s,a)<d^{p,\pi_{c}}(s,a)$, the reward can be overestimated, potentially resulting in policy updates to increase $\pi(a|s)$ that do not align with optimal policy directions.
Toy Examples.
To empirically validate our analysis of model bias and policy shift, we use a 1D grid world example, as depicted in Fig. 1 (a). In Fig. 1 (b), we compare the training process with and without model bias and find the value function overestimation and underestimation under model bias. These results reveal that even slight deviations in model dynamics can significantly distort value function estimates, leading to a biased training objective. Consequently, the policies learned by the agents deviate from the optimal policy. To assess the impact of the policy shift, we examine a model-free offline setting where data is pre-collected using a behavior policy $\pi_{b}$. Fig. 1 (c) showcases that the KL divergence between the learned policy $\pi$ and $\pi_{b}$ under policy shift is substantially reduced compared to the scenario without policy shift, sometimes even decreasing undesirably. This indicates that policy shift does compel $\pi$ to align closely with $\pi_{b}$, thereby impeding the training process and substantially slowing down the policy convergence.
These analyses and empirical evidence highlight the critical role of addressing model bias and policy shift in achieving accurate value learning and efficient policy training.
Shifts-aware Reward via Probabilistic Inference
This subsection introduces a novel shifts-aware reward through a unified probabilistic inference framework, serving as a lower-bound surrogate for the true objective. We further examine how this modified reward contributes to refining the value function and motivating $\pi$ to escape the limitations of behavior policy, thus mitigating model bias and policy shift.
Shifts-aware Reward.
We introduce the shifts-aware reward $\tilde{r}(s_{t},a_{t},s_{t+1})$ formulated within trajectories. It comprises the vanilla reward, adjustments for model bias, and modifications for policy shift, expressed as the following:
$\begin{split}&\tilde{r}(s_{t},a_{t},s_{t+1})=\log r(s_{t},a_{t})\\ &+\frac{1}{(1-\gamma)\gamma^{t}}\left[\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right].\end{split}$ (5)
This modified reward forms the basis of our surrogate objective function during training, which is formulated as:
$\mathcal{L}(\pi)=\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\sum_{t=0}^{\infty}\gamma^{t}\tilde{r}(s_{t},a_{t},s_{t+1})\right].$ (6)
Additionally, we establish that $\mathcal{L}(\pi)$ serves as a lower bound for the true objective, ensuring its rationality as a surrogate.
Theorem 1.
Let $\mathcal{L}(\pi)$ be the training objective defined in (6) and $\tilde{r}$ denote the shifts-aware reward (5). Then,
$\log\left[\mathcal{J}_{M}(\pi)\right]\geq(1-\gamma)\mathcal{L}(\pi),\quad\forall\ \pi.$
Optimization objective based on this shifts-aware reward provides a methodological foundation for integrating shift adjustments seamlessly into reinforcement learning algorithms to uniformly address model bias and policy shift.
Modal Bias Adjustment.
We now examine how this adjustment refines the value estimation and addresses the challenge of model bias. By rolling out the learned policy $\pi$ within the model dynamics $m$, we obtain a model-generated dataset $\mathcal{D}_{m}$. Furthermore, the modified training objective of the model dataset $\mathcal{D}_{m}$ can be computed as the following:
$\tilde{J}(\pi)=\mathbb{E}_{s\sim\mathcal{D}_{m},a\sim\pi(\cdot|s),s^{\prime}\sim\mathcal{D}_{m}}\left[\tilde{r}(s,a,s^{\prime})\right]$
$=\mathbb{E}_{(s,a)\sim\mathcal{D}_{m}}\left[\log r(s,a)-\alpha D_{\operatorname{KL}}(m(\cdot|s,a)\|p(\cdot|s,a))\right].$
This modified reward consists of the logarithmic vanilla reward penalized by the KL divergence between the model and environment dynamics. This adjustment reduces value function misestimations by accounting for the accuracy of the learned model, encouraging the development of both a high-return policy and an accurate model approximation.
Policy Shift Modification.
To delve into the positive effect of policy shift modification, we consider a dataset $\mathcal{D}_{b}$ generated by executing the behavior policy $\pi_{b}$ within the environment $p$. Moreover, the modified training objective of the dataset $\mathcal{D}_{b}$ is derived as the following formulation:
$\tilde{J}(\pi)=\mathbb{E}_{s\sim\mathcal{D}_{b},a\sim\pi(\cdot|s),s^{\prime}\sim\mathcal{D}_{b}}\left[\tilde{r}(s,a,s^{\prime})\right]$
$=\mathbb{E}_{s\sim\mathcal{D}_{b}}\left[\mathbb{E}_{a\sim\pi(\cdot|s)}\left[\log r(s,a)\right]+\beta D_{\operatorname{KL}}(\pi(\cdot|s)||\pi_{b}(\cdot|s))\right],$
which introduces a bonus to the vanilla reward based on the KL divergence between $\pi$ and $\pi_{b}$. This modification guides $\pi$ to break out of the confines of the behavior policy $\pi_{b}$, promoting better exploration and more effective learning.
Shifts-aware Model-based Offline RL
We apply the shifts-aware reward in model-based offline RL and develop the practical framework, SAMBO-RL.
To adapt the shifts-aware reward for practical implementation, where training data is typically stored in transition format rather than trajectories, we introduce two weighting factors, $\alpha$ and $\beta$. These factors replace the respective coefficients in the reward function, shaping it as the following:
$\tilde{r}(s,a,s^{\prime})=\log r(s,a)+\alpha\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}+\beta\log\frac{\pi(a|s)}{\pi_{c}(a|s)}.$
These weighting factors ensure the reward and adjustments are properly balanced in transition-based learning scenarios, which is empirically validated in the following experiments.
Furthermore, since the environment dynamics $p$ and behavior policy $\pi_{b}$ are typically unknown and cannot be exactly computed, we derive two classifiers to estimate the shifts-aware reward (Eysenbach et al. 2020, 2022). Details are shown in Appendix A3. Specifically, these classifiers estimate model bias adjustment $\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}$ and policy shift modification $\log\frac{\pi(a|s)}{\pi_{c}(a|s)}$ respectively. The transition classifier $C_{\phi}(s,a,s^{\prime})$ estimates the probability that the transition $(s,a,s^{\prime})$ comes from the environment, while the action classifier $C_{\psi}(s,a)$ estimate the probability that a given state-action pair $(s,a)$ is generated by the current policy $\pi$.
We train the classifiers by minimizing the loss functions:
$\mathcal{L}(\varphi)=-\mathbb{E}_{\mathcal{D}_{1}}[\log C_{\varphi}]-\mathbb{E}_{\mathcal{D}_{2}}[\log(1-C_{\varphi})].$ (7)
For the transition classifier $C_{\varphi}=C_{\phi}(s,a,s^{\prime})$, $\mathcal{D}_{1}=\mathcal{D}_{env}$ is the offline dataset, and $\mathcal{D}_{2}=\mathcal{D}_{m}$ is the model-generated dataset. For the action classifier $C_{\varphi}=C_{\psi}(s,a)$, $\mathcal{D}_{1}=\mathcal{D}_{\pi}$ is the dataset collected by executing the current policy $\pi$, and $\mathcal{D}_{2}=\mathcal{D}_{env}$ is the offline dataset. Based on these classifiers, we approximate the shifts-aware reward as the following:
$\tilde{r}(s,a,s^{\prime})\approx\log r(s,a)$ (8)
$+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})}+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$
The complete procedure of SAMBO-RL is outlined in Algorithm 1, with further details provided in Appendix B.
Algorithm 1 SAMBO-RL
Input: Offline dataset $\mathcal{D}_{env}$, learned dyanmics models $\{m_{\theta}^{i}\}_{i=1}^{N}$, initialized policy $\pi$, shifts-aware reward hyperparameters $\alpha$ and $\beta$.
1:  Train the probabilistic dynamics model $m_{\theta}(s^{\prime},r|s,a)=\mathcal{N}(\mu_{\theta}(s,a),\Sigma_{\theta}(s,a))$ on $\mathcal{D}_{env}$.
2:  Initialize the model dataset buffer $\mathcal{D}_{m}\leftarrow\varnothing$.
3:  for $i=1,2,...,N_{iter}$ do
4:     Initialize the policy dataset buffer $\mathcal{D}_{\pi}\leftarrow\varnothing$.
5:     Generate synthetic rollouts by model $m_{\theta}$. Add transition data in these rollouts to $\mathcal{D}_{m}$ and $\mathcal{D}_{\pi}$.
6:     Update classifiers $C_{\phi}$ and $C_{\psi}$ according to Eq. (7).
7:     Compute shifts-aware reward according to Eq. (8).
8:     Run SAC (Haarnoja et al. 2018) with shifts-aware reward to update policy $\pi$.
9:  end for
Experiments
In our experiments, we focus on three objectives: (1) validating the effectiveness of shifts-aware reward, (2) comparing the performance of SAMBO with previous methods, and (3) analyzing the contribution of each component and hyperparameter within SAMBO to its overall performance.
We evaluate our algorithm using standard offline RL benchmarks in Mujoco environments and conduct an ablation study to assess its performance. Our implementation is based on the OfflineRL-Kit library***https://github.com/yihaosun1124/OfflineRL-Kit, a comprehensive and high-performance library for implementing offline reinforcement learning algorithms. The basic parameters of our algorithm are consistent with the settings in this library.
Effectiveness of Shifts-aware Reward
(a) Model Bias Adjustment
(b) Policy Shift Modification
Figure 2: Assessment of the shifts-aware reward. (a) Model bias adjustment mitigates inaccuracies in value estimation, achieving a nearly optimal return. (b) Policy shift modification addresses implicit policy restrictions, leading to faster convergence.
We conduct an evaluation of the effectiveness of the Shifts-aware Reward in addressing model bias and policy shift issues within the 1D grid environment illustrated in Fig. 1 (a).
Model Bias Adjustment.
Model bias exacerbates inaccuracies in value estimates, ultimately resulting in a distorted optimization objective. To evaluate the effectiveness of model bias adjustment, we examine its amendment to the training process. As illustrated in Fig. 2 (a), we execute the policy gradient algorithm under four settings: using the vanilla reward with the overestimating model (OM), the shifts-aware reward with the overestimating model (SAR with OM), the vanilla reward with the underestimating model (UM), and the shifts-aware reward with the underestimating model (SAR with UM). We then measure the expected return of the learned policy after each update during training in the actual environment. The results show that while the OM initially achieves optimal returns, it rapidly diverges from the optimal policy. In contrast, SAR with OM effectively counters the performance degradation caused by model overestimation, thereby sustaining optimal returns. Similarly, SAR with UM mitigates the policy suboptimality due to model underestimation, achieving near highest return.
Policy Shift Modification.
Policy shift implicitly restricts policy updates. To assess the effectiveness of the policy shift modification, we apply the policy gradient algorithm to the offline objective $\mathcal{J}(\pi)=\mathbb{E}_{s\sim d^{p,\pi_{b}}}\left[V^{\pi}(s)\right]$ in the environment depicted in Fig. 1 (a). We test four settings: using either the vanilla reward or the shifts-aware reward, combined with a behavior policy that is either uniformly distributed or the opposite of the optimal policy. We measure the KL divergence between the learned policy and the behavior policy after each update. As illustrated in Fig. 2 (b), the algorithm using the vanilla reward is limited by the behavior policy, serving a lower KL divergence and slower convergence. In contrast, the algorithm employing the shifts-aware reward effectively motivates policy to be aggressive and optimistic and overcomes this restriction, displaying a higher and more stable divergence and achieving faster policy convergence.
Results on the D4RL Benchmark
Datasets.
In line with prior research, SAMBO is evaluated using the D4RL benchmark (Fu et al. 2020) on the MuJoCo simulator (Todorov, Erez, and Tassa 2012). The evaluation encompasses 12 datasets spanning three environments (HalfCheetah, Hopper, Walker2d) and four dataset types (random, medium, medium-replay, medium-expert). The “v2” versions of datasets are employed for assessment.
Baselines.
We compare SAMBO against multiple baselines, including both model-free and model-based offline RL algorithms. For model-free approaches, BC (Bain and Sammut 1995; Ross, Gordon, and Bagnell 2011) learns behavior policy, CQL (Kumar et al. 2020) penalizes the Q-function according to policy change, TD3+BC (Fujimoto and Gu 2021) adds a BC term to regularize the policy, and EDAC (An et al. 2021) quantifies the uncertainty of the Q-value using an ensemble of Q-function networks. For model-based approaches, MOPO (Yu et al. 2020) penalizes rewards via the predicted variance of models. COMBO (Yu et al. 2021) implements CQL within a model-based framework. TT (Janner, Li, and Levine 2021) uses a transformer to model offline trajectories and employs beam search for planning. RAMBO (Rigter, Lacerda, and Hawes 2022) adversarially trains the policy and the dynamics model within a robust framework. MOBILE (Sun et al. 2023) penalizes the value learning of model data based on the uncertainty of the Bellman Q-function, derived from an ensemble of models.
Comparison Results.
Results are presented in Table 1. Each number represents the normalized score, calculated as 100×100$\times$100 ×(score -- random policy score)/ (expert policy score -- random policy score) (Fu et al. 2020). SAMBO achieves the highest scores on two random datasets and demonstrates strong performance across both medium and medium-replay datasets. Additionally, SAMBO shows competitive results on medium-expert datasets. This performance aligns with our theoretical analysis. By addressing both model bias and policy shift, SAMBO effectively balances the exploitation of offline data with the need to overcome the limitations of the behavioral policy. This approach enables the algorithm to be more aggressive and optimistic in its search for the optimal policy. In scenarios involving low-quality datasets, the aggressive exploration of SAMBO is more likely to yield benefits, potentially leading to the discovery of a superior policy. Conversely, a conservative approach may impede the learning of optimal policy in such contexts. For high-quality datasets, the advantages of aggressive exploration tend to diminish. Fortunately, in these scenarios, methods that emphasize estimating Q-value uncertainty perform effectively. This suggests that our approach and conservative estimation methods can perform well on complementary datasets.
Task Name BC CQL TD3+BC EDAC MOPO MOPO* COMBO TT RAMBO MOBILE SAMBO (Ours)
halfcheetah-random 2.2 31.3 11.0 28.4 35.4 38.5 38.8 6.1 39.5 39.3 39.7$\pm$2.0
hopper-random 3.7 5.3 8.5 25.3 11.7 31.7 17.9 6.9 25.4 31.9 32.4$\pm$0.5
walker2d-random 1.3 5.4 1.6 16.6 13.6 7.4 7.0 5.9 0.0 17.9 7.0$\pm$8.2
halfcheetah-medium 43.2 46.9 48.3 65.9 42.3 72.4 54.2 46.9 77.9 74.6 72.5$\pm$3.8
hopper-medium 54.1 61.9 59.3 101.6 28.0 62.8 97.2 67.1 87.0 106.6 99.7$\pm$0.3
walker2d-medium 70.9 79.5 83.7 92.5 17.8 84.1 81.9 81.3 84.9 87.7 78.3$\pm$2.6
halfcheetah-medium-replay 37.6 45.3 44.6 61.3 53.1 72.1 55.1 44.1 68.7 71.7 70.6$\pm$1.6
hopper-medium-replay 16.6 86.3 60.9 110.0 67.5 92.7 89.5 99.4 99.5 103.9 101.5$\pm$1.5
walker2d-medium-replay 20.3 76.8 81.8 87.1 39.0 85.9 56.0 82.6 89.2 89.9 82.8$\pm$0.2
halfcheetah-medium-expert 44.0 95.0 90.0 106.3 63.3 83.6 90.0 95.0 95.4 108.2 94.8$\pm$2.8
hopper-medium-expert 53.9 96.9 98.0 110.7 23.7 74.6 111.0 110.0 88.2 112.6 106.5$\pm$5.6
walker2d-medium-expert 90.1 109.1 110.1 114.7 44.6 108.2 103.3 101.9 56.7 115.2 80.2$\pm$6.8
Table 1: Results on the D4RL Gym benchmark. MOPO’s results are obtained on the “v0” datasets; MOPO*’s results are obtained from experiments in the OfflineRL-Kit library on the “v2” datasets. The numbers reported for SAMBO are the normalized scores averaged over the final iteration of training across 4 seeds, with $\pm$ standard deviation. The top three scores are highlighted in bold. We use an asterisk “*” for the best score, an underscore “_” for the second, and bold text for the third.
Ablation Studies
We conducted ablation studies on our method to understand why SAMBO demonstrates competitive performance. The following experimental results are presented in Figure 3.
(a) Ablation study for SAMBO
(b) Hyperparameter tuning for $\alpha$
(c) Hyperparameter tuning for $\beta$
Figure 3: Illustration of hyperparameter experiments in the HalfCheetah task with a random dataset. (a) The effectiveness of model bias adjustment and policy shift modification in the shifts-aware reward. (b) A large model bias adjustment coefficient $\alpha$ increases conservativeness. (c) A large policy shift modification coefficient $\beta$ increases the aggressiveness of exploration.
Ablation.
Our shifts-aware reward integrates both the model bias adjustment and the policy shift modification. To elucidate the importance of these components, we perform ablation studies on the Halfcheetah task. We assess the performance under four configurations: SAMBO, SAMBO without model bias adjustment (wo/mb), SAMBO without policy shift modification (wo/ps), and SAMBO without any modifications (logr), while keeping all other parameters consistent. The ablation results are presented in Figure. 3 (a).
We find that, compared to the experiment denoted as logr, the performance of wo/mb deteriorates. This degradation arises because wo/mb employs policy shift modification to break the implicit limitations of the behavior policy and encourage more aggressive exploration. However, it does not address model bias, resulting in significant inaccuracies in value estimation, ultimately leading to suboptimal performance. In contrast, SAMBO incorporates both model bias adjustment and policy shift modification, facilitating aggressive exploration while addressing value estimation challenges, thus alleviating suboptimal performance. This underscores the effectiveness of model bias adjustment.
The ablation experiment wo/ps incorporates model bias adjustment but does not address policy shift. As a result, while wo/ps mitigates value misestimation caused by model bias, the lack of policy shift modification causes the algorithm to remain implicitly constrained by the behavioral policy. Therefore, wo/ps demonstrates reduced effectiveness in discovering a more optimal policy. In contrast, SAMBO integrates policy shift modification during conservative updates, which enables more aggressive exploration and results in superior performance compared to wo/ps. This highlights the effectiveness of incorporating policy shift modification.
Hyperparameter Tuning.
SAMBO is primarily regulated by two key hyperparameters: the model bias adjustment coefficient $\alpha$ and the policy shift modification coefficient $\beta$.
The model bias adjustment coefficient $\alpha$ regulates the extent of adjustment applied to model-generated data. A larger $\alpha$ imposes a significant penalty on the model, leading to excessively conservative utilization of the model. Consequently, even in regions where the model performs well, exploration may be restricted due to this excessive penalty, potentially resulting in suboptimal performance. Conversely, a smaller $\alpha$ can cause the algorithm to be overly optimistic about exploring risky regions, thereby impairing performance. To validate this, we set $\beta=0.01$ and varied $\alpha$ in the Halfcheetah task. The results, shown in Fig. 3 (b), indicate that both very low ($\alpha=0$) and high ($\alpha=0.02$ or $0.05$) values lead to decreased performance. Specifically, as $\alpha$ increases from 0 to 0.05, the algorithm initially improves but later declines as it becomes increasingly conservative, which impedes exploration. This underscores that both excessively low and high values of $\alpha$ are detrimental to policy learning.
The policy shift modification coefficient $\beta$ governs the extent of modification applied to the offline data, thereby influencing the degree of deviation from the behavior policy. A larger $\beta$ value promotes increased exploratory behavior by encouraging the algorithm to diverge from the behavior policy. However, this also introduces a heightened risk of taking unsafe actions. As illustrated in Fig. 3 (c), the algorithm exhibits some performance fluctuations as $\beta$ increases from 0 to 0.1, though it generally maintains robustness. A low value of $\beta$ may lead to suboptimal solutions, whereas a high value of $\beta$ can induce excessive aggressiveness of the algorithm, resulting in performance instability. Therefore, $\beta$ also needs to be set appropriately to balance exploration and stability.
Conclusion and Discussion
In this paper, we analyze the distribution shift problem, attributing its root causes to model bias and policy shift. Model bias results in distorted objectives, while policy shift obstructs the convergence of policies. We introduce a shifts-aware policy optimization framework to uniformly address these challenges and apply it within the model-based offline setting. Notably, our framework is versatile and can be integrated into general reinforcement learning algorithms. Our main limitation is classifier accuracy; imprecise classifiers struggle to estimate the shifts-aware reward accurately, leading to unstable training. In future work, We aim to explore alternative methods that directly address distribution shift using datasets, eliminating the need for classifier estimation. Additionally, we will extend our shifts-aware reward framework to tackle distribution shift challenges in online settings, with a particular focus on policy shift problems. Our novel analysis and results related to policy shift in this context open new avenues for better exploiting offline datasets and overcoming the restrictions imposed by behavior policies.
Appendix A A. Theoretical Analysis
A1. The Supplement Analysis of Distribution Shift
Dual Formulation of Model Bias.
The reinforcement learning optimization objective can be equivalently reformulated in its dual form, expressed as $\mathcal{J}_{M}(\pi)=\mathbb{E}_{(s,a)\sim d^{p,\pi}}[r(s,a)]$, where $d^{p,\pi}(s,a)$ denotes the state-action visitation distribution under dynamics $p$ and policy $\pi$. To illustrate the impact of model bias in the dual formulation, we similarly examine the scenario where the training policy $\pi$ is applied within the model dynamics $q$. The training objective $\mathcal{J}(\pi)$ is derived as follows:
$\mathcal{J}(\pi)=\mathbb{E}_{(s,a)\sim d^{q,\pi}}[r(s,a)]=\mathbb{E}_{(s,a)\sim d^{p,\pi}}[r(s,a)\frac{d^{q,\pi}(s,a)}{d^{p,\pi}(s,a)}].$
Model bias can result in inaccuracies in estimating the state-action visitation distribution, subsequently affecting the accuracy of value function estimation. Similarly, if the learned model overestimates or underestimates the state-action visitation distribution, it will lead to corresponding overestimations or underestimations of rewards, thereby distorting the objective.
Trajectory Formulation of Policy Shift.
Similarly, we consider the trajectory formulation of policy shift. We execute the sampling policy $\pi_{c}$ within the environment transition dynamics $p$. Then the training objective $\mathcal{J}(\pi)$ is then given by:
$\mathcal{J}(\pi)=\mathbb{E}_{\tau\sim p^{\pi_{c}}}[R(\tau)]=\mathbb{E}_{\tau\sim p^{\pi}}\left[R(\tau)\prod\limits_{t=0}^{\infty}\frac{\pi_{c}(a_{t}|s_{t})}{\pi(a_{t}|s_{t})}\right].$
The policy shift can influence the value estimation of trajectories, imposing implicit constraints on policy learning. For a high-reward trajectory $\tau$, if the current policy $\pi$ closely approximates the optimal policy and satisfies $\prod_{t}\pi(a_{t}\mid s_{t})>\prod_{t}\pi_{c}(a_{t}\mid s_{t})$, it may lead to an underestimation of the return for trajectory $\tau$. Conversely, for a low-reward trajectory $\tau$, if the current policy is relatively effective and satisfies $\prod_{t}\pi(a_{t}\mid s_{t})<\prod_{t}\pi_{c}(a_{t}\mid s_{t})$, this can result in an overestimation of the return for trajectory $\tau$. Both of these impede the effective learning of the policy.
A2. The Proof of Theorem 1
Theorem 1.
Let $\mathcal{L}(\pi)$ be the training objective defined in (6) and $\tilde{r}$ denote the shifts-aware reward (5). Then,
$\log\left[\mathcal{J}_{M}(\pi)\right]\geq(1-\gamma)\mathcal{L}(\pi),\quad\forall\ \pi.$
Proof.
The cumulative discounted reward $R(\tau)$ of trajectory $\tau=(s_{0},a_{0},r_{0},s_{1},...)$ can be written as:
$R(\tau)=\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})=\frac{1}{1-\gamma}\sum_{t=0}^{\infty}(1-\gamma)\gamma^{t}r(s_{t},a_{t}).$ (9)
As discussed in the main text, the probability distribution over a trajectory $\tau$ can be formulated as follows:
$p^{\pi}(\tau)=\mu_{0}(s_{0})\prod\limits_{t=0}^{\infty}p(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t})\text{ , }q^{\pi_{c}}(\tau)=\mu_{0}(s_{0})\prod\limits_{t=0}^{\infty}q(s_{t+1}|s_{t},a_{t})\pi_{c}(a_{t}|s_{t}).$ (10)
Then according to (9) and (10) and Jensen’s inequality, we have
$\log\left[\mathcal{J}_{M}(\pi)\right]$ $=\log\mathbb{E}_{p^{\pi}(\tau)}R(\tau)$
$=\log\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[R(\tau)\frac{p^{\pi}(\tau)}{q^{\pi_{c}}(\tau)}\right]$
$=\log\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[R(\tau)\frac{\mu(s_{0})\prod_{t=0}^{\infty}p(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t})}{\mu(s_{0})\prod_{t=0}^{\infty}q(s_{t+1}|s_{t},a_{t})\pi_{c}(a_{t}|s_{t})}\right]$
$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\log R(\tau)+\sum_{t=0}^{\infty}\left(\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right)\right]$
$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\log\frac{1}{1-\gamma}+\sum_{t=0}^{\infty}(1-\gamma)\gamma^{t}\log r(s_{t},a_{t})+\sum_{t=0}^{\infty}\left(\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right)\right]$
$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\sum_{t=0}^{\infty}\left((1-\gamma)\gamma^{t}\log r(s_{t},a_{t})+\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right)\right]$
$=(1-\gamma)\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(\log r(s_{t},a_{t})+\frac{1}{(1-\gamma)\gamma^{t}}\left(\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right)\right)\right]$
$=(1-\gamma)\mathcal{L}(\pi).$
A3. Classifiers Estimation Derivations
We assume that all steps in the derivation are appropriate, that the denominator is non-zero, and that the argument of the logarithm is not zero.
Transition Classifier Derivation.
Define two events $\mathbb{E}$ and $\mathbb{M}$, where
[…]
The transition classifier is defined as $C_{\phi}(s_{t},a_{t},s_{t+1})=P(\mathbb{E}|s_{t},a_{t},s_{t+1})$. Then
$C_{\phi}(s,a,s_{t+1})=P(\mathbb{E}|s_{t},a_{t},s_{t+1})=$ $\frac{P(\mathbb{E},s_{t},a_{t})P(s_{t+1}|s_{t},a_{t},\mathbb{E})}{P(s_{t},a_{t},s_{t+1})}$
$=\frac{P(\mathbb{E}|s_{t},a_{t})}{P(s_{t},a_{t})}\frac{P(s_{t+1}|s_{t},a_{t},\mathbb{E})}{P(s_{t},a_{t},s_{t+1})}.$ (11)
Events $\mathbb{E}$ and $\mathbb{M}$ form a partition of the probability space, which means that
$P(\mathbb{M}|s_{t},a_{t},s_{t+1})=1-P(\mathbb{E}|s_{t},a_{t},s_{t+1})=1-C_{\phi}(s,a,s_{t+1}).$
Similarly, we have
$1-C_{\phi}(s,a,s_{t+1})$ $=\frac{P(\mathbb{M},s_{t},a_{t})P(s_{t+1}|s_{t},a_{t},\mathbb{M})}{P(s_{t},a_{t},s_{t+1})}$
$=\frac{P(\mathbb{M}|s_{t},a_{t})}{P(s_{t},a_{t})}\frac{P(s_{t+1}|s_{t},a_{t},\mathbb{M})}{P(s_{t},a_{t},s_{t+1})}.$ (12)
According to (11) and (12), we have
$\frac{C_{\phi}(s,a,s_{t+1})}{1-C_{\phi}(s,a,s_{t+1})}=\frac{P(s_{t+1}|s_{t},a_{t},\mathbb{E})}{P(s_{t+1}|s_{t},a_{t},\mathbb{M})}\frac{P(\mathbb{E}|s_{t},a_{t})}{P(\mathbb{M}|s_{t},a_{t})}.$ (13)
Specifically, $P(s_{t+1}|s_{t},a_{t},\mathbb{E})=p(s_{t+1}|s_{t},a_{t})$, where $p$ represents the environment transition dynamic; $P(s_{t+1}|s_{t},a_{t},\mathbb{M})=q(s_{t+1}|s_{t},a_{t})$, where $q$ denotes the model transition dynamic. Then taking the logarithm of both sides of (13), we obtain:
$\log\frac{C_{\phi}(s,a,s_{t+1})}{1-C_{\phi}(s,a,s_{t+1})}=\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{P(\mathbb{E}|s_{t},a_{t})}{P(\mathbb{M}|s_{t},a_{t})}.$ (14)
Given the independence of the state-action pair $(s_{t},a_{t})$ from the transition dynamics, we assume that the visit distribution of $(s_{t},a_{t})$ remains identical. Then
$\frac{P(\mathbb{E}|s_{t},a_{t})}{P(\mathbb{M}|s_{t},a_{t})}\approx\frac{d^{\pi}(s_{t},a_{t})\left|\mathcal{D}_{env}\right|}{d^{\pi}(s_{t},a_{t})\left|\mathcal{D}_{m}\right|}=\frac{\left|\mathcal{D}_{env}\right|}{\left|\mathcal{D}_{m}\right|},$ (15)
where $\left|\mathcal{D}_{env}\right|$ and $\left|\mathcal{D}_{m}\right|$ represent the sizes of the datasets $\mathcal{D}_{env}$ and $\mathcal{D}_{m}$, respectively. Thus, $\frac{P(\mathbb{E}|s_{t},a_{t})}{P(\mathbb{M}|s_{t},a_{t})}$ can be approximated by a constant. Then according to (14) and (15), we have
$\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}\approx\log\frac{C_{\phi}(s,a,s_{t+1})}{1-C_{\phi}(s,a,s_{t+1})}-c_{1},$
where $c_{1}$ is a constant that does not influence the training of SAMBO.
Action Classifier Derivation.
The derivation of the action classifier is similar to that of the transition classifier. Define two events $\mathbb{A}$ and $\mathbb{C}$, where
[…]
={Given the state $\mathbb{C}=\left\{\text{Given the state }s_{t},\text{ the agent takes action }a_{t}\text{ according to the behavior policy }\pi_{c}\right\}.$
The action classifier is defined as $C_{\psi}(s_{t},a_{t})=P(\mathbb{A}|s_{t},a_{t})$, Then
$C_{\psi}(s,a)=P(\mathbb{A}|s_{t},a_{t})=\frac{P(a_{t}|s_{t},\mathbb{A})P(\mathbb{A}|s_{t})P(s_{t})}{P(s_{t},a_{t})},$ (16)
$1-C_{\psi}(s,a)=P(\mathbb{C}|s_{t},a_{t})=\frac{P(a_{t}|s_{t},\mathbb{C})P(\mathbb{C}|s_{t})P(s_{t})}{P(s_{t},a_{t})}.$ (17)
Specifically, $P(a_{t}|s_{t},\mathbb{A})=\pi(a_{t}|s_{t})$, $P(a_{t}|s_{t},\mathbb{C})=\pi_{c}(a_{t}|s_{t})$. Then according to (16) and (17), we have
$\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}=\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}.$ (18)
Similarly, $\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}$ can be also approximated by the constant $\frac{\left|\mathcal{D}_{\pi}\right|}{\left|\mathcal{D}_{env}\right|}$. According to the sampling method in Algorithm 2, we have
$\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}=\frac{d(s_{t})\left|\mathcal{D}_{\pi}\right|}{d(s_{t})\left|\mathcal{D}_{env}\right|}=\frac{\left|\mathcal{D}_{\pi}\right|}{\left|\mathcal{D}_{env}\right|},$
where $d(s_{t})$ denotes the distribution of states in the offline dataset, and $\mathcal{D}_{\pi}$ represents the dataset collected by policy $\pi$. Thus, taking the logarithm of both sides of (18), we get
$\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}=\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}-c_{2},$
where $c_{2}$ is a constant that does not influence the training process.
Appendix B B. Implementation Details
We present a detailed outline of the SAMBO algorithm in Algorithm 2.
Algorithm 2 SAMBO-RL
Input: Offline dataset $\mathcal{D}_{env}$, learned dyanmics models $\{m_{\theta}^{i}\}_{i=1}^{N}$, initialized policy $\pi$, shifts-aware reward hyperparameters $\alpha$ and $\beta$, rollout horizon $h$, rollout batch size $b$.
1:  Train the probabilistic dynamics model $m_{\theta}(s^{\prime},r|s,a)=\mathcal{N}(\mu_{\theta}(s,a),\Sigma_{\theta}(s,a))$ on $\mathcal{D}_{env}$.
2:  Initialize the model dataset buffer $\mathcal{D}_{m}\leftarrow\varnothing$.
3:  for $i=1,2,...,N_{iter}$ do
4:     Initialize the policy dataset buffer $\mathcal{D}_{\pi}\leftarrow\varnothing$.
5:     for $1,2,...,b$ (in parallel) do
6:        Sample state $s_{1}$ from $\mathcal{D}_{env}$ for the initialization of the rollout.
7:        for $j=1,2,...,h$ do
8:           Sample an action $a_{j}\sim\pi(s_{j})$.
9:           Randomly pick dynamics $m$ from $\{m_{\theta}^{i}\}_{i=1}^{N}$ and sample $s_{j+1},r_{j}\sim m(s_{j},a_{j})$.
10:           Add sample $(s_{j},a_{j},r_{j},s_{j+1})$ to $\mathcal{D}_{m}$ and $\mathcal{D}_{\pi}$.
11:        end for
12:     end for
13:     Update classifiers $C_{\phi}$ and $C_{\psi}$ according to Eq. (19) and Eq. (20), respectively.
14:     Drawing samples from $\mathcal{D}_{env}\cup\mathcal{D}_{m}$, compute shifts-aware reward according to Eq. (21) and Eq. (22), respectively.
15:     Run SAC with shifts-aware reward to update policy $\pi$.
16:  end for
B1. Model Training
In our approach, the model is represented by a probabilistic neural network that outputs a Gaussian distribution for both the next state and reward, given the current state and action:
$m_{\theta}\left(s_{t+1},r_{t}\mid s_{t},a_{t}\right)=\mathcal{N}\left(\mu_{\theta}\left(s_{t},a_{t}\right),\Sigma_{\theta}\left(s_{t},a_{t}\right)\right).$
Our model training approach is consistent with the methodology used in prior works (Janner et al. 2019; Yu et al. 2020). We train an ensemble of seven dynamics models and select the best five based on their validation prediction error from a held-out set containing 1000 transitions in the offline dataset $\mathcal{D}_{env}$. Each model in the ensemble is a 4-layer feedforward neural network with 200 hidden units. During model rollouts, we randomly choose one model from the best five models.
B2. Classifiers Training
The transition classifier $C_{\phi}(s,a,s^{\prime})\in[0,1]$ represents the probability that the transition originated from the environment. It is trained by minimizing the standard cross-entropy loss function:
$\mathcal{L}_{\phi}=-\mathbb{E}_{\mathcal{D}_{env}}\left[\log C_{\phi}(s,a,s^{\prime})\right]-\mathbb{E}_{\mathcal{D}_{m}}\left[\log(1-C_{\phi}(s,a,s^{\prime}))\right].$ (19)
The action classifier $C_{\psi}(s,a)\in[0,1]$ represents the probability that a given state-action pair $(s,a)$ was generated by the current policy $\pi$. This classifier is trained by minimizing:
$\mathcal{L}_{\psi}=-\mathbb{E}_{\mathcal{D}_{\pi}}\left[\log C_{\psi}(s,a)\right]-\mathbb{E}_{\mathcal{D}_{env}}\left[\log(1-C_{\psi}(s,a))\right].$ (20)
For data in the model dataset $\mathcal{D}_{m}$, the dynamics are governed by the model $q$. Furthermore, we approximate the training data collection policy for $\mathcal{D}_{m}$ using the current policy $\pi$, i.e. $\pi_{c}=\pi$. Therefore, $\forall(s,a,r,s^{\prime})\in\mathcal{D}_{m}$, the shifts-aware reward simplifies to
$\tilde{r}(s,a,s^{\prime})=\log r(s,a)+\alpha\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}.$
Then, for all $(s,a,s^{\prime})\in\mathcal{D}_{m}$, the shift-aware reward can be estimated using the following formulation,
$\tilde{r}(s,a,s^{\prime})\approx\log r(s,a)+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})}.$ (21)
For data in the offline dataset $\mathcal{D}_{env}$, the dynamics are governed by environment $p$, i.e. $q=p$, and the training data collection policy aligns with the behavioral policy, i.e. $\pi_{c}=\pi_{b}$. Therefore, $\forall(s,a,r,s^{\prime})\in\mathcal{D}_{env}$, the shifts-aware reward simplifies to
$\tilde{r}(s,a,s^{\prime})=\log r(s,a)+\beta\log\frac{\pi(a|s)}{\pi_{b}(a|s)},$
which is independent of the subsequent state $s^{\prime}$. Then, for all $(s,a,s^{\prime})\in\mathcal{D}_{env}$, the shift-aware reward can be estimated by
$\tilde{r}(s,a,s^{\prime})\approx\log r(s,a)+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$ (22)
B3. Policy Learning
The policy optimization in our approach is based on the SAC algorithm, with hyperparameter configurations aligning with those employed in MOBILE (Sun et al. 2023). For each update, we draw a batch of 256 transitions, with $5\%$ sourced from the offline dataset $\mathcal{D}_{env}$ and the remaining $95\%$ from the synthetic dataset $\mathcal{D}_{m}$. The detailed hyperparameter settings used for the D4RL benchmark are provided in Table 2.
Hyperparameters Value Description
K 2 The number of critics.
$\gamma$ 0.99 Discount factor.
$l_{r}$ of actor $1\times 10^{-4}$ Policy learning rate.
$l_{r}$ of critic $3\times 10^{-4}$ Critic learning rate.
Optimizer Adam Optimizers of the actor and critics.
$f$ 0.05 Ratio of the real samples.
Batch size 256 Batch size for each update.
$N_{\text{iter}}$ 3M Total gradient steps.
Table 2: Hyperparameters for policy optimization in SAMBO
Appendix C C. Experimental Details
C1. Benchmark
We conduct experiments on Gym tasks (“v2” version) within the D4RL benchmark and now provide the sources for the reported performance on this benchmark. For TD3+BC (Fujimoto and Gu 2021), EDAC (An et al. 2021), TT (Janner, Li, and Levine 2021), RAMBO (Rigter, Lacerda, and Hawes 2022), and MOBILE (Sun et al. 2023), we cite the scores reported in the original papers, as these evaluations were conducted on the “v2” datasets in Gym. Since COMBO (Yu et al. 2021) does not provide source code, we include the results as reported in the original paper. The results of CQL (Kumar et al. 2020) are taken from the performance table in MOBILE, where CQL was retrained on the “v2” datasets in Gym. The results of MOPO* are obtained from experiments using the OfflineRL-Kit libraryhttps://github.com/yihaosun1124/OfflineRL-Kit on the “v2” datasets. These scores are referenced in Table 1.
C2. Hyperparameters
We list the hyperparameters that have been tuned as follows. The specific settings are detailed in Table 3.
Task Name $\boldsymbol{\alpha}$ $\boldsymbol{\beta}$ $\boldsymbol{h}$ $\boldsymbol{c}$
halfcheetah-random 0.01 0.01 5 -0.2
hopper-random 0.02 0.01 5 0
walker2d-random 0.03 0.03 5 0
halfcheetah-medium 0.01 0.01 5 0
hopper-medium 0.02 0.01 5 0
walker2d-medium 0.035 0.01 5 0
halfcheetah-medium-replay 0.01 0.01 5 0
hopper-medium-replay 0.02 0.01 5 0
walker2d-medium-replay 0.01 0.01 1 -0.2
halfcheetah-medium-expert 0.04 0.01 5 0
hopper-medium-expert 0.04 0.02 5 0
walker2d-medium-expert 0.05 0.05 1 0
Table 3: Hyperparameters of SAMBO used in the D4RL datasets.
Model Bias Coefficient $\alpha$.
We tune $\alpha$ in the range of $[0.01,0.05]$. For training stability, in the walker2d-medium dataset, we fix the transition classifier after 1 million iterations by setting $\alpha=0$. In the walker2d-medium-expert dataset, we apply an annealing factor to the transition classifier. Initially set to 3, the factor remains constant for the first 1 million iterations and then decays reciprocally to 1 over the subsequent 2 million iterations.
Policy Shift Coefficient $\beta$.
We tune $\beta$ in the range of $[0.01,0.05]$.
Rollout Length $h$.
Similar to MOPO, We perform short-horizon branch rollouts in SAMBO. we tune $h$ in the range of $\{1,5\}$ for Gym tasks in the D4RL benchmark.
Reward Truncation Coefficient $c$.
To compute the logarithm of the reward term in the shifts-aware reward, we apply a translation to the real reward, which does not affect policy optimization. Specifically, given the offline dataset $\mathcal{D}_{env}$, let $r_{max}$ denote the maximum reward and $r_{min}$ denote the minimum reward in this dataset. Then we perform a translation and truncation on the real reward as follows:
$r^{\prime}=\min\{1\times 10^{-8},r-c(r_{max}-r_{min})+1\times 10^{-8}\}.$
Then we calculate the shifts-aware reward by
$\tilde{r}(s,a,s^{\prime})\approx\log r^{\prime}(s,a)+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})}+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$
In most datasets, we set $c=0$. However, we observe that employing a negative value for $c$ yields favorable results in certain datasets.
C3. Computing Infrastructure
All experiments are conducted using a single GeForce GTX 3090 GPU and an Intel(R) Xeon(R) Gold 6330 CPU @ 2.00GHz. Our implementation code is based on the OfflineRL-Kit library. The software libraries and frameworks utilized in SAMBO are consistent with those specified in this library.

### v2

@@
 Abstract
-Model-based Offline Reinforcement Learning trains policies based on offline datasets and model dynamics, without direct real-world environment interactions.
-However, this method is inherently challenged by distribution shift.
-Previous approaches have primarily focused on tackling this issue directly leveraging off-policy mechanisms and heuristic uncertainty in model dynamics, but they resulted in inconsistent objectives and lacked a unified theoretical foundation.
-This paper offers a comprehensive analysis that disentangles the problem into two key components: model bias and policy shift.
-We provide both theoretical insights and empirical evidence to demonstrate how these factors lead to inaccuracies in value function estimation and impose implicit restrictions on policy learning.
-To address these challenges, we derive adjustment terms for model bias and policy shift within a unified probabilistic inference framework.
-These adjustments are seamlessly integrated into the vanilla reward function to create a novel Shifts-aware Reward (SAR), aiming at refining value learning and facilitating policy training.
-Furthermore, we introduce Shifts-aware Model-based Offline Reinforcement Learning (SAMBO-RL), a practical framework that efficiently trains classifiers to approximate the SAR for policy optimization.
-Empirically, we show that SAR effectively mitigates distribution shift, and SAMBO-RL demonstrates superior performance across various benchmarks, underscoring its practical effectiveness and validating our theoretical analysis.
-footnotetext: footnotetext: Preprint. Under Review.
-Introduction
-Offline reinforcement learning (RL) (Lange, Gabel, and Riedmiller 2012; Levine et al.
-2020) learns policies from a pre-collected dataset generated by a behavior policy, avoiding additional online interaction with the environment.
-This approach showcases remarkable potential in data-driven decision scenarios (Emerson, Guy, and McConville 2023; Sinha, Mandlekar, and Garg 2022) where the exploration is costly or hazardous.
-The model-based framework (Luo et al.
-2018) is particularly effective for offline RL (Yu et al.
-2020, 2021; Sun et al.
-2023; Li et al.
-2024).
-It involves learning models of the environment from the given dataset and utilizing these models to generate data for policy training.
-Nevertheless, directly learning a policy using offline and model-generated data introduces the distribution shift challenge.
+Model-based offline reinforcement learning trains policies using pre-collected datasets and learned environment models, eliminating the need for direct real-world environment interaction.
+However, this paradigm is inherently challenged by distribution shift (DS).
+Existing methods address this issue by leveraging off-policy mechanisms and estimating model uncertainty, but they often result in inconsistent objectives and lack a unified theoretical foundation.
+This paper offers a comprehensive analysis that disentangles the problem into two fundamental components: model bias and policy shift.
+Our theoretical and empirical investigations reveal how these factors distort value estimation and restrict policy optimization.
+To tackle these challenges, we derive a novel Shifts-aware Reward (SAR) through a unified probabilistic inference framework, which modifies the vanilla reward to refine value learning and facilitate policy training.
+Building on this, we introduce Shifts-aware Model-based Offline Reinforcement Learning (SAMBO-RL), a practical framework that efficiently trains classifiers to approximate SAR for policy optimization.
+Empirical experiments show that SAR effectively mitigates DS, and SAMBO-RL achieves superior or comparable performance across various benchmarks, underscoring its effectiveness and validating our theoretical analysis.
+Machine Learning, ICML
+1 Introduction
+Offline reinforcement learning (RL) (Lange et al., 2012; Levine et al., 2020) learns policies from the offline dataset generated by a behavior policy, avoiding additional online interaction with the environment.
+This approach shows remarkable potential in data-driven decision scenarios (Emerson et al., 2023; Sinha et al., 2022) where exploration is costly or hazardous.
+The model-based framework (Luo et al., 2018) is particularly effective for offline RL (Yu et al., 2020, 2021; Sun et al., 2023; Li et al., 2024).
+It involves learning models of the environment from the offline dataset and utilizing these models to generate data for policy training.
+However, directly learning a policy using offline and synthetic data introduces the distribution shift challenge.
 This shift causes the training objective to deviate from the true objective of RL, leading to poor test performance.
 Prior methods in model-based offline RL focus on directly leveraging off-policy methods in online RL and heuristic model uncertainty for conservative learning to mitigate the impact of distribution shift.
-For instance, MOPO (Yu et al.
-2020) uses the predicted variance of the learned model as an uncertainty penalty on rewards.
-MOBILE (Sun et al.
-2023) penalizes value learning of model data according to the uncertainty of off-policy methods in online RL and the Bellman Q-function estimated by ensemble models.
-RAMBO (Rigter, Lacerda, and Hawes 2022) considers the trajectory distribution in the worst case and adversarially trains the policy and transition model while ensuring accurate predictions.
+For instance, MOPO (Yu et al., 2020) uses the predicted variance of the learned model as an uncertainty penalty on rewards.
+MOBILE (Sun et al., 2023) penalizes value learning of model data according to the uncertainty of the Bellman Q-function estimated by ensemble models.
+RAMBO (Rigter et al., 2022) considers the trajectory distribution in the worst case and adversarially trains the policy and transition model while ensuring accurate predictions.
 However, these approaches have resulted in a biased training objective due to the model uncertainty and remain the distribution shift problem to be solved.
+Figure 1: The negative impacts of model bias and policy shift.
+(a) In the 1D grid world, the agent can take two actions: move left (L) or move right (R) from the current state $S$ to the adjacent grid cells $S-1$ or $S+1$, with movement obstructed at the boundaries.
+The initial state is uniformly distributed, and the agent receives undiscounted rewards upon reaching specific states.
+(b) The left panel illustrates value estimates during training and the final policy convergence with and without model bias.
+Even slight model bias can significantly distort value function estimates, leading the learned policies to deviate from the true optimal policy in the actual environment.
+(c) The right panel showcases expected returns and the KL divergence between the learned policy $\pi$ and the behavior policy $\pi_{b}$ during training with and without policy shift.
+We find that policy shift implicitly restricts $\pi$ to remain close to $\pi_{b}$, thereby slowing policy convergence.
 In this paper, we demystify the distribution shift problem in model-based RL and study its fundamental properties.
@@
 In contrast, previous studies heuristically mitigate these two issues from off-policy algorithms and model uncertainty.
-Figure 1: The negative impacts of model bias and policy shift.
-(a) In the 1D grid world, the agent can take two actions: move left (L) or move right (R) to the adjacent grid, with movement obstructed at the boundaries.
-The initial state is uniformly distributed, and the agent receives rewards upon reaching specific states.
-(b) The left panel illustrates value estimates during training and the final policy convergence with and without model bias.
-Even slight model bias can significantly distort value function estimates, leading the learned policies to deviate from the true optimal policy in the actual environment.
-(c) The right panel showcases expected returns and the KL divergence between the learned policy $\pi$ and the behavior policy $\pi_{b}$ during training with and without policy shift.
-We find that policy shift implicitly restricts $\pi$ to remain close to $\pi_{b}$, thereby slowing policy convergence.
 To address these challenges, we theoretically derive adjustments for model bias and policy shift by employing a unified probabilistic inference framework, and introduce a shifts-aware reward, which integrates the vanilla reward with adjustments for model bias and policy shift.
@@
 This approach incorporates model bias adjustment for model-generated data and policy shift modification for offline data.
-Furthermore, it learns a transition classifier and an action classifier to approximate the shifts-aware reward and then performs policy optimization based on this adjusted reward.
+Furthermore, it learns a transition classifier and action classifiers to approximate the shifts-aware reward and then performs policy optimization based on this adjusted reward.
 Our contributions include:
-  • We demystify the distribution shift problem in model-based offline RL, i.e., model bias distorts the estimation of value function, while policy shift imposes an implicit policy limitation to hinder policy convergence.
-  • We upgrade the vanilla reward into a shifts-aware reward with model bias adjustment and policy shift modification.
-This shifts-aware reward is used for policy learning to mitigate distribution shift in model-based offline RL.
+  • We demystify the distribution shift problem in model-based offline RL, i.e., model bias distorts the estimation of the value function, while policy shift imposes an implicit policy limitation that hinders policy convergence.
+  • We upgrade the vanilla reward into a shifts-aware reward incorporating model bias adjustment and policy shift modification.
+This modified reward is used for policy learning to mitigate distribution shift in model-based offline RL.
   • We devise SAMBO-RL, which trains transition and action classifiers to approximate the shifts-aware reward, effectively addressing both model bias and policy shift.
-  • We show that shifts-aware reward effectively mitigates distribution shift, and SAMBO-RL demonstrates superior performance across various benchmarks.
-Related Work
-Model-free Offline RL encompasses two primary approaches: policy constraints and value regularization, both incorporating conservatism (Jin, Yang, and Wang 2021).
+  • We show that shifts-aware reward effectively mitigates distribution shift, and SAMBO-RL demonstrates superior or comparable performance across various benchmarks.
+2 Related Work
+Model-free Offline RL encompasses two primary approaches: policy constraints and value regularization, both incorporating conservatism (Jin et al., 2021).
 Policy constraint methods aim to keep the learned policy close to the behavioral policy, mitigating distribution shift.
-For instance, BEAR (Kumar et al.
-2019) and BRAC (Wu, Tucker, and Nachum 2019) constrain the learned policy by minimizing different divergences.
-BCQ (Fujimoto, Meger, and Precup 2019) perturbs actions during learning to ensure they remain within the action space of the offline dataset by learning a generative model of the behavior policy.
+For instance, BEAR (Kumar et al., 2019) and BRAC (Wu et al., 2019) constrain the learned policy by minimizing different divergences.
+BCQ (Fujimoto et al., 2019) perturbs actions during learning to ensure they remain within the action space of the offline dataset by learning a generative model of the behavior policy.
 On the other hand, value regularization methods add conservative regularization terms to the value optimization objective, avoiding out-of-distribution (OOD) actions.
-CQL (Kumar et al.
-2020) penalizes the Q-values of OOD samples by incorporating regularization into the value loss function.
-TD3+BC (Fujimoto and Gu 2021) introduces a behavior cloning regularization into the TD3 (Fujimoto, Hoof, and Meger 2018) objective.
-Similarly, EDAC (An et al.
-2021) and PBRL (Bai et al.
-2021) penalize the Q-function based on uncertainty derived from an ensemble of Q-networks.
-MCQ (Lyu et al.
-2022) actively trains OOD actions by constructing pseudo-target values to alleviate over-pessimism.
-Model-based Offline RL approximates the environment using learned models and performs conservative policy optimization (Lu et al.
-2021), achieving high data efficiency (Li et al.
-2024).
-MOPO (Yu et al.
-2020) incorporates conservatism by penalizing rewards based on the uncertainty of model predictions.
-MOReL (Kidambi et al.
-2020) introduces a pessimistic MDP to penalize rewards of state-action pairs in unexplored regions.
-COMBO (Yu et al.
-2021) extends CQL to the model-based setting, regularizing the value function on OOD samples.
-RAMBO (Rigter, Lacerda, and Hawes 2022) adopts an adversarial approach, training the policy and model within a robust framework while ensuring accurate transition predictions of the model.
-CBOP (Jeong et al.
-2022) adaptively weights multi-step returns in a model-based value expansion framework (Feinberg et al.
-2018).
-Lastly, MOBILE (Sun et al.
-2023) introduces penalties during value learning by leveraging the uncertainty of the Bellman Q-function estimates derived from ensemble models.
-Preliminaries
+CQL (Kumar et al., 2020) penalizes the Q-values of OOD samples by incorporating regularization into the value loss function.
+TD3+BC (Fujimoto & Gu, 2021) introduces a behavior cloning regularization into the TD3 (Fujimoto et al., 2018) objective.
+Similarly, EDAC (An et al., 2021) and PBRL (Bai et al., 2021) penalize the Q-function based on uncertainty derived from an ensemble of Q-networks.
+MCQ (Lyu et al., 2022) actively trains OOD actions by constructing pseudo-target values to alleviate over-pessimism.
+Model-based Offline RL approximates the environment using learned models and performs conservative policy optimization (Lu et al., 2021), achieving high data efficiency (Li et al., 2024).
+MOPO (Yu et al., 2020) incorporates conservatism by penalizing rewards based on the uncertainty of model predictions.
+MOReL (Kidambi et al., 2020) introduces a pessimistic MDP to penalize rewards of state-action pairs in unexplored regions.
+COMBO (Yu et al., 2021) extends CQL to the model-based setting, regularizing the value function on OOD samples.
+TT (Janner et al., 2021) uses a transformer to model offline trajectories and employs beam search for planning.
+RAMBO (Rigter et al., 2022) adopts an adversarial approach, training policy and model within a robust framework while ensuring accurate transition predictions of the model.
+CBOP (Jeong et al., 2022) adaptively weights multi-step returns in a model-based value expansion framework (Feinberg et al., 2018).
+Lastly, MOBILE (Sun et al., 2023) introduces penalties during value learning by leveraging the uncertainty of the Bellman Q-function estimates derived from ensemble learned dynamic models.
+3 Preliminaries
 MDP.
@@
 MBPO.
-Similar to previous works (Yu et al.
-2020, 2021; Rigter, Lacerda, and Hawes 2022; Sun et al.
-2023), we focus on model-based policy optimization (MBPO) (Janner et al.
-2019), which employs an actor-critic RL algorithm.
+Similar to previous works (Yu et al., 2020, 2021; Rigter et al., 2022; Sun et al., 2023), we focus on model-based policy optimization (MBPO) (Janner et al., 2019), which employs an actor-critic RL algorithm.
 MBPO first performs $k$-step rollouts using the model $m(s^{\prime}|s,a)$ starting from state $s\in\mathcal{D}_{env}$ and adds the generated data to a replay buffer $\mathcal{D}_{m}$.
@@
 where $R(\tau)=\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})$ represents the discounted return of the whole trajectory $\tau=(s_{0},a_{0},r_{0},s_{1},...)$.
-Shifts-aware Policy Optimization
+4 Shifts-aware Policy Optimization
 We first formalize the distribution shift problem as shift weighting and further delve into its negative effects on policy optimization from theory and experiments.
 To address these challenges, we employ probabilistic inference to derive a surrogate objective and formulate a shifts-aware reward.
-Finally, we implement this theoretical framework within model-based offline RL and introduce SAMBO-RL, a practical algorithm designed to mitigate distribution shift.
-Demystifying Distribution Shift
+Finally, we implement this theoretical framework within model-based offline RL and introduce SAMBO-RL, a practical algorithm designed to mitigate distribution shifts.
+4.1 Demystifying Distribution Shift
 We examine the distribution shift problem from an optimization perspective and identify its origins as model bias and policy shift.
-Furthermore, we explore how these factors affect accuracy in value estimation and impose implicit policy restrictions, resulting in a biased objective and slow convergence.
-The supplementary analysis is in Appendix A1.
-Shift Weighting.
-The training objective $\mathcal{J}(\pi)$ in (2) can be expressed as the following trajectory formulation:
+Then we explore how these factors affect accuracy in value estimation and impose implicit policy restrictions, resulting in a biased objective and slow convergence.
+The supplementary analysis is in Appendix A.1.
+Shift Weighting. The training objective $\mathcal{J}(\pi)$ in (2) can be expressed as the following trajectory formulation:
 $\mathcal{J}(\pi)=\mathbb{E}_{\tau\sim p^{\pi}(\tau)}\left[\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}R(\tau)\right].$ (3)
 Compared to (2), it indicates that for a given trajectory $\tau$, the true objective under practical sampling distributions differs in shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$ from the training objective, which formalizes the distribution shift in trajectory formulation.
-Substituting (1) into the shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$, we can attribute this weighting to two primary components: model bias and policy shift, reformulated as the following equation:
+Substituting (1) into the shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$, we can attribute this weighting to two primary components: model bias and policy shift, reformulated as the following:
 $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}=\underbrace{\prod_{t}\frac{q(s_{t+1}|s_{t},a_{t})}{p(s_{t+1}|s_{t},a_{t})}}_{\text{model bias}}\underbrace{\prod_{t}\frac{\pi_{c}(a_{t}|s_{t})}{\pi(a_{t}|s_{t})}}_{\text{policy shift}}.$ (4)
 Model bias illustrates the discrepancy between the dynamics used for training and those of the actual environment, while policy shift represents the divergence between the policies used for sampling training data and those being learned.
-Negative Impact of Shift Weighting.
-We analyze the impact of shift weighting from model bias and policy shift.
-(i) To illustrate the effect of model bias, we examine the scenario where the training policy $\pi$ operates within model dynamics $q$, which diverges from actual environment dynamics $p$.
+Negative Impact of Shift Weighting. We analyze the impact of shift weighting from model bias and policy shift.
+(i) To illustrate the effect of model bias, we examine the scenario where the training policy $\pi$ operates within model dynamics $q$, which diverges from the actual environment dynamics $p$.
 In this context, the training objective $\mathcal{J}(\pi)$ is:
@@
 Compared to the true objective (3), the compounded prediction bias of the model significantly affects value estimation and leads to a distorted objective.
-Specifically, when the model $q$ overestimates or underestimates the likelihood of generating trajectory $\tau$, it consequently overestimates or underestimates the expected return of this trajectory in the real environment.
-This essentially introduces a bias for the optimization objective, resulting in wrong training directions.
+Specifically, when the model $q$ overestimates or underestimates the probability of generating trajectory $\tau$, it consequently overestimates or underestimates the expected return of this trajectory in the real environment.
+This essentially introduces bias for the optimization objective, resulting in wrong training directions.
 (ii) To delve deeper into the policy optimization process under policy shift, we consider the setting where a sampling policy $\pi_{c}$ is executed within the true environment transition dynamics $p$.
@@
 This indicates that policy shift does compel $\pi$ to align closely with $\pi_{b}$, thereby impeding the training process and substantially slowing down the policy convergence.
-These analyses and empirical evidence highlight the critical role of addressing model bias and policy shift in achieving accurate value learning and efficient policy training.
-Shifts-aware Reward via Probabilistic Inference
+4.2 Shifts-aware Reward via Probabilistic Inference
 This subsection introduces a novel shifts-aware reward through a unified probabilistic inference framework, serving as a lower-bound surrogate for the true objective.
-We further examine how this modified reward contributes to refining the value function and motivating $\pi$ to escape the limitations of behavior policy, thus mitigating model bias and policy shift.
+We examine how this modified reward contributes to refining the value function and motivating $\pi$ to escape the limitations of behavior policy, thus mitigating model bias and policy shift.
 Shifts-aware Reward.
@@
 Additionally, we establish that $\mathcal{L}(\pi)$ serves as a lower bound for the true objective, ensuring its rationality as a surrogate.
-Theorem 1.
+Theorem 4.1.
 Let $\mathcal{L}(\pi)$ be the training objective defined in (6) and $\tilde{r}$ denote the shifts-aware reward (5). Then,
@@
 This modification guides $\pi$ to break out of the confines of the behavior policy $\pi_{b}$, promoting better exploration and more effective learning.
-Shifts-aware Model-based Offline RL
+4.3 Shifts-aware Model-based Offline RL
 We apply the shifts-aware reward in model-based offline RL and develop the practical framework, SAMBO-RL.
@@
 These weighting factors ensure the reward and adjustments are properly balanced in transition-based learning scenarios, which is empirically validated in the following experiments.
-Furthermore, since the environment dynamics $p$ and behavior policy $\pi_{b}$ are typically unknown and cannot be exactly computed, we derive two classifiers to estimate the shifts-aware reward (Eysenbach et al.
-2020, 2022).
-Details are shown in Appendix A3.
+Furthermore, since the environment dynamics $p$ and behavior policy $\pi_{b}$ are typically unknown and cannot be exactly computed, we derive two classifiers to estimate the shifts-aware reward (Eysenbach et al., 2020, 2022).
+Details are shown in Appendix A.3.
 Specifically, these classifiers estimate model bias adjustment $\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}$ and policy shift modification $\log\frac{\pi(a|s)}{\pi_{c}(a|s)}$ respectively.
@@
 The complete procedure of SAMBO-RL is outlined in Algorithm 1, with further details provided in Appendix B.
-Algorithm 1 SAMBO-RL
-Input: Offline dataset $\mathcal{D}_{env}$, learned dyanmics models $\{m_{\theta}^{i}\}_{i=1}^{N}$, initialized policy $\pi$, shifts-aware reward hyperparameters $\alpha$ and $\beta$.
-1:  Train the probabilistic dynamics model $m_{\theta}(s^{\prime},r|s,a)=\mathcal{N}(\mu_{\theta}(s,a),\Sigma_{\theta}(s,a))$ on $\mathcal{D}_{env}$.
-2:  Initialize the model dataset buffer $\mathcal{D}_{m}\leftarrow\varnothing$.
-3:  for $i=1,2,...,N_{iter}$ do
-4:     Initialize the policy dataset buffer $\mathcal{D}_{\pi}\leftarrow\varnothing$.
-5:     Generate synthetic rollouts by model $m_{\theta}$. Add transition data in these rollouts to $\mathcal{D}_{m}$ and $\mathcal{D}_{\pi}$.
-6:     Update classifiers $C_{\phi}$ and $C_{\psi}$ according to Eq. (7).
-7:     Compute shifts-aware reward according to Eq. (8).
-8:     Run SAC (Haarnoja et al. 2018) with shifts-aware reward to update policy $\pi$.
-9:  end for
-Experiments
+5 Experiments
 In our experiments, we focus on three objectives: (1) validating the effectiveness of shifts-aware reward, (2) comparing the performance of SAMBO with previous methods, and (3) analyzing the contribution of each component and hyperparameter within SAMBO to its overall performance.
 We evaluate our algorithm using standard offline RL benchmarks in Mujoco environments and conduct an ablation study to assess its performance.
-Our implementation is based on the OfflineRL-Kit library***https://github.com/yihaosun1124/OfflineRL-Kit, a comprehensive and high-performance library for implementing offline reinforcement learning algorithms.
+Our implementation is based on the OfflineRL-Kit library [^1] https://github.com/yihaosun1124/OfflineRL-Kit, a comprehensive and high-performance library for implementing offline reinforcement learning algorithms.
 The basic parameters of our algorithm are consistent with the settings in this library.
-Effectiveness of Shifts-aware Reward
+5.1 Effectiveness of Shifts-aware Reward
 (a) Model Bias Adjustment
@@
 (b) Policy shift modification addresses implicit policy restrictions, leading to faster convergence.
-We conduct an evaluation of the effectiveness of the Shifts-aware Reward in addressing model bias and policy shift issues within the 1D grid environment illustrated in Fig. 1 (a).
+We conduct an evaluation of the effectiveness of the shifts-aware Reward in addressing model bias and policy shift issues within the 1D grid world illustrated in Fig. 1 (a).
+Table 1: Results on the D4RL Gym benchmark.
+MOPO’s results are obtained on the “v0” datasets; MOPO*’s results are obtained from experiments in the OfflineRL-Kit library on the “v2” datasets.
+The numbers reported for SAMBO are the normalized scores averaged over the final iteration of training across 4 seeds, with $\pm$ standard deviation.
+The top three scores are highlighted in bold.
+We use an asterisk “*” for the best score, an underscore “_” for the second, and bold text for the third.
+Task Name BC CQL TD3+BC EDAC MOPO MOPO* COMBO TT RAMBO MOBILE SAMBO (Ours)
+halfcheetah-random 2.2 31.3 11.0 28.4 35.4 38.5 38.8 6.1 39.5 39.3 39.7$\pm$2.0
+hopper-random 3.7 5.3 8.5 25.3 11.7 31.7 17.9 6.9 25.4 31.9 32.4$\pm$0.5
+walker2d-random 1.3 5.4 1.6 16.6 13.6 7.4 7.0 5.9 0.0 17.9 8.1 $\pm$ 7.8
+halfcheetah-medium 43.2 46.9 48.3 65.9 42.3 72.4 54.2 46.9 77.9 74.6 72.5$\pm$3.8
+hopper-medium 54.1 61.9 59.3 101.6 28.0 62.8 97.2 67.1 87.0 106.6 99.7$\pm$0.3
+walker2d-medium 70.9 79.5 83.7 92.5 17.8 84.1 81.9 81.3 84.9 87.7 78.3$\pm$2.6
+halfcheetah-medium-replay 37.6 45.3 44.6 61.3 53.1 72.1 55.1 44.1 68.7 71.7 70.6$\pm$1.6
+hopper-medium-replay 16.6 86.3 60.9 110.0 67.5 92.7 89.5 99.4 99.5 103.9 101.5$\pm$1.5
+walker2d-medium-replay 20.3 76.8 81.8 87.1 39.0 85.9 56.0 82.6 89.2 89.9 92.3$\pm$ 1.5
+halfcheetah-medium-expert 44.0 95.0 90.0 106.3 63.3 83.6 90.0 95.0 95.4 108.2 97.8 $\pm$0.3
+hopper-medium-expert 53.9 96.9 98.0 110.7 23.7 74.6 111.0 110.0 88.2 112.6 111.0 $\pm$1.2
+walker2d-medium-expert 90.1 109.1 110.1 114.7 44.6 108.2 103.3 101.9 56.7 115.2 106.1 $\pm$ 2.9
 Model Bias Adjustment.
 Model bias exacerbates inaccuracies in value estimates, ultimately resulting in a distorted optimization objective.
-To evaluate the effectiveness of model bias adjustment, we examine its amendment to the training process.
+To evaluate the effects of model bias adjustment on the training process, we conduct experiments.
 As illustrated in Fig.
-2 (a), we execute the policy gradient algorithm under four settings: using the vanilla reward with the overestimating model (OM), the shifts-aware reward with the overestimating model (SAR with OM), the vanilla reward with the underestimating model (UM), and the shifts-aware reward with the underestimating model (SAR with UM).
-We then measure the expected return of the learned policy after each update during training in the actual environment.
+2 (a), we execute the policy gradient algorithm under four settings: using the vanilla reward with the overestimating model (OM) or underestimating model (UM), the shifts-aware reward with the overestimating model (SAR with OM) underestimating model (SAR with UM).
+We then evaluate the expected return of the learned policy after each training update in the actual environment.
 The results show that while the OM initially achieves optimal returns, it rapidly diverges from the optimal policy.
-In contrast, SAR with OM effectively counters the performance degradation caused by model overestimation, thereby sustaining optimal returns.
-Similarly, SAR with UM mitigates the policy suboptimality due to model underestimation, achieving near highest return.
+In contrast, SAR with OM effectively counters the performance degradation caused by model overestimation, sustaining optimal returns.
+Similarly, SAR with UM mitigates the policy suboptimality due to model underestimation, achieving a return near the theoretical optimum.
+(a) MOBILE
+(b) SAMBO
+Figure 3: The $(s,a,s^{\prime})$ distributions of offline RL algorithms result in the Hopper-M task.
+Model Data represents the distribution of data generated by executing the final learned policy in the dynamic model, while Real Data represents the distribution of data generated by executing the final learned policy in the environment.
 Policy Shift Modification.
 Policy shift implicitly restricts policy updates.
-To assess the effectiveness of the policy shift modification, we apply the policy gradient algorithm to the offline objective $\mathcal{J}(\pi)=\mathbb{E}_{s\sim d^{p,\pi_{b}}}\left[V^{\pi}(s)\right]$ in the environment depicted in Fig.
+To assess the efficacy of the policy shift modification, we apply the policy gradient algorithm to the offline objective $\mathcal{J}(\pi)=\mathbb{E}_{s\sim d^{p,\pi_{b}}}\left[V^{\pi}(s)\right]$ in the environment depicted in Fig.
 1 (a).
-We test four settings: using either the vanilla reward or the shifts-aware reward, combined with a behavior policy that is either uniformly distributed or the opposite of the optimal policy.
+We test four settings: the vanilla reward or SAR, combined with a behavior policy that is either uniformly distributed or the opposite of the optimal policy.
 We measure the KL divergence between the learned policy and the behavior policy after each update.
@@
 2 (b), the algorithm using the vanilla reward is limited by the behavior policy, serving a lower KL divergence and slower convergence.
-In contrast, the algorithm employing the shifts-aware reward effectively motivates policy to be aggressive and optimistic and overcomes this restriction, displaying a higher and more stable divergence and achieving faster policy convergence.
-Results on the D4RL Benchmark
+In contrast, the algorithm employing the SAR effectively motivates policy to be optimistic.
+It overcomes the restriction, displaying a higher and more stable divergence and achieving faster convergence.
+Effectiveness in Complex Tasks.
+We also conducted experiments on complex tasks.
+Fig.
+3 shows a visual comparison in the NeoRL task (details in Sec.
+5.2.2).
+SAMBO effectively constrains model-generated data (i.e., Model Data) to align closely with the Offline Data, whereas the previous state-of-the-art method, MOBILE, imposes weaker constraints.
+Furthermore, the real-world data (i.e., Real Data) generated by SAMBO is well-aligned with the Offline Data and remains tightly concentrated, indicating the SAMBO’s effectiveness.
+In contrast, the Real Data from MOBILE deviates significantly from both the Offline and Model Data, still suffering from distribution shift, leading to diminished performance.
+We attribute this discrepancy to SAMBO’s robust handling of model bias and policy shift, ensuring that the Real Data, Model Data, and Offline Data remain closely aligned, effectively mitigating the distribution shift.
+5.2 Benchmark Results
+5.2.1 D4RL
 Datasets.
-In line with prior research, SAMBO is evaluated using the D4RL benchmark (Fu et al.
-2020) on the MuJoCo simulator (Todorov, Erez, and Tassa 2012).
+In line with prior research, SAMBO is evaluated using the D4RL benchmark (Fu et al., 2020) on the MuJoCo simulator (Todorov et al., 2012).
 The evaluation encompasses 12 datasets spanning three environments (halfcheetah, hopper, walker2d) and four dataset types (random, medium, medium-replay, medium-expert).
@@
 Baselines.
-We compare SAMBO against multiple baselines, including both model-free and model-based offline RL algorithms.
-For model-free approaches, BC (Bain and Sammut 1995; Ross, Gordon, and Bagnell 2011) learns behavior policy, CQL (Kumar et al.
-2020) penalizes the Q-function according to policy change, TD3+BC (Fujimoto and Gu 2021) adds a BC term to regularize the policy, and EDAC (An et al.
-2021) quantifies the uncertainty of the Q-value using an ensemble of Q-function networks.
-For model-based approaches, MOPO (Yu et al.
-2020) penalizes rewards via the predicted variance of models.
-COMBO (Yu et al.
-2021) implements CQL within a model-based framework.
-TT (Janner, Li, and Levine 2021) uses a transformer to model offline trajectories and employs beam search for planning.
-RAMBO (Rigter, Lacerda, and Hawes 2022) adversarially trains the policy and the dynamics model within a robust framework.
-MOBILE (Sun et al.
-2023) penalizes the value learning of model data based on the uncertainty of the Bellman Q-function, derived from an ensemble of models.
+We evaluate SAMBO against several baseline methods, including model-free methods such as BC (Bain & Sammut, 1995; Ross et al., 2011), which learns the behavior policy, CQL (Kumar et al., 2020), TD3+BC (Fujimoto & Gu, 2021), and EDAC (An et al., 2021), as well as model-based methods such as MOPO (Yu et al., 2020), COMBO (Yu et al., 2021), TT (Janner et al., 2021), RAMBO (Rigter et al., 2022), and MOBILE (Sun et al., 2023).
+Table 2: Normalized average returns in NeoRL benchmark over 4 random seeds.
+The results for MOBILE are reported from the original paper, and MOBILE represents the results obtained by executing the official implementation.
+Task Name BC CQL TD3+BC EDAC MOPO MOBILE MOBILE SAMBO (Ours)
+HalfCheetah-L 29.1 38.2 30.0 31.3 40.1 54.7 54.8 39.8 $\pm$ 2.6
+Hopper-L 15.1 16.0 15.8 18.3 6.2 17.4 18.2 19.4 $\pm$ 1.3
+Walker2d-L 28.5 44.7 43.0 40.2 11.6 37.6 23.7 46.1 $\pm$ 5.7
+HalfCheetah-M 49.0 54.6 52.3 54.9 62.3 77.8 79.0 62.5 $\pm$1.6
+Hopper-M 51.3 64.5 70.3 44.9 1.0 51.1 43.4 67.6 $\pm$ 12.2
+Walker2d-M 48.7 57.3 58.6 57.6 39.9 62.2 60.1 63.4 $\pm$ 4.0
+HalfCheetah-H 71.3 77.4 75.3 81.4 65.9 83.0 71.8 71.6 $\pm$ 13.9
+Hopper-H 43.1 76.6 75.3 52.5 11.5 87.8 42.3 93.9 $\pm$ 15.2
+Walker2d-H 72.6 75.3 69.6 75.5 18.0 74.9 71.9 79.2 $\pm$ 0.6
+(a) Hopper-L
+(b) Hopper-M
+(c) Hopper-H
+Figure 4: Performance comparison of Hopper tasks in NeoRL benchmark.
 Comparison Results.
 Results are presented in Table 1.
-Each number represents the normalized score, calculated as 100×100$\times$100 ×(score -- random policy score)/ (expert policy score -- random policy score) (Fu et al.
-2020).
-SAMBO achieves the highest scores on two random datasets and demonstrates strong performance across both medium and medium-replay datasets.
-Additionally, SAMBO shows competitive results on medium-expert datasets.
+Each number represents the normalized score, calculated as 100×100$\times$100 ×(score -- random policy score)/ (expert policy score -- random policy score) (Fu et al., 2020).
+SAMBO achieves strong performance on random datasets and demonstrates competitive performance across medium, medium-replay, and medium-expert datasets.
 This performance aligns with our theoretical analysis.
@@
 This suggests that our approach and conservative estimation methods can perform well on complementary datasets.
-Task Name BC CQL TD3+BC EDAC MOPO MOPO* COMBO TT RAMBO MOBILE SAMBO (Ours)
-halfcheetah-random 2.2 31.3 11.0 28.4 35.4 38.5 38.8 6.1 39.5 39.3 39.7$\pm$2.0
-hopper-random 3.7 5.3 8.5 25.3 11.7 31.7 17.9 6.9 25.4 31.9 32.4$\pm$0.5
-walker2d-random 1.3 5.4 1.6 16.6 13.6 7.4 7.0 5.9 0.0 17.9 7.0$\pm$8.2
-halfcheetah-medium 43.2 46.9 48.3 65.9 42.3 72.4 54.2 46.9 77.9 74.6 72.5$\pm$3.8
-hopper-medium 54.1 61.9 59.3 101.6 28.0 62.8 97.2 67.1 87.0 106.6 99.7$\pm$0.3
-walker2d-medium 70.9 79.5 83.7 92.5 17.8 84.1 81.9 81.3 84.9 87.7 78.3$\pm$2.6
-halfcheetah-medium-replay 37.6 45.3 44.6 61.3 53.1 72.1 55.1 44.1 68.7 71.7 70.6$\pm$1.6
-hopper-medium-replay 16.6 86.3 60.9 110.0 67.5 92.7 89.5 99.4 99.5 103.9 101.5$\pm$1.5
-walker2d-medium-replay 20.3 76.8 81.8 87.1 39.0 85.9 56.0 82.6 89.2 89.9 82.8$\pm$0.2
-halfcheetah-medium-expert 44.0 95.0 90.0 106.3 63.3 83.6 90.0 95.0 95.4 108.2 94.8$\pm$2.8
-hopper-medium-expert 53.9 96.9 98.0 110.7 23.7 74.6 111.0 110.0 88.2 112.6 106.5$\pm$5.6
-walker2d-medium-expert 90.1 109.1 110.1 114.7 44.6 108.2 103.3 101.9 56.7 115.2 80.2$\pm$6.8
-Table 1: Results on the D4RL Gym benchmark.
-MOPO’s results are obtained on the “v0” datasets; MOPO*’s results are obtained from experiments in the OfflineRL-Kit library on the “v2” datasets.
-The numbers reported for SAMBO are the normalized scores averaged over the final iteration of training across 4 seeds, with $\pm$ standard deviation.
-The top three scores are highlighted in bold.
-We use an asterisk “*” for the best score, an underscore “_” for the second, and bold text for the third.
-Ablation Studies
-We conducted ablation studies on our method to understand why SAMBO demonstrates competitive performance. The following experimental results are presented in Figure 3.
-(a) Ablation study for SAMBO
-(b) Hyperparameter tuning for $\alpha$
-(c) Hyperparameter tuning for $\beta$
-Figure 3: Illustration of hyperparameter experiments in the HalfCheetah task with a random dataset.
-(a) The effectiveness of model bias adjustment and policy shift modification in the shifts-aware reward.
+5.2.2 NEORL
+Datasets.
+NeoRL (Qin et al., 2022) is a benchmark designed to emulate real-world scenarios by generating datasets through a more conservative policy, aligning closely with real-world data-collection scenarios.
+The datasets, characterized by their narrow scope and limited coverage, present a substantial challenge for offline RL algorithms.
+In this work, we examine nine datasets across three environments (HalfCheetah-v3, Hopper-v3, Walker2d-v3) and three dataset quality levels (L, M, H), corresponding to low, medium, and high-quality datasets.
+NeoRL offers different numbers of trajectories for each task (100, 1000, and 10000), and we consistently selected 1000 trajectories in our test.
+Baselines.
+We compare SAMBO with baseline methods that align with the d4rl experiments in Section 5.2.1, excluding TT, COMBO, and RAMBO, as no results for these methods are reported in their original papers or the NeoRL paper, and determining suitable hyperparameters for them would require an excessive amount of computational time.
+Comparison Results.
+Results are presented in Table 2 and Figure 4.
+We compare the learning curves of model-based approaches in the NeoRL benchmark.
+Since MOPO does not have an official implementation on NeoRL, we report its score based on the original NeoRL paper.
+The experiments show that Our algorithm achieves superior or competitive performance on the more challenging and realistic NeoRL datasets, which can be seen as lower-quality datasets compared to the D4RL datasets.
+In particular, SAMBO demonstrates exceptional performance in the Walker2d and Hopper environments, which, in contrast to the HalfCheetah environment, exhibit higher control complexity, more intricate action dimensions, and greater training difficulty.
+These offline datasets more accurately capture the dynamic control challenges encountered in real-world scenarios, making them significantly demanding and highlighting that SAMBO has strong potential for real-world applications.
+More experiments and discussions are provided in Appendix D.3.
+(a) Ablation in halfcheetah-random dataset
+(b) Tuning for $\alpha$
+(c) Tuning for $\beta$
+(d) Ablation study in Walker2d-H dataset
+Figure 5: Illustration of hyperparameter experiments in D4RL benchmark and NeoRL benchmark.
+(a) and (d) The effectiveness of model bias adjustment and policy shift modification in the shifts-aware reward.
 (b) A large model bias adjustment coefficient $\alpha$ increases conservativeness.
 (c) A large policy shift modification coefficient $\beta$ increases the aggressiveness of exploration.
+5.3 Ablation Studies
+We conducted ablation studies on our method to understand why SAMBO demonstrates competitive performance. The following experimental results are presented in Figure 5.
 Ablation.
 Our shifts-aware reward integrates both the model bias adjustment and the policy shift modification.
-To elucidate the importance of these components, we perform ablation studies on the Halfcheetah task.
-We assess the performance under four configurations: SAMBO, SAMBO without model bias adjustment (wo/mb), SAMBO without policy shift modification (wo/ps), and SAMBO without any modifications (logr), while keeping all other parameters consistent.
-The ablation results are presented in Figure.
-3 (a).
-We find that, compared to the experiment denoted as logr, the performance of wo/mb deteriorates.
+To elucidate the importance of these components, we perform ablation studies on the Halfcheetah-random task in D4RL benchmark and Walker2d-H task in NeoRL benchmark.
+As shown in Figure.
+5 (a) and (d), we assess the performance under four configurations: SAMBO, SAMBO without model bias adjustment (wo/mb), SAMBO without policy shift modification (wo/ps), and SAMBO without any modifications (logr), while keeping all other parameters consistent.
+Figure.
+5 (a) show that, compared to the experiment denoted as logr, the performance of wo/mb deteriorates.
 This degradation arises because wo/mb employs policy shift modification to break the implicit limitations of the behavior policy and encourage more aggressive exploration.
 However, it does not address model bias, resulting in significant inaccuracies in value estimation, ultimately leading to suboptimal performance.
-In contrast, SAMBO incorporates both model bias adjustment and policy shift modification, facilitating aggressive exploration while addressing value estimation challenges, thus alleviating suboptimal performance.
-This underscores the effectiveness of model bias adjustment.
-The ablation experiment wo/ps incorporates model bias adjustment but does not address policy shift.
-As a result, while wo/ps mitigates value misestimation caused by model bias, the lack of policy shift modification causes the algorithm to remain implicitly constrained by the behavioral policy.
-Therefore, wo/ps demonstrates reduced effectiveness in discovering a more optimal policy.
-In contrast, SAMBO integrates policy shift modification during conservative updates, which enables more aggressive exploration and results in superior performance compared to wo/ps.
-This highlights the effectiveness of incorporating policy shift modification.
+This negative effect is particularly evident in the narrow Walker2d-H dataset, as seen in Figure 5 (d).
+Conversely, the wo/ps ablation addresses model bias but neglects policy shift adjustments.
+While this reduces value misestimation, the algorithm remains constrained by the behavior policy due to the lack of policy shift modification.
+Consequently, wo/ps is less effective in discovering a more optimal policy.
+As shown in Figure 5 (a), the performance of wo/ps is inferior to sambo, and Figure 5 (d) illustrates that wo/ps converges slowly compared to sambo.
+In contrast, SAMBO incorporates not only model bias adjustment but also policy shift modification, facilitating aggressive exploration while addressing value estimation challenges, thus alleviating suboptimal performance.
 Hyperparameter Tuning.
 SAMBO is primarily regulated by two key hyperparameters: the model bias adjustment coefficient $\alpha$ and the policy shift modification coefficient $\beta$.
-The model bias adjustment coefficient $\alpha$ regulates the extent of adjustment applied to model-generated data.
-A larger $\alpha$ imposes a significant penalty on the model, leading to excessively conservative utilization of the model.
-Consequently, even in regions where the model performs well, exploration may be restricted due to this excessive penalty, potentially resulting in suboptimal performance.
-Conversely, a smaller $\alpha$ can cause the algorithm to be overly optimistic about exploring risky regions, thereby impairing performance.
-To validate this, we set $\beta=0.01$ and varied $\alpha$ in the Halfcheetah task.
-The results, shown in Fig.
-3 (b), indicate that both very low ($\alpha=0$) and high ($\alpha=0.02$ or $0.05$) values lead to decreased performance.
-Specifically, as $\alpha$ increases from 0 to 0.05, the algorithm initially improves but later declines as it becomes increasingly conservative, which impedes exploration.
-This underscores that both excessively low and high values of $\alpha$ are detrimental to policy learning.
-The policy shift modification coefficient $\beta$ governs the extent of modification applied to the offline data, thereby influencing the degree of deviation from the behavior policy.
-A larger $\beta$ value promotes increased exploratory behavior by encouraging the algorithm to diverge from the behavior policy.
-However, this also introduces a heightened risk of taking unsafe actions.
-As illustrated in Fig.
-3 (c), the algorithm exhibits some performance fluctuations as $\beta$ increases from 0 to 0.1, though it generally maintains robustness.
-A low value of $\beta$ may lead to suboptimal solutions, whereas a high value of $\beta$ can induce excessive aggressiveness of the algorithm, resulting in performance instability.
-Therefore, $\beta$ also needs to be set appropriately to balance exploration and stability.
-Conclusion and Discussion
+The model bias adjustment coefficient $\alpha$ adjusts model-generated data, with larger values overly penalizing exploration and smaller values inducing over-optimism, both harming performance.
+To validate this, we fixed $\beta=0.01$ and varied $\alpha$ in the halfcheetah task.
+As shown in Fig.
+5 (b), increasing $\alpha$ from 0 to 0.05 initially improves performance but later hinders it due to excessive conservativeness, highlighting the need for a balanced $\alpha$ in policy learning.
+The policy shift coefficient $\beta$ regulates modifications to offline data, controlling deviation from the behavior policy.
+Larger $\beta$ values enhance exploration by encouraging greater deviation from the behavior policy, but also increase the risk of unsafe or suboptimal actions.
+As shown in Fig.
+5 (c), the algorithm exhibits performance fluctuations as $\beta$ increases from 0 to 0.1.
+Low $\beta$ may lead to suboptimal solutions, while high $\beta$ can induce excessive aggressiveness and instability.
+Therefore, $\beta$ also needs to be set appropriately.
+6 Conclusion and Discussion
 In this paper, we analyze the distribution shift problem, attributing its root causes to model bias and policy shift.
 Model bias results in distorted objectives, while policy shift obstructs the convergence of policies.
-We introduce a shifts-aware policy optimization framework to uniformly address these challenges and apply it within the model-based offline setting.
+We introduce a shifts-aware policy optimization framework to uniformly address these challenges through shift-aware rewards (SAR) and apply it within the model-based offline setting, thereby proposing our SAMBO-RL algorithm.
 Notably, our framework is versatile and can be integrated into general reinforcement learning algorithms.
 Our main limitation is classifier accuracy; imprecise classifiers struggle to estimate the shifts-aware reward accurately, leading to unstable training.
-In future work, We aim to explore alternative methods that directly address distribution shift using datasets, eliminating the need for classifier estimation.
+In future work, We aim to explore alternative methods that directly address distribution shifts using datasets, thereby eliminating the need for classifier estimation.
 Additionally, we will extend our shifts-aware reward framework to tackle distribution shift challenges in online settings, with a particular focus on policy shift problems.
-Our novel analysis and results related to policy shift in this context open new avenues for better exploiting offline datasets and overcoming the restrictions imposed by behavior policies.
-Appendix A A. Theoretical Analysis
-A1. The Supplement Analysis of Distribution Shift
+Our novel analysis and results related to policy shifts in this context open new avenues for better exploiting offline datasets and overcoming the implicit restrictions imposed by behavior policies.
+Impact Statement
+This study addresses the distribution shift problem in offline model-based reinforcement learning (RL), with the potential to facilitate the deployment of RL algorithms in real-world applications, particularly in scenarios where data collection is expensive.
+The research aims to enhance the robustness and reliability of RL systems in real world.
+However, the direct deployment of RL systems in high-risk domains, such as autonomous driving, may introduce substantial risks, especially if algorithms fail to adapt to unforeseen distribution shifts in real-world environments.
+Furthermore, the application of offline RL models, particularly in sensitive areas like healthcare, may raise significant ethical concerns.
+In such domains, the training data may not fully represent all patient groups.
+If distribution shift issues are not adequately addressed, the resulting treatment policies may fail for certain patient populations.
+Therefore, to ensure the safe, effective, and equitable deployment of RL systems in real world, rigorous testing, validation, and continuous monitoring are essential.
+Appendix A Theoretical Analysis
+A.1 The Supplement Analysis of Distribution Shift
 Dual Formulation of Model Bias.
@@
 Both of these impede the effective learning of the policy.
-A2. The Proof of Theorem 1
-Theorem 1.
+A.2 The Proof of Theorem 1
+Theorem A.1.
 Let $\mathcal{L}(\pi)$ be the training objective defined in (6) and $\tilde{r}$ denote the shifts-aware reward (5). Then,
@@
 $=(1-\gamma)\mathcal{L}(\pi).$
-A3. Classifiers Estimation Derivations
+A.3 Classifiers Estimation Derivations
 We assume that all steps in the derivation are appropriate, that the denominator is non-zero, and that the argument of the logarithm is not zero.
-Transition Classifier Derivation.
+A.3.1 Transition Classifier Derivation.
 Define two events $\mathbb{E}$ and $\mathbb{M}$, where
@@
 where $c_{1}$ is a constant that does not influence the training of SAMBO.
-Action Classifier Derivation.
+A.3.2 Action Classifier Derivation.
 The derivation of the action classifier is similar to that of the transition classifier. Define two events $\mathbb{A}$ and $\mathbb{C}$, where
@@
 ={Given the state $\mathbb{C}=\left\{\text{Given the state }s_{t},\text{ the agent takes action }a_{t}\text{ according to the behavior policy }\pi_{c}\right\}.$
-The action classifier is defined as $C_{\psi}(s_{t},a_{t})=P(\mathbb{A}|s_{t},a_{t})$, Then
-$C_{\psi}(s,a)=P(\mathbb{A}|s_{t},a_{t})=\frac{P(a_{t}|s_{t},\mathbb{A})P(\mathbb{A}|s_{t})P(s_{t})}{P(s_{t},a_{t})},$ (16)
-$1-C_{\psi}(s,a)=P(\mathbb{C}|s_{t},a_{t})=\frac{P(a_{t}|s_{t},\mathbb{C})P(\mathbb{C}|s_{t})P(s_{t})}{P(s_{t},a_{t})}.$ (17)
+The action classifier is defined as $C_{\psi_{1}}(s_{t},a_{t})=P(\mathbb{A}|s_{t},a_{t})$, Then
+$C_{\psi_{1}}(s,a)=P(\mathbb{A}|s_{t},a_{t})=\frac{P(a_{t}|s_{t},\mathbb{A})P(\mathbb{A}|s_{t})P(s_{t})}{P(s_{t},a_{t})},$ (16)
+$1-C_{\psi_{1}}(s,a)=P(\mathbb{C}|s_{t},a_{t})=\frac{P(a_{t}|s_{t},\mathbb{C})P(\mathbb{C}|s_{t})P(s_{t})}{P(s_{t},a_{t})}.$ (17)
 Specifically, $P(a_{t}|s_{t},\mathbb{A})=\pi(a_{t}|s_{t})$, $P(a_{t}|s_{t},\mathbb{C})=\pi_{c}(a_{t}|s_{t})$. Then according to (16) and (17), we have
-$\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}=\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}.$ (18)
-Similarly, $\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}$ can be also approximated by the constant $\frac{\left|\mathcal{D}_{\pi}\right|}{\left|\mathcal{D}_{env}\right|}$.
-According to the sampling method in Algorithm 2, we have
-$\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}=\frac{d(s_{t})\left|\mathcal{D}_{\pi}\right|}{d(s_{t})\left|\mathcal{D}_{env}\right|}=\frac{\left|\mathcal{D}_{\pi}\right|}{\left|\mathcal{D}_{env}\right|},$
-where $d(s_{t})$ denotes the distribution of states in the offline dataset, and $\mathcal{D}_{\pi}$ represents the dataset collected by policy $\pi$.
-Thus, taking the logarithm of both sides of (18), we get
-$\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}=\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}-c_{2},$
-where $c_{2}$ is a constant that does not influence the training process.
-Appendix B B. Implementation Details
+$\frac{C_{\psi_{1}}(s,a)}{1-C_{\psi_{1}}(s,a)}=\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}.$ (18)
+Similarly, $\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}$ can also be approximated by a classifier $C_{\psi_{2}}(s)$, i.e.
+$\frac{P(\mathbb{A}|s_{t})}{P(\mathbb{C}|s_{t})}=\frac{C_{\psi_{2}}(s)}{1-C_{\psi_{2}}(s)}$.
+Thus, we can approximate $\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}$ using two classifiers,
+$\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}=\log\frac{C_{\psi_{1}}(s,a)}{1-C_{\psi_{1}}(s,a)}-\log\frac{C_{\psi_{2}}(s)}{1-C_{\psi_{2}}(s)}.$
+For simplicity, we use the notation
+$\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}=\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$
+Appendix B Implementation Details
 We present a detailed outline of the SAMBO algorithm in Algorithm 1.
@@
 16:  end for
-B1. Model Training
+B.1 Model Training
 In our approach, the model is represented by a probabilistic neural network that outputs a Gaussian distribution for both the next state and reward, given the current state and action:
 $m_{\theta}\left(s_{t+1},r_{t}\mid s_{t},a_{t}\right)=\mathcal{N}\left(\mu_{\theta}\left(s_{t},a_{t}\right),\Sigma_{\theta}\left(s_{t},a_{t}\right)\right).$
-Our model training approach is consistent with the methodology used in prior works (Janner et al.
-2019; Yu et al.
-2020).
+Our model training approach is consistent with the methodology used in prior works (Janner et al., 2019; Yu et al., 2020).
 We train an ensemble of seven dynamics models and select the best five based on their validation prediction error from a held-out set containing 1000 transitions in the offline dataset $\mathcal{D}_{env}$.
@@
 During model rollouts, we randomly choose one model from the best five models.
-B2. Classifiers Training
+B.2 Classifiers Training
 The transition classifier $C_{\phi}(s,a,s^{\prime})\in[0,1]$ represents the probability that the transition originated from the environment.
@@
 The action classifier $C_{\psi}(s,a)\in[0,1]$ represents the probability that a given state-action pair $(s,a)$ was generated by the current policy $\pi$. This classifier is trained by minimizing:
-$\mathcal{L}_{\psi}=-\mathbb{E}_{\mathcal{D}_{\pi}}\left[\log C_{\psi}(s,a)\right]-\mathbb{E}_{\mathcal{D}_{env}}\left[\log(1-C_{\psi}(s,a))\right].$ (20)
+[…]
 For data in the model dataset $\mathcal{D}_{m}$, the dynamics are governed by the model $q$.
@@
 $\tilde{r}(s,a,s^{\prime})\approx\log r(s,a)+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$ (22)
-B3. Policy Learning
-The policy optimization in our approach is based on the SAC algorithm, with hyperparameter configurations aligning with those employed in MOBILE (Sun et al.
-2023).
+B.3 Policy Learning
+The policy optimization in our approach is based on the SAC algorithm, with hyperparameter configurations aligning with those employed in MOBILE (Sun et al., 2023).
 For each update, we draw a batch of 256 transitions, with $5\%$ sourced from the offline dataset $\mathcal{D}_{env}$ and the remaining $95\%$ from the synthetic dataset $\mathcal{D}_{m}$.
 The detailed hyperparameter settings used for the D4RL benchmark are provided in Table 3.
+Table 3: Hyperparameters for policy optimization in SAMBO
 Hyperparameters Value Description
@@
 $N_{\text{iter}}$ 3M Total gradient steps.
-Table 2: Hyperparameters for policy optimization in SAMBO
-Appendix C C. Experimental Details
-C1. Benchmark
-We conduct experiments on Gym tasks (“v2” version) within the D4RL benchmark and now provide the sources for the reported performance on this benchmark.
-For TD3+BC (Fujimoto and Gu 2021), EDAC (An et al.
-2021), TT (Janner, Li, and Levine 2021), RAMBO (Rigter, Lacerda, and Hawes 2022), and MOBILE (Sun et al.
-2023), we cite the scores reported in the original papers, as these evaluations were conducted on the “v2” datasets in Gym.
-Since COMBO (Yu et al.
-2021) does not provide source code, we include the results as reported in the original paper.
-The results of CQL (Kumar et al.
-2020) are taken from the performance table in MOBILE, where CQL was retrained on the “v2” datasets in Gym.
-The results of MOPO* are obtained from experiments using the OfflineRL-Kit libraryhttps://github.com/yihaosun1124/OfflineRL-Kit on the “v2” datasets.
+Appendix C Experimental Details
+C.1 Benchmark
+We conduct experiments on Gym tasks (“v2” version) within the D4RL (Fu et al., 2020) benchmark.
+We also evaluated SAMBO on the NeoRL (Qin et al., 2022) benchmark, a more challenging dataset generated using a more conservative policy to better reflect real-world data collection scenarios.
+These datasets are characterized by their narrow scope and limited coverage.
+D4RL.
+We now provide the sources for the reported performance on this benchmark.
+For TD3+BC (Fujimoto & Gu, 2021), EDAC (An et al., 2021), TT (Janner et al., 2021), RAMBO (Rigter et al., 2022), and MOBILE (Sun et al., 2023), we cite the scores reported in the original papers, as these evaluations were conducted on the “v2” datasets in Gym.
+Since COMBO (Yu et al., 2021) does not provide source code, we include the results as reported in the original paper.
+The results of CQL (Kumar et al., 2020) are taken from the performance table in MOBILE, where CQL was retrained on the “v2” datasets in Gym.
+The results of MOPO* are obtained from experiments using the OfflineRL-Kit library [^2] https://github.com/yihaosun1124/OfflineRL-Kit on the “v2” datasets.
 These scores are referenced in Table 1.
-C2. Hyperparameters
-We list the hyperparameters that have been tuned as follows. The specific settings are detailed in Table 3.
+NeoRL.
+We report the performance of BC, CQL, and MOPO based on the original NeoRL paper, while the performance of TD3+BC, EDAC, and MOBILE is derived from the scores reported in the MOBILE (Sun et al., 2023) paper.
+C.2 Hyperparameters
+We list the hyperparameters that have been tuned as follows. The specific settings are detailed in Table 4 and Table 5.
+Table 4: Hyperparameters of SAMBO used in the D4RL datasets.
 Task Name $\boldsymbol{\alpha}$ $\boldsymbol{\beta}$ $\boldsymbol{h}$ $\boldsymbol{c}$
@@
 hopper-random 0.02 0.01 5 0
-walker2d-random 0.03 0.03 5 0
-halfcheetah-medium 0.01 0.01 5 0
+walker2d-random 0.3 0.3 5 0
+halfcheetah-medium 0.03 0.03 5 0
 hopper-medium 0.02 0.01 5 0
@@
 halfcheetah-medium-replay 0.01 0.01 5 0
-hopper-medium-replay 0.02 0.01 5 0
-walker2d-medium-replay 0.01 0.01 1 -0.2
-halfcheetah-medium-expert 0.04 0.01 5 0
-hopper-medium-expert 0.04 0.02 5 0
-walker2d-medium-expert 0.05 0.05 1 0
-Table 3: Hyperparameters of SAMBO used in the D4RL datasets.
-Model Bias Coefficient $\alpha$.
-We tune $\alpha$ in the range of $[0.01,0.05]$.
-For training stability, in the walker2d-medium dataset, we fix the transition classifier after 1 million iterations by setting $\alpha=0$.
-In the walker2d-medium-expert dataset, we apply an annealing factor to the transition classifier.
-Initially set to 3, the factor remains constant for the first 1 million iterations and then decays reciprocally to 1 over the subsequent 2 million iterations.
-Policy Shift Coefficient $\beta$.
-We tune $\beta$ in the range of $[0.01,0.05]$.
+hopper-medium-replay 0.05 0.05 5 0
+walker2d-medium-replay 1 1 1 0
+halfcheetah-medium-expert 0.3 0.3 5 0
+hopper-medium-expert 1 1 5 0
+walker2d-medium-expert 2 2 1 0
+Table 5: Hyperparameters of SAMBO used in the NeoRL datasets.
+Task Name $\boldsymbol{\alpha}$ $\boldsymbol{\beta}$ $\boldsymbol{h}$ $\boldsymbol{c}$
+Halfcheetah-L 0.01 0.01 5 0
+Hopper-L 0.3 0.3 5 0
+Walker2d-L 3 1 1 0
+Halfcheetah-M 0.1 0.1 5 0
+Hopper-M 0.3 0.3 5 0
+Walker2d-M 3 1 1 0
+Halfcheetah-H 5 1 5 0
+Hopper-H 0.3 0.3 5 0
+Walker2d-H 3 1 1 0
+Coefficient $\alpha$ and $\beta$.
+The model bias coefficient, $\alpha$, determines the degree of correction for model bias, while the policy shift coefficient, $\beta$, regulates the extent of correction for policy shift.
 Rollout Length $h$.
@@
 In most datasets, we set $c=0$. However, we observe that employing a negative value for $c$ yields favorable results in certain datasets.
-C3. Computing Infrastructure
+We terminate training at 1M to prevent potential performance degradation due to the rapid convergence of the HalfCheetah environment in NeoRL.
+Additionally, to maintain training stability in the HalfCheetah-high dataset, we freeze classifiers after 0.5 million iterations.
+Similarly, in the walker2d-medium dataset of the D4RL benchmark, we ensure training stability by fixing the transition classifier after 1 million iterations.
+C.3 Reward Function Discussion
+Based on our derivation, if we start with $\mathbb{E}_{p^{\pi}(\tau)}e^{R(\tau)}$, we can obtain a new surrogate objective, which we refer to as SAMBO-r.
+In scenarios with sparse rewards and narrow datasets, where $\log r$ may distort the reward, SAMBO-r may be more suitable as a surrogate objective.
+  • SAMBO. The modified reward is
+    $\tilde{r}(s,a,s^{\prime})=\log r^{\prime}(s,a)+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})}+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$
+  • SAMBO-r. The modified reward is
+    $\tilde{r}(s,a,s^{\prime})=r^{\prime}(s,a)+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})}+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$
+C.4 Computing Infrastructure
 All experiments are conducted using a single GeForce GTX 3090 GPU and an Intel(R) Xeon(R) Gold 6330 CPU @ 2.00GHz.
@@
 The software libraries and frameworks utilized in SAMBO are consistent with those specified in this library.
+Appendix D Omitted Experiments
+D.1 Visualization
+Figure 6: The $(s,a,s^{\prime})$ distributions generated by SAMBO result in the walker2d-medium-expert task in D4RL benchmark.
+Model Data represents the distribution of data generated by executing the final learned policy in the dynamic model, while Real Data represents the distribution of data generated by executing the final learned policy in the environment.
+We also conducted a visualization experiment on the walker2d-medium-expert dataset within the D4RL benchmark.
+As shown in Figure 6, SAMBO effectively constrains the Model Data to closely match the Offline Data, and the Real Data generated by executing the learned policy in the actual environment does not significantly deviate from either the Offline Data or Model Data.
+This highlights the effectiveness of SAMBO in mitigating distribution shift across both benchmarks.
+D.2 Detailed Ablation Discussion
+SAMBO is regulated by the model bias adjustment coefficient $\alpha$ and the policy shift modification coefficient $\beta$.
+The model bias adjustment coefficient $\alpha$ controls the degree of adjustment to model-generated data.
+A larger $\alpha$ leads to excessive conservatism, restricting exploration even in well-performing regions and potentially causing suboptimal performance.
+Conversely, a smaller $\alpha$ may encourage excessive optimism in exploring risky regions, impairing performance.
+To validate this, we set $\beta=0.01$ and varied $\alpha$ in the Halfcheetah task.
+The results, shown in Fig.
+7 (a), indicate that both very low ($\alpha=0$) and high ($\alpha=0.02$ or $0.05$) values lead to decreased performance.
+Specifically, as $\alpha$ increases from 0 to 0.05, the algorithm initially improves but later declines as it becomes increasingly conservative, which impedes exploration.
+This underscores that both excessively low and high values of $\alpha$ are detrimental to policy learning.
+The policy shift modification coefficient $\beta$ governs the extent of modification applied to the offline data, thereby influencing the degree of deviation from the behavior policy.
+A larger $\beta$ value promotes increased exploratory behavior by encouraging the algorithm to diverge from the behavior policy.
+However, this also introduces a heightened risk of taking unsafe actions.
+As shown in Fig.
+7 (b), the algorithm exhibits some performance fluctuations as $\beta$ increases from 0 to 0.1, though it generally maintains robustness.
+A low value of $\beta$ may lead to suboptimal solutions, whereas a high value of $\beta$ can induce excessive aggressiveness of the algorithm, resulting in performance instability.
+Thus, $\beta$ should also be set appropriately to balance exploration and stability.
+(a) Tuning for $\alpha$
+(b) Tuning for $\beta$
+Figure 7: Illustration of hyperparameter experiments in the HalfCheetah-random task in D4RL benchmark.
+D.3 Performance Compare
+We also compared the performance curves of model-based approaches on other tasks in the NeoRL benchmark.
+The curve for MOBILE was obtained by executing the official implementation.
+Since MOPO does not have an official implementation on NeoRL, we report its score based on the results presented in the original NeoRL paper.
+As shown in Figures 10 to 16, SAMBO outperforms in all Hopper and Walker2d tasks, both of which are more complex and challenging.
+Although it exhibits slower convergence initially in the Walker2d-M and Walker2d-H tasks, its performance improves rapidly as the experiment progresses, ultimately achieving state-of-the-art (SOTA) levels.
+On the other hand, SAMBO also demonstrates comparable performance in simpler HalfCheetah tasks.
+Figure 8: Hopper-L
+Figure 9: Hopper-M
+Figure 10: Hopper-H
+Figure 11: Walker2d-L
+Figure 12: Walker2d-M
+Figure 13: Walker2d-H
+Figure 14: HalfCheetah-L
+Figure 15: HalfCheetah-M
+Figure 16: HalfCheetah-H

### v3

@@
-SAMBO-RL: Shifts-aware Model-based Offline Reinforcement Learning
-Abstract
+Mitigating Distribution Shift in Model-based Offline RL via Shifts-aware Reward Learning
+Abstract.
 Model-based offline reinforcement learning trains policies using pre-collected datasets and learned environment models, eliminating the need for direct real-world environment interaction.
@@
 Our theoretical and empirical investigations reveal how these factors distort value estimation and restrict policy optimization.
-To tackle these challenges, we derive a novel Shifts-aware Reward (SAR) through a unified probabilistic inference framework, which modifies the vanilla reward to refine value learning and facilitate policy training.
-Building on this, we introduce Shifts-aware Model-based Offline Reinforcement Learning (SAMBO-RL), a practical framework that efficiently trains classifiers to approximate SAR for policy optimization.
-Empirical experiments show that SAR effectively mitigates DS, and SAMBO-RL achieves superior or comparable performance across various benchmarks, underscoring its effectiveness and validating our theoretical analysis.
-Machine Learning, ICML
-1 Introduction
-Offline reinforcement learning (RL) (Lange et al., 2012; Levine et al., 2020) learns policies from the offline dataset generated by a behavior policy, avoiding additional online interaction with the environment.
-This approach shows remarkable potential in data-driven decision scenarios (Emerson et al., 2023; Sinha et al., 2022) where exploration is costly or hazardous.
-The model-based framework (Luo et al., 2018) is particularly effective for offline RL (Yu et al., 2020, 2021; Sun et al., 2023; Li et al., 2024).
+To tackle these challenges, we derive a novel shifts-aware reward through a unified probabilistic inference framework, which modifies the vanilla reward to refine value learning and facilitate policy training.
+Building on this, we develop a practical implementation that leverages classifier-based techniques to approximate the adjusted reward for effective policy optimization.
+Empirical results across multiple benchmarks demonstrate that the proposed approach mitigates distribution shift and achieves superior or comparable performance, validating our theoretical insights.
+Key words and phrases:
+Offline RL, Distribution Shift, Model-based RL
+1. Introduction
+Offline reinforcement learning (RL) Lange et al.
+(2012); Levine et al.
+(2020) learns policies from the offline dataset generated by a behavior policy, avoiding additional online interaction with the environment.
+This approach shows remarkable potential in data-driven decision scenarios Emerson et al.
+(2023); Sinha et al.
+(2022) where exploration is costly or hazardous.
+The model-based framework Luo et al.
+(2018) is particularly effective for offline RL Yu et al.
+(2020, 2021); Sun et al.
+(2023); Li et al.
+(2024).
 It involves learning models of the environment from the offline dataset and utilizing these models to generate data for policy training.
 However, directly learning a policy using offline and synthetic data introduces the distribution shift challenge.
-This shift causes the training objective to deviate from the true objective of RL, leading to poor test performance.
+This shift causes the training objective to significantly deviate from the true objective of reinforcement learning, leading to poor performance during testing.
 Prior methods in model-based offline RL focus on directly leveraging off-policy methods in online RL and heuristic model uncertainty for conservative learning to mitigate the impact of distribution shift.
-For instance, MOPO (Yu et al., 2020) uses the predicted variance of the learned model as an uncertainty penalty on rewards.
-MOBILE (Sun et al., 2023) penalizes value learning of model data according to the uncertainty of the Bellman Q-function estimated by ensemble models.
-RAMBO (Rigter et al., 2022) considers the trajectory distribution in the worst case and adversarially trains the policy and transition model while ensuring accurate predictions.
-However, these approaches have resulted in a biased training objective due to the model uncertainty and remain the distribution shift problem to be solved.
-Figure 1: The negative impacts of model bias and policy shift.
-(a) In the 1D grid world, the agent can take two actions: move left (L) or move right (R) from the current state $S$ to the adjacent grid cells $S-1$ or $S+1$, with movement obstructed at the boundaries.
+For instance, MOPO Yu et al.
+(2020) uses the predicted variance of the learned model as an uncertainty penalty on rewards.
+MOBILE Sun et al.
+(2023) penalizes value learning of model data according to the uncertainty of the Bellman Q-function estimated by ensemble models.
+RAMBO Rigter et al.
+(2022) considers the trajectory distribution in the worst case and adversarially trains the policy and model while ensuring accurate predictions.
+However, these approaches result in a biased objective by only addressing the distribution shift from model inaccuracies, leaving the broader distribution shift unresolved.
+Figure 1.
+The negative impacts of model bias and policy shift.
+(a) In the 1D grid world, the agent can take two actions: move left (L) or move right (R) from the current state $S$ to the adjacent grid cells S1S-1 or $S+1$, with movement obstructed at the boundaries.
 The initial state is uniformly distributed, and the agent receives undiscounted rewards upon reaching specific states.
@@
 Even slight model bias can significantly distort value function estimates, leading the learned policies to deviate from the true optimal policy in the actual environment.
-(c) The right panel showcases expected returns and the KL divergence between the learned policy $\pi$ and the behavior policy $\pi_{b}$ during training with and without policy shift.
-We find that policy shift implicitly restricts $\pi$ to remain close to $\pi_{b}$, thereby slowing policy convergence.
+(c) The right panel showcases expected returns and the KL divergence between the learned policy and the behavior policy with and without policy shift.
+We find that policy shift implicitly slows policy convergence.
 In this paper, we demystify the distribution shift problem in model-based RL and study its fundamental properties.
@@
 This bias leads to inaccurate estimations of trajectory returns, resulting in distorted objectives.
-If the learned model overestimates or underestimates the probability of generating a trajectory, it consequently causes the overestimation or underestimation of trajectory returns.
-Additionally, policy shift depicts the changes in state-action visitation distribution caused by the deviation between the learned policy $\pi$ and behavior policy $\pi_{b}$.
-This shift implicitly forces $\pi$ towards $\pi_{b}$, hindering the learning process and slowing down policy convergence.
-In contrast, previous studies heuristically mitigate these two issues from off-policy algorithms and model uncertainty.
-To address these challenges, we theoretically derive adjustments for model bias and policy shift by employing a unified probabilistic inference framework, and introduce a shifts-aware reward, which integrates the vanilla reward with adjustments for model bias and policy shift.
-The model bias adjustment penalizes the reward to mitigate issues of overestimation or underestimation, thereby refining value learning and preventing its convergence to local optima.
-Meanwhile, the policy shift modification revises the reward to motivate $\pi$ to escape the limitation of $\pi_{b}$, accelerating policy convergence.
-Building upon our theoretical framework, we develop Shifts-aware Model-based Offline Reinforcement Learning (SAMBO-RL).
-This approach incorporates model bias adjustment for model-generated data and policy shift modification for offline data.
+If the learned model overestimates or underestimates the probability of generating a trajectory, the resulting distribution shift during training consequently leads to overestimation or underestimation of trajectory returns.
+Additionally, policy shift depicts the changes in state-action visitation distribution caused by the deviation between the training data collection policy and the learned policy.
+This distribution shift between training and deployment implicitly perturbs policy updates and thereby hinders learning.
+Existing studies heuristically mitigate these issues with off-policy mechanisms and uncertainty penalties, lacking a consistent, theoretically grounded framework.
+To address these challenges, we theoretically derive adjustments for model bias and policy shift by employing a unified probabilistic inference framework, and introduce a shifts-aware reward (SAR), which integrates the vanilla reward with adjustments for model bias and policy shift.
+The model bias adjustment penalizes the reward to mitigate issues of overestimation or underestimation, thereby refining value learning and alleviating distribution shift during training.
+Meanwhile, the policy shift modification adjusts the reward to mitigate distribution shift between training and deployment, stabilizing the policy updates and accelerating policy convergence.
+Building on our theoretical framework, we develop Shifts-aware Model-based Offline Reinforcement Learning (SAMBO-RL).
+This approach incorporates model bias adjustment for model-generated data and policy shift modification for training data.
 Furthermore, it learns a transition classifier and action classifiers to approximate the shifts-aware reward and then performs policy optimization based on this adjusted reward.
 Our contributions include:
-  • We demystify the distribution shift problem in model-based offline RL, i.e., model bias distorts the estimation of the value function, while policy shift imposes an implicit policy limitation that hinders policy convergence.
-  • We upgrade the vanilla reward into a shifts-aware reward incorporating model bias adjustment and policy shift modification.
-This modified reward is used for policy learning to mitigate distribution shift in model-based offline RL.
-  • We devise SAMBO-RL, which trains transition and action classifiers to approximate the shifts-aware reward, effectively addressing both model bias and policy shift.
-  • We show that shifts-aware reward effectively mitigates distribution shift, and SAMBO-RL demonstrates superior or comparable performance across various benchmarks.
-2 Related Work
-Model-free Offline RL encompasses two primary approaches: policy constraints and value regularization, both incorporating conservatism (Jin et al., 2021).
-Policy constraint methods aim to keep the learned policy close to the behavioral policy, mitigating distribution shift.
-For instance, BEAR (Kumar et al., 2019) and BRAC (Wu et al., 2019) constrain the learned policy by minimizing different divergences.
-BCQ (Fujimoto et al., 2019) perturbs actions during learning to ensure they remain within the action space of the offline dataset by learning a generative model of the behavior policy.
-On the other hand, value regularization methods add conservative regularization terms to the value optimization objective, avoiding out-of-distribution (OOD) actions.
-CQL (Kumar et al., 2020) penalizes the Q-values of OOD samples by incorporating regularization into the value loss function.
-TD3+BC (Fujimoto & Gu, 2021) introduces a behavior cloning regularization into the TD3 (Fujimoto et al., 2018) objective.
-Similarly, EDAC (An et al., 2021) and PBRL (Bai et al., 2021) penalize the Q-function based on uncertainty derived from an ensemble of Q-networks.
-MCQ (Lyu et al., 2022) actively trains OOD actions by constructing pseudo-target values to alleviate over-pessimism.
-Model-based Offline RL approximates the environment using learned models and performs conservative policy optimization (Lu et al., 2021), achieving high data efficiency (Li et al., 2024).
-MOPO (Yu et al., 2020) incorporates conservatism by penalizing rewards based on the uncertainty of model predictions.
-MOReL (Kidambi et al., 2020) introduces a pessimistic MDP to penalize rewards of state-action pairs in unexplored regions.
-COMBO (Yu et al., 2021) extends CQL to the model-based setting, regularizing the value function on OOD samples.
-TT (Janner et al., 2021) uses a transformer to model offline trajectories and employs beam search for planning.
-RAMBO (Rigter et al., 2022) adopts an adversarial approach, training policy and model within a robust framework while ensuring accurate transition predictions of the model.
-CBOP (Jeong et al., 2022) adaptively weights multi-step returns in a model-based value expansion framework (Feinberg et al., 2018).
-Lastly, MOBILE (Sun et al., 2023) introduces penalties during value learning by leveraging the uncertainty of the Bellman Q-function estimates derived from ensemble learned dynamic models.
-3 Preliminaries
+  • We demystify the distribution shift in model-based offline RL, i.e., model bias causes distribution inconsistency in the training data, ultimately distorting value estimation, while policy shift leads to distribution inconsistency between training and deployment, hindering policy convergence.
+  • We upgrade the vanilla reward into a shifts-aware reward that incorporates model bias adjustment and policy shift modification, effectively mitigating distribution shift.
+  • We devise SAMBO-RL, which trains a transition classifier and an action classifier to approximate the shifts-aware reward, effectively addressing both model bias and policy shift.
+  • We empirically show that shifts-aware reward effectively mitigates distribution shift, and SAMBO-RL demonstrates comparable performance across various benchmarks.
+2. Preliminaries
 MDP.
-We focus on a Markov decision process (MDP) specified by the tuple $M=(\mathcal{S},\mathcal{A},p,r,\mu_{0},\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $p(s^{\prime}|s,a)$ is the environment transition dynamics, positive $r(s,a)>0$ is the reward function, $\mu_{0}(s)$ is the initial state distribution, and $\gamma\in(0,1)$ is the discount factor.
-Given an MDP, we define the state value function $V^{\pi}(s)=\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right]$ and the $Q$ function $Q^{\pi}(s,a)=\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right]$.
+We focus on a Markov decision process (MDP) specified by the six-elements tuple $M=(\mathcal{S},\mathcal{A},p,r,\mu_{0},\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $p(s^{\prime}|s,a)$ is the environment transition dynamics, positive $r(s,a)>0$ is the reward function, $\mu_{0}(s)$ is the initial state distribution, and $\gamma\in(0,1)$ is the discount factor typically used for future reward weighting.
+Given an MDP, we define the state value function of the policy $\pi$: $V^{\pi}(s)=\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right]$ and the $Q$ function of the policy $\pi$: $Q^{\pi}(s,a)=\mathbb{E}_{\pi,p}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a\right]$.
 MBPO.
-Similar to previous works (Yu et al., 2020, 2021; Rigter et al., 2022; Sun et al., 2023), we focus on model-based policy optimization (MBPO) (Janner et al., 2019), which employs an actor-critic RL algorithm.
-MBPO first performs $k$-step rollouts using the model $m(s^{\prime}|s,a)$ starting from state $s\in\mathcal{D}_{env}$ and adds the generated data to a replay buffer $\mathcal{D}_{m}$.
-Then it optimizes the policy using mini-batches of data sampled from $\mathcal{D}_{env}\cup\mathcal{D}_{m}$, where each datapoint is sampled from the real dataset $\mathcal{D}_{env}$ with probability $f$ and the model dataset $\mathcal{D}_{m}$ with probability $1-f$.
-Trajectory Formulation.
-To demystify the distribution shift, We consider the trajectory formulation of optimization objectives.
-Let $p^{\pi}(\cdot)$ represent the probability distribution over trajectories =$\mathcal{H}=\{\tau=(s_{0},a_{0},r_{0},s_{1},...)\}$ generated by executing policy $\pi$ under dynamics $p$.
-This is given by:
+Similar to previous works Yu et al.
+(2020, 2021); Rigter et al.
+(2022); Sun et al.
+(2023) in offline RL, we focus on model-based policy optimization (MBPO) Janner et al.
+(2019), which uses an actor-critic RL algorithm.
+Define $\mathcal{D}_{env}$ as the offline dataset collected by behavior policy $\pi_{b}$ in the real environment, and model dataset $\mathcal{D}_{m}$ as the buffer storing synthetic data.
+MBPO first performs $h$-step rollouts using the model $m(s^{\prime}|s,a)$ starting from state $s\in\mathcal{D}_{env}$ and adds the generated data to the buffer $\mathcal{D}_{m}$.
+It then optimizes the policy using data sampled from $\mathcal{D}_{env}\cup\mathcal{D}_{m}$, where each datapoint is drawn from $\mathcal{D}_{env}$ with probability $f$ and $\mathcal{D}_{m}$ with probability 1f1-f.
+We denote the data used for training as $\mathcal{D}_{train}$.
+[…]
 $p^{\pi}(\tau)=\mu_{0}(s_{0})\prod\limits_{t=0}^{\infty}p(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t}).$ (1)
-The training data $\mathcal{D}_{train}$ can be interpreted as being generated by a data collection policy $\pi_{c}$ through rollouts in dynamics $q$.
-Consequently, the probability of observing trajectory $\tau$ in the training data can be formulated as $q^{\pi_{c}}(\tau)$.
-The reinforcement learning objective and the practical training objective are then formalized as the following equations:
+The training data $\mathcal{D}_{train}$ can be interpreted as generated by a data collection policy $\pi_{c}$ through rollouts in an unknown dynamics $q$.
+$\pi_{c}$ is inherently related to $\pi$, as the dataset $\mathcal{D}_{train}$ includes samples generated by the policy $\pi$.
+Consequently, the probability of observing trajectory $\tau$ in the training data is formulated as $q^{\pi_{c}}(\tau)$.
+Then the RL objective and the practical training objective are
 $\mathcal{J}_{M}(\pi)=\mathbb{E}_{\tau\sim p^{\pi}}\left[R(\tau)\right]\text{and }\mathcal{J}(\pi)=\mathbb{E}_{\tau\sim q^{\pi_{c}}}[R(\tau)],$ (2)
-where $R(\tau)=\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})$ represents the discounted return of the whole trajectory $\tau=(s_{0},a_{0},r_{0},s_{1},...)$.
-4 Shifts-aware Policy Optimization
+[…]
+3. Related Work
+Model-free Offline RL encompasses two primary approaches: policy constraints and value regularization, both incorporating conservatism Jin et al.
+(2021).
+Policy constraints keep the learned policy close to the behavior policy, while value regularization adds conservative terms to the value optimization.
+BEAR Kumar et al.
+(2019) and BRAC Wu et al.
+(2019) constrain the learned policy by minimizing different divergences.
+BCQ Fujimoto et al.
+(2019) perturbs actions during learning to ensure they remain within the action space of the offline dataset by learning a generative model of the behavior policy.
+SACS Mao et al.
+(2024a) simultaneously performs out-of-distribution (OOD) state correction using a model that is not employed for data generation and OOD action suppression in offline RL.
+CQL Kumar et al.
+(2020) penalizes the Q-values of OOD samples by incorporating regularization into the value loss function.
+TD3+BC Fujimoto and Gu (2021) introduces a behavior cloning regularization into the TD3 Fujimoto et al.
+(2018) objective.
+Similarly, EDAC An et al.
+(2021) and MCQ Lyu et al.
+(2022b) actively trains OOD actions by constructing pseudo-target values to alleviate over-pessimism.
+DMG Mao et al.
+(2024b) leverages mild action generalization and generalization propagation to enhance generalization in offline RL.
+SEABO assigns larger rewards to transitions closer to their nearest expert demonstrations, alleviating offline RL’s heavy reliance on the offline data annotated with reward labels.
+A2PR Liu et al.
+(2024) selects high-advantage actions as an imitation learning regularizer to mitigate performance degradation from excessive conservatism.
+Model-based Offline RL approximates the environment using learned models and performs conservative policy optimization Lu et al.
+(2021), achieving high data efficiency Li et al.
+(2024).
+MOPO Yu et al.
+(2020) incorporates conservatism by penalizing rewards based on the uncertainty of model predictions.
+MOReL Kidambi et al.
+(2020) introduces a pessimistic MDP to penalize rewards of state-action pairs in unexplored regions.
+COMBO Yu et al.
+(2021) extends CQL to the model-based setting, regularizing the value function on OOD samples.
+RAMBO Rigter et al.
+(2022) adopts an adversarial approach, training the policy and model within a robust framework while ensuring accurate transition predictions of the model.
+CBOP Jeong et al.
+(2022) adaptively weights multi-step returns in a model-based value expansion framework Feinberg et al.
+(2018).
+CABI Lyu et al.
+(2022a) improves the reliability of synthesized data by training bidirectional dynamics models and trusting only the samples where the forward and backward models are consistent.
+C-LAP Alles et al.
+(2024) learns a generative model of the joint distribution of observations and actions, and imposes an implicit constraint on the generated actions to enable constrained policy learning.
+Lastly, MOBILE Sun et al.
+(2023) introduces penalties during value learning by leveraging the uncertainty of the Bellman Q-function estimates derived from ensemble models, achieving SOTA performance among these model-based and model-free approaches.
+Cross-domain policy transfer aims to transfer a policy learned in a source environment to a target environment with different dynamics.
+This is similar to model-based approaches, which learn a model to approximate the environment dynamics and derive high-performing policies in the real environment by training within the learned model.
+GARAT Desai et al.
+(2020) learns an adversarial imitation-from-observation policy by discriminating between generated actions and target environment actions.
+DARC Eysenbach et al.
+(2020) addresses cross-domain RL via reward correction to enable domain adaptation under changing environment dynamics.
+SAIL Liu et al.
+(2019) performs state alignment in cross-domain imitation learning to align state sequences across different domains.
+Similarly, SRPO Xue et al.
+(2023) introduces a regularization term based on the KL divergence between state distributions of different domains in the RL objective to improve data efficiency.
+OTDF Lyu et al.
+(2025) uses cross-domain offline datasets and mitigates dynamics mismatches via optimal transport, facilitating policy learning.
+4. Shifts-aware Policy Optimization
 We first formalize the distribution shift problem as shift weighting and further delve into its negative effects on policy optimization from theory and experiments.
 To address these challenges, we employ probabilistic inference to derive a surrogate objective and formulate a shifts-aware reward.
-Finally, we implement this theoretical framework within model-based offline RL and introduce SAMBO-RL, a practical algorithm designed to mitigate distribution shifts.
-4.1 Demystifying Distribution Shift
+Then we implement this theoretical framework within model-based offline RL and introduce SAMBO-RL, a practical algorithm to mitigate distribution shifts.
+Finally, we provide a theoretical analysis of SAR.
+4.1. Demystifying Distribution Shift
 We examine the distribution shift problem from an optimization perspective and identify its origins as model bias and policy shift.
-Then we explore how these factors affect accuracy in value estimation and impose implicit policy restrictions, resulting in a biased objective and slow convergence.
-The supplementary analysis is in Appendix A.1.
-Shift Weighting. The training objective $\mathcal{J}(\pi)$ in (2) can be expressed as the following trajectory formulation:
+Then we explore how these factors affect accuracy in value estimation and induce instability in policy updates.
+Shift Weighting. The training objective $\mathcal{J}(\pi)$ in (2) can be expressed as the trajectory formulation:
 $\mathcal{J}(\pi)=\mathbb{E}_{\tau\sim p^{\pi}(\tau)}\left[\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}R(\tau)\right].$ (3)
-Compared to (2), it indicates that for a given trajectory $\tau$, the true objective under practical sampling distributions differs in shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$ from the training objective, which formalizes the distribution shift in trajectory formulation.
-Substituting (1) into the shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$, we can attribute this weighting to two primary components: model bias and policy shift, reformulated as the following:
+Compared to $\mathcal{J}_{M}(\pi)$ in (2), it indicates that for a given trajectory $\tau$, the true objective under practical sampling distributions differs in shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$ from the training objective, which formalizes the distribution shift in trajectory formulation.
+Substituting (1) into the shift weighting $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}$, we can attribute it to two primary components: model bias and policy shift.
+Then the shift weighting can be reformulated as the following:
 $\frac{q^{\pi_{c}}(\tau)}{p^{\pi}(\tau)}=\underbrace{\prod_{t}\frac{q(s_{t+1}|s_{t},a_{t})}{p(s_{t+1}|s_{t},a_{t})}}_{\text{model bias}}\underbrace{\prod_{t}\frac{\pi_{c}(a_{t}|s_{t})}{\pi(a_{t}|s_{t})}}_{\text{policy shift}}.$ (4)
 Model bias illustrates the discrepancy between the dynamics used for training and those of the actual environment, while policy shift represents the divergence between the policies used for sampling training data and those being learned.
+Both components ultimately lead to the distribution shift issue.
 Negative Impact of Shift Weighting. We analyze the impact of shift weighting from model bias and policy shift.
-(i) To illustrate the effect of model bias, we examine the scenario where the training policy $\pi$ operates within model dynamics $q$, which diverges from the actual environment dynamics $p$.
-In this context, the training objective $\mathcal{J}(\pi)$ is:
-$\mathcal{J}(\pi)$ $=\mathbb{E}_{\tau\sim p^{\pi}}\left[R(\tau)\prod\limits_{t=0}^{\infty}\frac{q(s_{t+1}|s_{t},a_{t})}{p(s_{t+1}|s_{t},a_{t})}\right].$
-Compared to the true objective (3), the compounded prediction bias of the model significantly affects value estimation and leads to a distorted objective.
-Specifically, when the model $q$ overestimates or underestimates the probability of generating trajectory $\tau$, it consequently overestimates or underestimates the expected return of this trajectory in the real environment.
-This essentially introduces bias for the optimization objective, resulting in wrong training directions.
-(ii) To delve deeper into the policy optimization process under policy shift, we consider the setting where a sampling policy $\pi_{c}$ is executed within the true environment transition dynamics $p$.
-In this scenario, the training objective $\mathcal{J}(\pi)$ is:
-$\mathcal{J}(\pi)$ $=\mathbb{E}_{(s,a)\sim d^{p,\pi}}\left[r(s,a)\frac{d^{p,\pi_{c}}(s,a)}{d^{p,\pi}(s,a)}\right].$
-Here, $d^{p,\pi}(s,a)$ is the state-action visitation distribution under dynamics $p$ and policy $\pi$, i.e.
-$d^{p,\pi}(s,a)=(1-\gamma)\mathbb{E}_{s_{0}\sim\mu_{0},s_{t}\sim p,a_{t}\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{I}(s_{t}=s,a_{t}=a)\right]$, where $\mathbb{I}$ represents the indicator function.
+(i) To illustrate the effect of model bias, we examine the synthetic data generation scenario in model-based approaches, where the training policy $\pi$ operates within model dynamics $m$ (i.e.
+$q=m$), which diverges from the environment dynamics $p$.
+In this context, according to (4), the training objective $\mathcal{J}(\pi)$ is:
+$\mathcal{J}(\pi)$ $=\mathbb{E}_{\tau\sim p^{\pi}}\left[R(\tau)\prod\limits_{t=0}^{\infty}\frac{m(s_{t+1}|s_{t},a_{t})}{p(s_{t+1}|s_{t},a_{t})}\right].$
+Compared to the true objective, the compounded prediction bias of the model significantly affects value estimation and leads to a distorted objective.
+Specifically, when the model $m$ overestimates or underestimates the probability of generating trajectory $\tau$, it consequently overestimates or underestimates the expected return of this trajectory in the real environment.
+The inconsistency between the model and the environment in predicting future dynamics leads to a mismatch in trajectory distributions during training.
+This distribution shift in training introduces bias into the optimization objective, thereby guiding the policy toward suboptimal or incorrect updates.
+(ii) To delve deeper into the policy optimization process under policy shift, We consider a typical offline reinforcement learning scenario, in which policy optimization is constrained to the training dataset $\mathcal{D}_{train}$, and assume that the learned model perfectly fits the environment dynamics.
+In this scenario, the training objective $\mathcal{J}(\pi)$ can be formulated as:
+$\mathcal{J}(\pi)$ $=\mathbb{E}_{(s,a)\sim d^{p,\pi}}\left[r(s,a)(d^{p,\pi_{c}}(s,a)/d^{p,\pi}(s,a))\right].$
+Here, $d^{p,\pi}(s,a)$ is the state-action visitation distribution under dynamics $p$ and policy $\pi$
+$d^{p,\pi}(s,a)=(1-\gamma)\mathbb{E}_{s_{0}\sim\mu_{0},s_{t}\sim p,a_{t}\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{I}(s_{t}=s,a_{t}=a)\right],$
+where $\mathbb{I}$ represents the indicator function.
 Note that $\pi_{c}$ is related to $\pi$.
-Compared to the true objective, the policy shift implicitly forces policy $\pi$ to approach $\pi_{c}$, hindering the learning process and slowing down policy convergence.
+Compared to the true objective, the policy shift implicitly forces policy $\pi$ to approach $\pi_{c}$, perturbing policy updates.
 Specifically, for a state-action pair $(s,a)$ associated with a high reward, ideal policy updates should increase $\pi(a|s)$.
@@
 Similarly, for the state-action pair $(s,a)$ with a low reward, if the currently $\pi$ is reasonably good and leads to $d^{p,\pi}(s,a)<d^{p,\pi_{c}}(s,a)$, the reward can be overestimated, potentially resulting in policy updates to increase $\pi(a|s)$ that do not align with optimal policy directions.
+The mismatch between the deployed policy and the training distribution can disrupt updates.
 Toy Examples.
-To empirically validate our analysis of model bias and policy shift, we use a 1D grid world example, as depicted in Fig.
+To empirically validate our analysis of model bias and policy shift, we use a 1D grid world example depicted in Fig.
 1 (a).
@@
 Consequently, the policies learned by the agents deviate from the optimal policy.
-To assess the impact of the policy shift, we examine a model-free offline setting where data is pre-collected using a behavior policy $\pi_{b}$.
+To assess the impact of the policy shift, we examine a model-free offline setting where offline data is pre-collected using a behavior policy $\pi_{b}$.
 Fig.
-1 (c) showcases that the KL divergence between the learned policy $\pi$ and $\pi_{b}$ under policy shift is substantially reduced compared to the scenario without policy shift, sometimes even decreasing undesirably.
-This indicates that policy shift does compel $\pi$ to align closely with $\pi_{b}$, thereby impeding the training process and substantially slowing down the policy convergence.
-4.2 Shifts-aware Reward via Probabilistic Inference
+1 (c) showcases that the KL divergence between the learned policy $\pi$ and behavior policy $\pi_{b}$ under policy shift is substantially reduced compared to the scenario without policy shift, sometimes even decreasing undesirably.
+This indicates that the policy shift is impeding the training process and substantially slowing down the policy convergence.
+4.2. Shifts-aware Reward via Probabilistic Inference
 This subsection introduces a novel shifts-aware reward through a unified probabilistic inference framework, serving as a lower-bound surrogate for the true objective.
-We examine how this modified reward contributes to refining the value function and motivating $\pi$ to escape the limitations of behavior policy, thus mitigating model bias and policy shift.
+We examine how this modified reward contributes to refining the value function and constraining unstable policy updates, thereby mitigating the distribution shift in both training and real-world deployment.
 Shifts-aware Reward.
 We introduce the shifts-aware reward $\tilde{r}(s_{t},a_{t},s_{t+1})$ formulated within trajectories.
-It comprises the vanilla reward, adjustments for model bias, and modifications for policy shift, expressed as the following:
-$\begin{split}&\tilde{r}(s_{t},a_{t},s_{t+1})=\log r(s_{t},a_{t})\\ &+\frac{1}{(1-\gamma)\gamma^{t}}\left[\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right].\end{split}$ (5)
-This modified reward forms the basis of our surrogate objective function during training, which is formulated as:
+It comprises the vanilla reward, adjustments for model bias and modifications for policy shift, expressed as the following equation:
+$\begin{split}\tilde{r}(s_{t},a_{t},s_{t+1})&=\log r(s_{t},a_{t})\\ &+\frac{1}{(1-\gamma)\gamma^{t}}\left[\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right].\end{split}$ (5)
+This modified reward forms the basis of our surrogate objective function during training, which is defined as:
 $\mathcal{L}(\pi)=\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\sum_{t=0}^{\infty}\gamma^{t}\tilde{r}(s_{t},a_{t},s_{t+1})\right].$ (6)
-Additionally, we establish that $\mathcal{L}(\pi)$ serves as a lower bound for the true objective, ensuring its rationality as a surrogate.
-Theorem 4.1.
+Additionally, we establish that the surrogate objective $\mathcal{L}(\pi)$ serves as a lower bound for the true objective.
+Theorem. 1.
 Let $\mathcal{L}(\pi)$ be the training objective defined in (6) and $\tilde{r}$ denote the shifts-aware reward (5). Then,
 $\log\left[\mathcal{J}_{M}(\pi)\right]\geq(1-\gamma)\mathcal{L}(\pi),\quad\forall\ \pi.$
-Optimization objective based on this shifts-aware reward provides a methodological foundation for integrating shift adjustments seamlessly into reinforcement learning algorithms to uniformly address model bias and policy shift.
-Modal Bias Adjustment.
+The theorem ensures that $\mathcal{L}(\pi)$ is a rational surrogate objective for $\mathcal{J}_{M}(\pi)$.
+Optimization objective based on this shifts-aware reward provides a methodological foundation for integrating shift adjustments seamlessly to uniformly address distribution shift caused by model bias and policy shift.
+Model Bias Adjustment.
 We now examine how this adjustment refines the value estimation and addresses the challenge of model bias.
-By rolling out the learned policy $\pi$ within the model dynamics $m$, we obtain a model-generated dataset $\mathcal{D}_{m}$.
-Furthermore, the modified training objective of the model dataset $\mathcal{D}_{m}$ can be computed as the following:
+By rolling out the learned policy $\pi$ in model $m$, we obtain a model-generated dataset $\mathcal{D}_{m}$.
+Then, when training on dataset $\mathcal{D}_{train}$, the modified objective restricted to $\mathcal{D}_{m}$ is computed as:
 $\tilde{J}(\pi)$ $=\mathbb{E}_{s\sim\mathcal{D}_{m},a\sim\pi(\cdot|s),s^{\prime}\sim\mathcal{D}_{m}}\left[\tilde{r}(s,a,s^{\prime})\right]$
@@
 This modified reward consists of the logarithmic vanilla reward penalized by the KL divergence between the model and environment dynamics.
-This adjustment reduces value function misestimations by accounting for the accuracy of the learned model, encouraging the development of both a high-return policy and an accurate model approximation.
+This adjustment reduces value function misestimation by accounting for the accuracy of the learned model.
+By emphasizing data from regions with reliable model predictions, it mitigates the distribution shift in training data caused by model bias.
 Policy Shift Modification.
-To delve into the positive effect of policy shift modification, we consider a dataset $\mathcal{D}_{b}$ generated by executing the behavior policy $\pi_{b}$ within the environment $p$.
-Moreover, the modified training objective of the dataset $\mathcal{D}_{b}$ is derived as the following formulation:
-$\tilde{J}(\pi)=\mathbb{E}_{s\sim\mathcal{D}_{b},a\sim\pi(\cdot|s),s^{\prime}\sim\mathcal{D}_{b}}\left[\tilde{r}(s,a,s^{\prime})\right]$
-$=\mathbb{E}_{s\sim\mathcal{D}_{b}}\left[\mathbb{E}_{a\sim\pi(\cdot|s)}\left[\log r(s,a)\right]+\beta D_{\operatorname{KL}}(\pi(\cdot|s)||\pi_{b}(\cdot|s))\right],$
-which introduces a bonus to the vanilla reward based on the KL divergence between $\pi$ and $\pi_{b}$.
-This modification guides $\pi$ to break out of the confines of the behavior policy $\pi_{b}$, promoting better exploration and more effective learning.
-4.3 Shifts-aware Model-based Offline RL
-We apply the shifts-aware reward in model-based offline RL and develop the practical framework, SAMBO-RL.
+To delve into the positive effect of policy shift modification, we consider the modified training objective in the training dataset $\mathcal{D}_{train}$ restricted to $\mathcal{D}_{env}$, in which case the dynamic $q$ satisfies $q=p$ and the data collection policy satisfies $\pi_{c}=\pi_{b}$ in equation (4).
+Then the modified objective is
+$\tilde{J}(\pi)$ $=\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}_{env}}\left[\tilde{r}(s,a,s^{\prime})\right]$
+$=\mathbb{E}_{s\sim\mathcal{D}_{env}}\left[\mathbb{E}_{a\sim\pi_{b}(\cdot|s)}\left[\log r(s,a)\right]-\beta D_{\operatorname{KL}}(\pi_{b}(\cdot|s)||\pi(\cdot|s))\right].$
+By introducing a penalty to the vanilla reward based on the KL divergence between offline data collection policy $\pi_{b}$ and testing policy $\pi$, this modification prevents $\pi$ from making overly aggressive updates, thus stabilizing the learning process.
+It reduces the bias between the offline data and the testing policy, helping to mitigate the distribution shift caused by policy shift.
+4.3. Shifts-aware Model-based Offline RL
+We apply the shifts-aware reward in model-based offline RL and develop the algorithm SAMBO-RL.
 To adapt the shifts-aware reward for practical implementation, where training data is typically stored in transition format rather than trajectories, we introduce two weighting factors, $\alpha$ and $\beta$.
-These factors replace the respective coefficients in the reward function, shaping it as the following:
-$\tilde{r}(s,a,s^{\prime})=\log r(s,a)+\alpha\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}+\beta\log\frac{\pi(a|s)}{\pi_{c}(a|s)}.$
+These factors replace respective coefficients in reward function (5):
+$\tilde{r}(s,a,s^{\prime})=\log r(s,a)+\alpha\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}+\beta\log\frac{\pi(a|s)}{\pi_{c}(a|s)}.$ (7)
 These weighting factors ensure the reward and adjustments are properly balanced in transition-based learning scenarios, which is empirically validated in the following experiments.
-Furthermore, since the environment dynamics $p$ and behavior policy $\pi_{b}$ are typically unknown and cannot be exactly computed, we derive two classifiers to estimate the shifts-aware reward (Eysenbach et al., 2020, 2022).
+Moreover, as discussed in Section 4.2, for data in $\mathcal{D}_{env}$, the dynamics satisfy $q=p$ and the policy satisfies $\pi_{c}=\pi_{b}$.
+For data in $\mathcal{D}_{m}$, we have $q=m$, and the data-collection policy can be approximated by $\pi$, i.e., $\pi_{c}=\pi$.
+As discussed in detail in Appendix B.2, we have
+$\tilde{r}(s,a,s^{\prime})=\begin{cases}\log r(s,a)+\beta\log\frac{\pi(a|s)}{\pi_{b}(a|s)},&\forall(s,a,s^{\prime})\in\mathcal{D}_{env}\\ \log r(s,a)+\alpha\log\frac{p(s^{\prime}|s,a)}{m(s^{\prime}|s,a)},&\forall(s,a,s^{\prime})\in\mathcal{D}_{m}\end{cases}$
+Furthermore, since the environment dynamics $p$ and the policy $\pi_{b}$ are typically unknown and cannot be exactly computed, we derive classifiers to estimate the shifts-aware reward Eysenbach et al.
+(2020, 2022); Liu et al.
+(2021).
 Details are shown in Appendix A.2.
-Specifically, these classifiers estimate model bias adjustment $\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}$ and policy shift modification $\log\frac{\pi(a|s)}{\pi_{c}(a|s)}$ respectively.
+Specifically, these classifiers estimate $\log\frac{p(s^{\prime}|s,a)}{m(s^{\prime}|s,a)}$ and $\log\frac{\pi(a|s)}{\pi_{b}(a|s)}$ respectively.
 The transition classifier $C_{\phi}(s,a,s^{\prime})$ estimates the probability that the transition $(s,a,s^{\prime})$ comes from the environment, while the action classifier $C_{\psi}(s,a)$ estimate the probability that a given state-action pair $(s,a)$ is generated by the current policy $\pi$.
-We train the classifiers by minimizing the loss functions:
+We train these classifiers by minimizing the cross-entropy loss:
 $\mathcal{L}(\varphi)=-\mathbb{E}_{\mathcal{D}_{1}}[\log C_{\varphi}]-\mathbb{E}_{\mathcal{D}_{2}}[\log(1-C_{\varphi})].$ (8)
-For the transition classifier $C_{\varphi}=C_{\phi}(s,a,s^{\prime})$, $\mathcal{D}_{1}=\mathcal{D}_{env}$ is the offline dataset, and $\mathcal{D}_{2}=\mathcal{D}_{m}$ is the model-generated dataset.
-For the action classifier $C_{\varphi}=C_{\psi}(s,a)$, $\mathcal{D}_{1}=\mathcal{D}_{\pi}$ is the dataset collected by executing the current policy $\pi$, and $\mathcal{D}_{2}=\mathcal{D}_{env}$ is the offline dataset.
-Based on these classifiers, we approximate the shifts-aware reward as the following:
-$\tilde{r}(s,a,s^{\prime})\approx\log r(s,a)$ (8)
-$+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})}+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$
-The complete procedure of SAMBO-RL is outlined in Algorithm 1, with further details provided in Appendix B.
-5 Experiments
-In our experiments, we focus on three objectives: (1) validating the effectiveness of shifts-aware reward, (2) comparing the performance of SAMBO with previous methods, and (3) analyzing the contribution of each component and hyperparameter within SAMBO to its overall performance.
+For the transition classifier $C_{\varphi}=C_{\phi}(s,a,s^{\prime})$, $\mathcal{D}_{1}=\mathcal{D}_{env}$ and $\mathcal{D}_{2}=\mathcal{D}_{m}$.
+For the action classifier $C_{\varphi}=C_{\psi}(s,a)$, $\mathcal{D}_{1}=\mathcal{D}_{\pi}$ is the dataset collected by executing the current policy $\pi$, and $\mathcal{D}_{2}=\mathcal{D}_{env}$.
+Based on these classifiers, we approximate the shift-aware reward by
+$\tilde{r}(s,a,s^{\prime})=\begin{cases}\log r(s,a)+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)},&\forall(s,a,s^{\prime})\in\mathcal{D}_{env}\\ \log r(s,a)+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})},&\forall(s,a,s^{\prime})\in\mathcal{D}_{m}\end{cases}$ (9)
+The procedure of SAMBO-RL is outlined in Algorithm 1, with implementation details provided in Appendix B.
+Algorithm 1 SAMBO-RL
+1:Offline dataset $\mathcal{D}_{env}$, learned dyanmics models $\{m_{\theta}^{i}\}_{i=1}^{N}$, initialized policy $\pi$, hyperparameters $\alpha$ and $\beta$.
+2:Train the probabilistic dynamics model $m_{\theta}(s^{\prime},r|s,a)=\mathcal{N}(\mu_{\theta}(s,a),\Sigma_{\theta}(s,a))$ on $\mathcal{D}_{env}$.
+3:Initialize the model dataset buffer $\mathcal{D}_{m}\leftarrow\varnothing$.
+4:for $i=1,2,...,N_{iter}$ do
+5:  Initialize the policy dataset buffer $\mathcal{D}_{\pi}\leftarrow\varnothing$.
+6:  Generate synthetic rollouts by model $m_{\theta}$. Add transition data in these rollouts to $\mathcal{D}_{m}$ and $\mathcal{D}_{\pi}$.
+7:  Update classifiers $C_{\phi}$ and $C_{\psi}$ according to Eq. (8).
+8:  Compute shifts-aware reward according to Eq. (9).
+9:  Run SAC Haarnoja et al. (2018) with shifts-aware reward to update policy $\pi$.
+10:end for
+4.4. Theoretical Analysis
+We analyze the theoretical properties of SAR, with detailed proofs provided in Appendix A. We first introduce $\xi$-uncertainty quantifier, which characterizes the suboptimality of the derived policy.
+Definition 2.
+($\xi$-Uncertainty Quantifier Jin et al. (2021)). The set of penalization $\{\Gamma_{h}\}_{h\in[H]}$ forms a $\xi$-uncertainty quantifier if it holds with probability at least $1-\xi$ that
+$\left|\widehat{\mathbb{B}}\widehat{V}_{t+1}(s,a)-\mathbb{B}\widehat{V}_{t+1}(s,a)\right|\leq\Gamma_{t}(s,a),\quad\forall(s,a)\in\mathcal{S}\times\mathcal{A}$
+where $\widehat{V}_{t+1}$ is an estimated value function at step $t+1$ and $\widehat{\mathbb{B}}$ is an empirical Bellman operator estimating the Bellman operator $\mathbb{B}$.
+We can reformulate (7) in a state-action style:
+$\tilde{r}(s,a)$ $=\mathbb{E}_{s^{\prime}\sim q(\cdot|s)}\left[\tilde{r}(s,a,s^{\prime})\right]$
+$=\log r(s,a)-\alpha D_{\operatorname{KL}}(q(\cdot|s,a)\|p(\cdot|s,a))-\beta\log\frac{\pi_{c}(a|s)}{\pi(a|s)}.$
+Since the model cannot perfectly fit the environment dynamics $p$, we assume that the KL divergence satisfies $D_{\operatorname{KL}}(m(\cdot\mid s,a)\,\|\,p(\cdot\mid s,a))\geq\epsilon_{KL}>0$.
+Under this KL divergence assumption, we prove that the penalty of SAR $\tilde{r}(s,a)$ forms a $\xi$-uncertainty quantifier.
+Theorem. 3.
+Under the assumption that $D_{\operatorname{KL}}(m(\cdot\mid s,a)\,\|\,p(\cdot\mid s,a))\geq\epsilon_{KL}>0$, there exist $\alpha>0$ and $\beta>0$ such that the penalty of SAR forms a $\xi$-uncertainty quantifier.
+Based on this theorem, we derive the following value bound from the perspective of pessimistic value iteration (PEVI) Jin et al. (2021), which demonstrates that SAR is an effective corrective reward.
+Theorem. 4.
+Let $\pi^{*}$ be the optimal policy in the environment, $\pi^{*}_{G}$ be the optimal policy in MDP $G=(\mathcal{S},\mathcal{A},p,\log r,\mu_{0},\gamma)$, $V_{G}^{\pi}(s)$ the value function in $G$ and $\epsilon_{s}\geq D_{\operatorname{KL}}(m(\cdot\mid s,a)\,\|\,p(\cdot\mid s,a))\geq\epsilon_{KL}>0$.
+Then the derived policy $\widehat{\pi}$ according to SAR satisfies
+$\left|V^{\pi^{*}}(s_{1})-V^{\widehat{\pi}}(s_{1})\right|\leq 2\alpha H\epsilon_{s}+\epsilon(G;s_{1}),$
+with probability at least $1-\xi$ for all $s\in\mathcal{S}$. Here,
+$\epsilon(G;s)=|V^{\widehat{\pi}}(s)+V_{G}^{\pi^{*}_{G}}(s)-V^{\pi^{*}}(s)-V_{G}^{\widehat{\pi}}(s)|,$
+which quantify the value error between different MDPs and $\epsilon(M;s)=0,\forall s\in\mathcal{S}$.
+5. Experiments
+In our experiments, we focus on three objectives: (1) validating the effectiveness of shifts-aware reward, (2) comparing the performance of SAMBO with previous model-free and model-based methods, and (3) analyzing the contribution of each component in SAMBO.
 We evaluate our algorithm using standard offline RL benchmarks in Mujoco environments and conduct an ablation study to assess its performance.
-Our implementation is based on the OfflineRL-Kit library [^1] https://github.com/yihaosun1124/OfflineRL-Kit, a comprehensive and high-performance library for implementing offline reinforcement learning algorithms.
+Our implementation is based on the OfflineRL-Kit library [^1] https://github.com/yihaosun1124/OfflineRL-Kit, a comprehensive and high-performance library for implementing offline RL algorithms.
 The basic parameters of our algorithm are consistent with the settings in this library.
-5.1 Effectiveness of Shifts-aware Reward
-(a) Model Bias Adjustment
-(b) Policy Shift Modification
-Figure 2: Assessment of the shifts-aware reward.
-(a) Model bias adjustment mitigates inaccuracies in value estimation, achieving a nearly optimal return.
-(b) Policy shift modification addresses implicit policy restrictions, leading to faster convergence.
-We conduct an evaluation of the effectiveness of the shifts-aware Reward in addressing model bias and policy shift issues within the 1D grid world illustrated in Fig. 1 (a).
-Table 1: Results on the D4RL Gym benchmark.
-MOPO’s results are obtained on the “v0” datasets; MOPO*’s results are obtained from experiments in the OfflineRL-Kit library on the “v2” datasets.
+Table 1.
+Results on the D4RL Gym benchmark.
+MOPO are obtained on the “v0” datasets; MOPO* are obtained from experiments in the OfflineRL-Kit library on the “v2” datasets.
 The numbers reported for SAMBO are the normalized scores averaged over the final iteration of training across 4 seeds, with $\pm$ standard deviation.
 The top three scores are highlighted in bold.
-We use an asterisk “*” for the best score, an underscore “_” for the second, and bold text for the third.
-Task Name BC CQL TD3+BC EDAC MOPO MOPO* COMBO TT RAMBO MOBILE SAMBO (Ours)
-halfcheetah-random 2.2 31.3 11.0 28.4 35.4 38.5 38.8 6.1 39.5 39.3 39.7$\pm$2.0
-hopper-random 3.7 5.3 8.5 25.3 11.7 31.7 17.9 6.9 25.4 31.9 32.4$\pm$0.5
+Task Name BC CQL TD3+BC EDAC MOPO MOPO* COMBO $T$ RAMBO MOBILE(SOTA) SAMBO (Ours)
+[…]
+[…]
 walker2d-random 1.3 5.4 1.6 16.6 13.6 7.4 7.0 5.9 0.0 17.9 8.1 $\pm$ 7.8
-halfcheetah-medium 43.2 46.9 48.3 65.9 42.3 72.4 54.2 46.9 77.9 74.6 72.5$\pm$3.8
-hopper-medium 54.1 61.9 59.3 101.6 28.0 62.8 97.2 67.1 87.0 106.6 99.7$\pm$0.3
-walker2d-medium 70.9 79.5 83.7 92.5 17.8 84.1 81.9 81.3 84.9 87.7 78.3$\pm$2.6
-halfcheetah-medium-replay 37.6 45.3 44.6 61.3 53.1 72.1 55.1 44.1 68.7 71.7 70.6$\pm$1.6
-hopper-medium-replay 16.6 86.3 60.9 110.0 67.5 92.7 89.5 99.4 99.5 103.9 101.5$\pm$1.5
-walker2d-medium-replay 20.3 76.8 81.8 87.1 39.0 85.9 56.0 82.6 89.2 89.9 92.3$\pm$ 1.5
-halfcheetah-medium-expert 44.0 95.0 90.0 106.3 63.3 83.6 90.0 95.0 95.4 108.2 97.8 $\pm$0.3
-hopper-medium-expert 53.9 96.9 98.0 110.7 23.7 74.6 111.0 110.0 88.2 112.6 111.0 $\pm$1.2
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+halfcheetah-medium-expert 44.0 95.0 90.7 106.3 63.3 83.6 90.0 95.0 95.4 108.2 97.8 $\pm$0.3
+hopper-medium-expert 53.9 96.9 98.0 110.7 23.7 74.6 111.1 110.0 88.2 112.6 111.0 $\pm$1.2
 walker2d-medium-expert 90.1 109.1 110.1 114.7 44.6 108.2 103.3 101.9 56.7 115.2 106.1 $\pm$ 2.9
-Model Bias Adjustment.
-Model bias exacerbates inaccuracies in value estimates, ultimately resulting in a distorted optimization objective.
-To evaluate the effects of model bias adjustment on the training process, we conduct experiments.
+[…]
+(a) SAR in Toy Example
+(b) MOBILE
+(c) SAMBO
+Figure 2.
+Effectiveness of the shifts-aware reward.
+(a) SAR is capable of achieving a near-optimal return.
+(b) and (c) The $(s,a,s^{\prime})$ distributions of offline RL algorithms result in the Hopper-M task in NeoRL benchmark.
+Model Data represents the distribution of data generated by executing the final learned policy in the dynamic model, while Real Data represents the distribution of data generated by executing the final learned policy in the real environment.
+5.1. Effectiveness of Shifts-aware Reward
+Effectiveness in Toy Example.
+To assess the efficacy of shifts-aware reward in different models, we examine its amendment to the training process.
 As illustrated in Fig.
-2 (a), we execute the policy gradient algorithm under four settings: using the vanilla reward with the overestimating model (OM) or underestimating model (UM), the shifts-aware reward with the overestimating model (SAR with OM) underestimating model (SAR with UM).
-We then evaluate the expected return of the learned policy after each training update in the actual environment.
+2 (a), we execute the policy gradient algorithm under four settings: using the vanilla reward with the overestimating model (OM) or underestimating model (UM), the SAR with the overestimating model (SAR with OM) or underestimating model (SAR with UM).
+We then evaluate the expected return of the learned policy after each training update in the real environment.
 The results show that while the OM initially achieves optimal returns, it rapidly diverges from the optimal policy.
@@
 Similarly, SAR with UM mitigates the policy suboptimality due to model underestimation, achieving a return near the theoretical optimum.
-(a) MOBILE
-(b) SAMBO
-Figure 3: The $(s,a,s^{\prime})$ distributions of offline RL algorithms result in the Hopper-M task.
-Model Data represents the distribution of data generated by executing the final learned policy in the dynamic model, while Real Data represents the distribution of data generated by executing the final learned policy in the environment.
-Policy Shift Modification.
-Policy shift implicitly restricts policy updates.
-To assess the efficacy of the policy shift modification, we apply the policy gradient algorithm to the offline objective $\mathcal{J}(\pi)=\mathbb{E}_{s\sim d^{p,\pi_{b}}}\left[V^{\pi}(s)\right]$ in the environment depicted in Fig.
-1 (a).
-We test four settings: the vanilla reward or SAR, combined with a behavior policy that is either uniformly distributed or the opposite of the optimal policy.
-We measure the KL divergence between the learned policy and the behavior policy after each update.
-As illustrated in Fig.
-2 (b), the algorithm using the vanilla reward is limited by the behavior policy, serving a lower KL divergence and slower convergence.
-In contrast, the algorithm employing the SAR effectively motivates policy to be optimistic.
-It overcomes the restriction, displaying a higher and more stable divergence and achieving faster convergence.
+Both demonstrate the effectiveness.
 Effectiveness in Complex Tasks.
@@
 Fig.
-3 shows a visual comparison in the NeoRL task (details in Sec.
+2 (c) and (d) shows a visual comparison in the NeoRL task (details in Sec.
 5.2.2).
-SAMBO effectively constrains model-generated data (i.e., Model Data) to align closely with the Offline Data, whereas the previous state-of-the-art method, MOBILE, imposes weaker constraints.
-Furthermore, the real-world data (i.e., Real Data) generated by SAMBO is well-aligned with the Offline Data and remains tightly concentrated, indicating the SAMBO’s effectiveness.
-In contrast, the Real Data from MOBILE deviates significantly from both the Offline and Model Data, still suffering from distribution shift, leading to diminished performance.
-We attribute this discrepancy to SAMBO’s robust handling of model bias and policy shift, ensuring that the Real Data, Model Data, and Offline Data remain closely aligned, effectively mitigating the distribution shift.
-5.2 Benchmark Results
-5.2.1 D4RL
-Datasets.
-In line with prior research, SAMBO is evaluated using the D4RL benchmark (Fu et al., 2020) on the MuJoCo simulator (Todorov et al., 2012).
-The evaluation encompasses 12 datasets spanning three environments (halfcheetah, hopper, walker2d) and four dataset types (random, medium, medium-replay, medium-expert).
-The “v2” versions of datasets are employed for assessment.
-Baselines.
-We evaluate SAMBO against several baseline methods, including model-free methods such as BC (Bain & Sammut, 1995; Ross et al., 2011), which learns the behavior policy, CQL (Kumar et al., 2020), TD3+BC (Fujimoto & Gu, 2021), and EDAC (An et al., 2021), as well as model-based methods such as MOPO (Yu et al., 2020), COMBO (Yu et al., 2021), TT (Janner et al., 2021), RAMBO (Rigter et al., 2022), and MOBILE (Sun et al., 2023).
-Table 2: Normalized average returns in NeoRL benchmark over 4 random seeds.
-The results for MOBILE are reported from the original paper, and MOBILE represents the results obtained by executing the official implementation.
-Task Name BC CQL TD3+BC EDAC MOPO MOBILE MOBILE SAMBO (Ours)
-HalfCheetah-L 29.1 38.2 30.0 31.3 40.1 54.7 54.8 39.8 $\pm$ 2.6
-Hopper-L 15.1 16.0 15.8 18.3 6.2 17.4 18.2 19.4 $\pm$ 1.3
-Walker2d-L 28.5 44.7 43.0 40.2 11.6 37.6 23.7 46.1 $\pm$ 5.7
-HalfCheetah-M 49.0 54.6 52.3 54.9 62.3 77.8 79.0 62.5 $\pm$1.6
-Hopper-M 51.3 64.5 70.3 44.9 1.0 51.1 43.4 67.6 $\pm$ 12.2
-Walker2d-M 48.7 57.3 58.6 57.6 39.9 62.2 60.1 63.4 $\pm$ 4.0
-HalfCheetah-H 71.3 77.4 75.3 81.4 65.9 83.0 71.8 71.6 $\pm$ 13.9
-Hopper-H 43.1 76.6 75.3 52.5 11.5 87.8 42.3 93.9 $\pm$ 15.2
-Walker2d-H 72.6 75.3 69.6 75.5 18.0 74.9 71.9 79.2 $\pm$ 0.6
+SAMBO effectively constrains model-generated data (i.e., Model Data) to align closely with the Offline Data by model bias adjustment, whereas the previous state-of-the-art (SOTA) method, MOBILE, imposes weaker constraints.
+This enables SAMBO to effectively address the distribution shift between offline data and model data during training, thereby allowing for more efficient and reliable model utilization.
+Benefiting from the policy shift modification, the real-world data (i.e., Real Data) generated by SAMBO is well-aligned with both the Offline Data and the Model Data.
+This enables SAMBO to align with the deployment distribution, effectively addressing the distribution shift between the training data and the data encountered during real deployment, facilitating an efficient connection from training to testing.
+These ultimately result in an effective resolution to the distribution shift.
+Table 2. Normalized average returns in NeoRL benchmark over 4 random seeds. The hightest scores are highlighted in bold.
+Task Name BC CQL TD3+BC EDAC MOPO MOBILE SAMBO (Ours)
+HalfCheetah-L 29.1 38.2 30.0 31.3 40.1 54.8 38.4 $\pm$ 3.2
+Hopper-L 15.1 16.0 15.8 18.3 6.2 18.2 19.4 $\pm$ 1.3
+Walker2d-L 28.5 44.7 43.0 40.2 11.6 23.7 46.1 $\pm$ 5.7
+HalfCheetah-M 49.0 54.6 52.3 54.9 62.3 79.0 62.5 $\pm$1.6
+Hopper-M 51.3 64.5 70.3 44.9 1.0 43.4 67.6 $\pm$ 12.2
+Walker2d-M 48.7 57.3 58.6 57.6 39.9 60.1 63.4 $\pm$ 4.0
+HalfCheetah-H 71.3 77.4 75.3 81.4 65.9 71.8 71.6 $\pm$ 13.9
+Hopper-H 43.1 76.6 75.3 52.5 11.5 42.3 93.9 $\pm$ 15.2
+Walker2d-H 72.6 75.3 69.6 75.5 18.0 71.9 79.2 $\pm$ 0.6
+Average 45.4 56.1 54.5 50.7 28.5 51.7 60.2 $\pm$ 6.4
 (a) Hopper-L
@@
 (c) Hopper-H
-Figure 4: Performance comparison of Hopper tasks in NeoRL benchmark.
+Figure 3. Performance comparison of Hopper tasks in NeoRL benchmark.
+5.2. Benchmark Results
+5.2.1. D4RL
+Datasets.
+In line with prior research, SAMBO is evaluated using the D4RL benchmark Fu et al.
+(2020) on the MuJoCo simulator Todorov et al.
+(2012).
+The evaluation encompasses 12 datasets spanning three environments (halfcheetah, hopper, walker2d) and four dataset types (random, medium, medium-replay, medium-expert).
+To ensure consistency, we employ the “v2” versions of datasets for standardized evaluation.
+Baselines.
+We evaluate SAMBO against several baseline methods, including model-free methods such as BC Bain and Sammut (1995); Ross et al.
+(2011) which learns the behavior policy, CQL Kumar et al.
+(2020) that penalizes Q-values on OOD samples, TD3+BC Fujimoto and Gu (2021) that adopts a BC constraint during optimizing policy, and EDAC An et al.
+(2021) that penalizes the Q-function based on uncertainty derived from ensemble networks, as well as model-based methods such as MOPO Yu et al.
+(2020) that penalize rewards based on the uncertainty of the model predictions, COMBO Yu et al.
+(2021) that extends the penalty of CQL in the model-based setting to regularize OOD samples, TT Janner et al.
+(2021) that employs a Transformer to model distributions and uses beam search to plan, RAMBO Rigter et al.
+(2022) that trains the policy and model within a robust framework, and the prior SOTA algorithm MOBILE Sun et al.
+(2023) that penalizes value function during training by the uncertainty of Bellman Q-function estimates derived from ensemble models.
 Comparison Results.
 Results are presented in Table 1.
-Each number represents the normalized score, calculated as 100×100$\times$100 ×(score -- random policy score)/ (expert policy score -- random policy score) (Fu et al., 2020).
+Each number represents the normalized score, calculated as $100\times$(score - random policy score)/ (expert policy score - random policy score) Fu et al.
+(2020).
 SAMBO achieves strong performance on random datasets and demonstrates competitive performance across medium, medium-replay, and medium-expert datasets.
-This performance aligns with our theoretical analysis.
-By addressing both model bias and policy shift, SAMBO effectively balances the exploitation of offline data with the need to overcome the limitations of the behavioral policy.
-This approach enables the algorithm to be more aggressive and optimistic in its search for the optimal policy.
-In scenarios involving low-quality datasets, the aggressive exploration of SAMBO is more likely to yield benefits, potentially leading to the discovery of a superior policy.
-Conversely, a conservative approach may impede the learning of optimal policy in such contexts.
-For high-quality datasets, the advantages of aggressive exploration tend to diminish.
-Fortunately, in these scenarios, methods that emphasize estimating Q-value uncertainty perform effectively.
-This suggests that our approach and conservative estimation methods can perform well on complementary datasets.
-5.2.2 NEORL
+Furthermore, SAMBO demonstrates highly competitive average performance, approaching that of the model-free ensemble method EDAC An et al.
+(2021), and second only to the current SOTA model-based ensemble method MOBILE Sun et al.
+(2023).
+However, unlike these methods, SAMBO achieves this competitive performance without relying on value ensemble techniques.
+(a) Walker2d-L
+(b) Walker2d-M
+(c) Walker2d-H
+Figure 4. Illustration of ablation experiments in NeoRL benchmark.
+5.2.2. NEORL
 Datasets.
-NeoRL (Qin et al., 2022) is a benchmark designed to emulate real-world scenarios by generating datasets through a more conservative policy, aligning closely with real-world data-collection scenarios.
+NeoRL Qin et al.
+(2022) is a benchmark carefully designed to emulate real-world scenarios by generating datasets through a more conservative policy, thereby closely aligning with real-world data-collection scenarios.
 The datasets, characterized by their narrow scope and limited coverage, present a substantial challenge for offline RL algorithms.
 In this work, we examine nine datasets across three environments (HalfCheetah-v3, Hopper-v3, Walker2d-v3) and three dataset quality levels (L, M, H), corresponding to low, medium, and high-quality datasets.
-NeoRL offers different numbers of trajectories for each task (100, 1000, and 10000), and we consistently selected 1000 trajectories in our test.
+NeoRL provides varying numbers of trajectories for each task (100, 1000, and 10000), and for consistency, we uniformly selected 1000 trajectories in all our experiments.
 Baselines.
-We compare SAMBO with baseline methods that align with the d4rl experiments in Section 5.2.1, excluding TT, COMBO, and RAMBO, as no results for these methods are reported in their original papers or the NeoRL paper, and determining suitable hyperparameters for them would require an excessive amount of computational time.
+We compare SAMBO with baseline methods that align with the D4RL experiments in Section 5.2.1, excluding TT, COMBO, and RAMBO, as no results for these methods are reported in their original papers or the NeoRL paper, and determining suitable hyperparameters for them would be computationally expensive.
 Comparison Results.
@@
 Since MOPO does not have an official implementation on NeoRL, we report its score based on the original NeoRL paper.
-The experiments show that Our algorithm achieves superior or competitive performance on the more challenging and realistic NeoRL datasets, which can be seen as lower-quality datasets compared to the D4RL datasets.
-In particular, SAMBO demonstrates exceptional performance in the Walker2d and Hopper environments, which, in contrast to the HalfCheetah environment, exhibit higher control complexity, more intricate action dimensions, and greater training difficulty.
-These offline datasets more accurately capture the dynamic control challenges encountered in real-world scenarios, making them significantly demanding and highlighting that SAMBO has strong potential for real-world applications.
+Compared to other model-based and model-free methods, including ensemble approaches, SAMBO achieves superior performance on the more challenging and realistic NeoRL benchmarks.
+Moreover, SAMBO consistently outperforms Behavior Cloning (BC), while most baselines show degraded performance compared to BC on certain tasks.
+In particular, SAMBO demonstrates exceptional performance improvements in the Walker2d and Hopper environments.
+These results highlight that SAMBO has strong potential for applications on challenging datasets.
 More experiments and discussions are provided in Appendix D.2.
-(a) Ablation in halfcheetah-random dataset
-(b) Tuning for $\alpha$
-(c) Tuning for $\beta$
-(d) Ablation study in Walker2d-H dataset
-Figure 5: Illustration of hyperparameter experiments in D4RL benchmark and NeoRL benchmark.
-(a) and (d) The effectiveness of model bias adjustment and policy shift modification in the shifts-aware reward.
-(b) A large model bias adjustment coefficient $\alpha$ increases conservativeness.
-(c) A large policy shift modification coefficient $\beta$ increases the aggressiveness of exploration.
-5.3 Ablation Studies
-We conducted ablation studies on our method to understand why SAMBO demonstrates competitive performance. The following experimental results are presented in Figure 5.
-Ablation.
+5.3. Ablation and Tuning Studies
 Our shifts-aware reward integrates both the model bias adjustment and the policy shift modification.
-To elucidate the importance of these components, we perform ablation studies on the Halfcheetah-random task in D4RL benchmark and Walker2d-H task in NeoRL benchmark.
+To elucidate the importance of these components, we perform ablation studies on the Walker2d task in NeoRL benchmark.
 As shown in Figure.
-5 (a) and (d), we assess the performance under four configurations: SAMBO, SAMBO without model bias adjustment (wo/mb), SAMBO without policy shift modification (wo/ps), and SAMBO without any modifications (logr), while keeping all other parameters consistent.
-Figure.
-5 (a) show that, compared to the experiment denoted as logr, the performance of wo/mb deteriorates.
-This degradation arises because wo/mb employs policy shift modification to break the implicit limitations of the behavior policy and encourage more aggressive exploration.
-However, it does not address model bias, resulting in significant inaccuracies in value estimation, ultimately leading to suboptimal performance.
-This negative effect is particularly evident in the narrow Walker2d-H dataset, as seen in Figure 5 (d).
+4, we assess the performance under four configurations: SAMBO, SAMBO without model bias adjustment (wo/mb), SAMBO without policy shift modification (wo/ps), while keeping all other parameters consistent.
+The ablations show that, The wo/mb employs policy shift modification to stabilize policy updates.
+However, it does not address model bias and leads to significant inaccuracies in value estimation, ultimately resulting in poor performance.
 Conversely, the wo/ps ablation addresses model bias but neglects policy shift adjustments.
-While this reduces value misestimation, the algorithm remains constrained by the behavior policy due to the lack of policy shift modification.
-Consequently, wo/ps is less effective in discovering a more optimal policy.
-As shown in Figure 5 (a), the performance of wo/ps is inferior to sambo, and Figure 5 (d) illustrates that wo/ps converges slowly compared to sambo.
-In contrast, SAMBO incorporates not only model bias adjustment but also policy shift modification, facilitating aggressive exploration while addressing value estimation challenges, thus alleviating suboptimal performance.
-Hyperparameter Tuning.
-SAMBO is primarily regulated by two key hyperparameters: the model bias adjustment coefficient $\alpha$ and the policy shift modification coefficient $\beta$.
-The model bias adjustment coefficient $\alpha$ adjusts model-generated data, with larger values overly penalizing exploration and smaller values inducing over-optimism, both harming performance.
-To validate this, we fixed $\beta=0.01$ and varied $\alpha$ in the halfcheetah task.
-As shown in Fig.
-5 (b), increasing $\alpha$ from 0 to 0.05 initially improves performance but later hinders it due to excessive conservativeness, highlighting the need for a balanced $\alpha$ in policy learning.
-The policy shift coefficient $\beta$ regulates modifications to offline data, controlling deviation from the behavior policy.
-Larger $\beta$ values enhance exploration by encouraging greater deviation from the behavior policy, but also increase the risk of unsafe or suboptimal actions.
-As shown in Fig.
-5 (c), the algorithm exhibits performance fluctuations as $\beta$ increases from 0 to 0.1.
-Low $\beta$ may lead to suboptimal solutions, while high $\beta$ can induce excessive aggressiveness and instability.
-Therefore, $\beta$ also needs to be set appropriately.
-6 Conclusion and Discussion
-In this paper, we analyze the distribution shift problem, attributing its root causes to model bias and policy shift.
-Model bias results in distorted objectives, while policy shift obstructs the convergence of policies.
-We introduce a shifts-aware policy optimization framework to uniformly address these challenges through shift-aware rewards (SAR) and apply it within the model-based offline setting, thereby proposing our SAMBO-RL algorithm.
-Notably, our framework is versatile and can be integrated into general reinforcement learning algorithms.
+While this reduces value misestimation, the algorithm remains unstable due to the induced distribution shift between training and deployment.
+Consequently, wo/ps exhibits lower stability, and its final performance is slightly worse than that of SAMBO.
+In contrast, SAMBO incorporates both model bias adjustment and policy shift modification, facilitating stable updates and addressing value estimation challenges, thereby alleviating suboptimal performance.
+Furthermore, we performed parameter tuning experiments on $\alpha$ and $\beta$ to examine the robustness of our method with respect to parameter choices.
+As illustrated in Figure 5, we fix $\beta=0.01$ in example (a) and $\alpha=0.01$ in example (b).
+The results indicate that the performance of our algorithm remains generally stable and consistent within a certain range of parameter variations.
+(a) Tuning for $\alpha$
+(b) Tuning for $\beta$
+Figure 5. The tuning experiment was performed on halfchetah-random datasets in D4RL benchmark.
+6. Conclusion and Discussion
+In this paper, we analyze the distribution shift problem in model-based offline RL, attributing its root causes to model bias and policy shift.
+Model bias induces a distribution shift among training data originating from different sources, which distorts value estimation, while policy shift results in a distribution shift between training and deployment, hindering policy convergence.
+We introduce a shifts-aware policy optimization framework to uniformly address these challenges through shift-aware rewards (SAR).
+Theoretical and empirical analyses have demonstrated that SAR serves as an efficient surrogate for the true rewards.
+Building on SAR, we apply it within the model-based offline RL setting, proposing our SAMBO-RL algorithm.
+Empirical results demonstrate that SAMBO-RL consistently delivers superior or competitive performance, especially on the more challenging NeoRL benchmarks, underscoring SAMBO’s practical effectiveness and validating our theoretical analysis.
 Our main limitation is classifier accuracy; imprecise classifiers struggle to estimate the shifts-aware reward accurately, leading to unstable training.
-In future work, We aim to explore alternative methods that directly address distribution shifts using datasets, thereby eliminating the need for classifier estimation.
-Additionally, we will extend our shifts-aware reward framework to tackle distribution shift challenges in online settings, with a particular focus on policy shift problems.
-Our novel analysis and results related to policy shifts in this context open new avenues for better exploiting offline datasets and overcoming the implicit restrictions imposed by behavior policies.
-Impact Statement
-This study addresses the distribution shift problem in offline model-based reinforcement learning (RL), with the potential to facilitate the deployment of RL algorithms in real-world applications, particularly in scenarios where data collection is expensive.
-The research aims to enhance the robustness and reliability of RL systems in real world.
-However, the direct deployment of RL systems in high-risk domains, such as autonomous driving, may introduce substantial risks, especially if algorithms fail to adapt to unforeseen distribution shifts in real-world environments.
-Furthermore, the application of offline RL models, particularly in sensitive areas like healthcare, may raise significant ethical concerns.
-In such domains, the training data may not fully represent all patient groups.
-If distribution shift issues are not adequately addressed, the resulting treatment policies may fail for certain patient populations.
-Therefore, to ensure the safe, effective, and equitable deployment of RL systems in real world, rigorous testing, validation, and continuous monitoring are essential.
+In future work, We aim to explore alternative methods that directly address distribution shifts using datasets, thereby eliminating the need for explicit classifier estimation.
 Appendix A Theoretical Analysis
-A.1 The Supplement Analysis of Distribution Shift
-Dual Formulation of Model Bias.
-The reinforcement learning optimization objective can be equivalently reformulated in its dual form, expressed as $\mathcal{J}_{M}(\pi)=\mathbb{E}_{(s,a)\sim d^{p,\pi}}[r(s,a)]$, where $d^{p,\pi}(s,a)$ denotes the state-action visitation distribution under dynamics $p$ and policy $\pi$.
-To illustrate the impact of model bias in the dual formulation, we similarly examine the scenario where the training policy $\pi$ is applied within the model dynamics $q$.
-The training objective $\mathcal{J}(\pi)$ is derived as follows:
-$\mathcal{J}(\pi)=\mathbb{E}_{(s,a)\sim d^{q,\pi}}[r(s,a)]=\mathbb{E}_{(s,a)\sim d^{p,\pi}}[r(s,a)\frac{d^{q,\pi}(s,a)}{d^{p,\pi}(s,a)}].$
-Model bias can result in inaccuracies in estimating the state-action visitation distribution, subsequently affecting the accuracy of value function estimation.
-Similarly, if the learned model overestimates or underestimates the state-action visitation distribution, it will lead to corresponding overestimations or underestimations of rewards, thereby distorting the objective.
-Trajectory Formulation of Policy Shift.
-Similarly, we consider the trajectory formulation of policy shift.
-We execute the sampling policy $\pi_{c}$ within the environment transition dynamics $p$.
-Then the training objective $\mathcal{J}(\pi)$ is then given by:
-$\mathcal{J}(\pi)=\mathbb{E}_{\tau\sim p^{\pi_{c}}}[R(\tau)]=\mathbb{E}_{\tau\sim p^{\pi}}\left[R(\tau)\prod\limits_{t=0}^{\infty}\frac{\pi_{c}(a_{t}|s_{t})}{\pi(a_{t}|s_{t})}\right].$
-The policy shift can influence the value estimation of trajectories, imposing implicit constraints on policy learning.
-For a high-reward trajectory $\tau$, if the current policy $\pi$ closely approximates the optimal policy and satisfies $\prod_{t}\pi(a_{t}\mid s_{t})>\prod_{t}\pi_{c}(a_{t}\mid s_{t})$, it may lead to an underestimation of the return for trajectory $\tau$.
-Conversely, for a low-reward trajectory $\tau$, if the current policy is relatively effective and satisfies $\prod_{t}\pi(a_{t}\mid s_{t})<\prod_{t}\pi_{c}(a_{t}\mid s_{t})$, this can result in an overestimation of the return for trajectory $\tau$.
-Both of these impede the effective learning of the policy.
-A.2 The Proof of Theorem 1
-Theorem A.1.
+A.1. The Proof of Theorem 1
+Theorem. 1.
 Let $\mathcal{L}(\pi)$ be the training objective defined in (6) and $\tilde{r}$ denote the shifts-aware reward (5). Then,
@@
 Proof.
-The cumulative discounted reward $R(\tau)$ of trajectory $\tau=(s_{0},a_{0},r_{0},s_{1},...)$ can be written as:
+[…]
 $R(\tau)=\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})=\frac{1}{1-\gamma}\sum_{t=0}^{\infty}(1-\gamma)\gamma^{t}r(s_{t},a_{t}).$ (10)
 As discussed in the main text, the probability distribution over a trajectory $\tau$ can be formulated as follows:
-$p^{\pi}(\tau)=\mu_{0}(s_{0})\prod\limits_{t=0}^{\infty}p(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t})\text{ , }q^{\pi_{c}}(\tau)=\mu_{0}(s_{0})\prod\limits_{t=0}^{\infty}q(s_{t+1}|s_{t},a_{t})\pi_{c}(a_{t}|s_{t}).$ (10)
-Then according to (9) and (10) and Jensen’s inequality, we have
-$\log\left[\mathcal{J}_{M}(\pi)\right]$ $=\log\mathbb{E}_{p^{\pi}(\tau)}R(\tau)$
-$=\log\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[R(\tau)\frac{p^{\pi}(\tau)}{q^{\pi_{c}}(\tau)}\right]$
-$=\log\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[R(\tau)\frac{\mu(s_{0})\prod_{t=0}^{\infty}p(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t})}{\mu(s_{0})\prod_{t=0}^{\infty}q(s_{t+1}|s_{t},a_{t})\pi_{c}(a_{t}|s_{t})}\right]$
-$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\log R(\tau)+\sum_{t=0}^{\infty}\left(\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right)\right]$
-$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\log\frac{1}{1-\gamma}+\sum_{t=0}^{\infty}(1-\gamma)\gamma^{t}\log r(s_{t},a_{t})+\sum_{t=0}^{\infty}\left(\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right)\right]$
-$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\sum_{t=0}^{\infty}\left((1-\gamma)\gamma^{t}\log r(s_{t},a_{t})+\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right)\right]$
-$=(1-\gamma)\mathbb{E}_{q^{\pi_{c}}(\tau)}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(\log r(s_{t},a_{t})+\frac{1}{(1-\gamma)\gamma^{t}}\left(\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right)\right)\right]$
-$=(1-\gamma)\mathcal{L}(\pi).$
-A.3 Classifiers Estimation Derivations
+$p^{\pi}(\tau)=\mu_{0}(s_{0})\prod\limits_{t=0}^{\infty}p(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t})$ (11)
+$q^{\pi_{c}}(\tau)=\mu_{0}(s_{0})\prod\limits_{t=0}^{\infty}q(s_{t+1}|s_{t},a_{t})\pi_{c}(a_{t}|s_{t}).$ (12)
+Then according to (10), (11) and (12) and Jensen’s inequality, we have
+$\log[\mathcal{J}_{M}(\pi)]$
+$=\log\mathbb{E}_{p^{\pi}(\tau)}R(\tau)$
+$=\log\mathbb{E}_{q^{\pi_{c}}(\tau)}[R(\tau)\tfrac{p^{\pi}(\tau)}{q^{\pi_{c}}(\tau)}]$
+$=\log\mathbb{E}_{q^{\pi_{c}}(\tau)}\!\Big[R(\tau)\tfrac{\mu(s_{0})\prod_{t=0}^{\infty}p(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t})}{\mu(s_{0})\prod_{t=0}^{\infty}q(s_{t+1}|s_{t},a_{t})\pi_{c}(a_{t}|s_{t})}\Big]$
+$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\!\Big[\log R(\tau)+\sum_{t=0}^{\infty}(\log\tfrac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\tfrac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})})\Big]$
+$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\!\Big[\log\tfrac{1}{1-\gamma}+\sum_{t=0}^{\infty}(1-\gamma)\gamma^{t}\log r(s_{t},a_{t})$
+$\quad+\sum_{t=0}^{\infty}(\log\tfrac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\tfrac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})})\Big]$
+$\geq\mathbb{E}_{q^{\pi_{c}}(\tau)}\!\Big[\sum_{t=0}^{\infty}((1-\gamma)\gamma^{t}\log r(s_{t},a_{t})+\log\tfrac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\tfrac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})})\Big]$
+$=(1-\gamma)\mathbb{E}_{q^{\pi_{c}}(\tau)}\!\Big[\sum_{t=0}^{\infty}\gamma^{t}\Big(\log r(s_{t},a_{t})+\tfrac{1}{(1-\gamma)\gamma^{t}}(\log\tfrac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}$
+$\quad+\log\tfrac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})})\Big)\Big]$
+$=(1-\gamma)\mathcal{L}(\pi)$
+A.2. Classifiers Estimation Derivations
 We assume that all steps in the derivation are appropriate, that the denominator is non-zero, and that the argument of the logarithm is not zero.
-A.3.1 Transition Classifier Derivation.
-Define two events $\mathbb{E}$ and $\mathbb{M}$, where
+A.2.1. Transition Classifier Derivation.
+Define two events $\mathbb{E}$ and $\mathbb{M}$.
 […]
+$\text{the environmen }p\text{ to generate the subsequent state.}\}$
 […]
+$\text{the model }m\text{ to generate the subsequent state.}\}$
 The transition classifier is defined as $C_{\phi}(s_{t},a_{t},s_{t+1})=P(\mathbb{E}|s_{t},a_{t},s_{t+1})$. Then
@@
 $\frac{C_{\phi}(s,a,s_{t+1})}{1-C_{\phi}(s,a,s_{t+1})}=\frac{P(s_{t+1}|s_{t},a_{t},\mathbb{E})}{P(s_{t+1}|s_{t},a_{t},\mathbb{M})}\frac{P(\mathbb{E}|s_{t},a_{t})}{P(\mathbb{M}|s_{t},a_{t})}.$ (15)
-Specifically, $P(s_{t+1}|s_{t},a_{t},\mathbb{E})=p(s_{t+1}|s_{t},a_{t})$, where $p$ represents the environment transition dynamic; $P(s_{t+1}|s_{t},a_{t},\mathbb{M})=q(s_{t+1}|s_{t},a_{t})$, where $q$ denotes the model transition dynamic.
+Specifically, $P(s_{t+1}|s_{t},a_{t},\mathbb{E})=p(s_{t+1}|s_{t},a_{t})$, where $p$ represents the environment transition dynamic; $P(s_{t+1}|s_{t},a_{t},\mathbb{M})=m(s_{t+1}|s_{t},a_{t})$, where $m$ denotes the model transition dynamic.
 Then taking the logarithm of both sides of (15), we obtain:
-$\log\frac{C_{\phi}(s,a,s_{t+1})}{1-C_{\phi}(s,a,s_{t+1})}=\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{P(\mathbb{E}|s_{t},a_{t})}{P(\mathbb{M}|s_{t},a_{t})}.$ (14)
+$\log\frac{C_{\phi}(s,a,s_{t+1})}{1-C_{\phi}(s,a,s_{t+1})}=\log\frac{p(s_{t+1}|s_{t},a_{t})}{m(s_{t+1}|s_{t},a_{t})}+\log\frac{P(\mathbb{E}|s_{t},a_{t})}{P(\mathbb{M}|s_{t},a_{t})}.$ (16)
 Given the independence of the state-action pair $(s_{t},a_{t})$ from the transition dynamics, we assume that the visit distribution of $(s_{t},a_{t})$ remains identical. Then
@@
 Then according to (16) and (17), we have
-$\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}\approx\log\frac{C_{\phi}(s,a,s_{t+1})}{1-C_{\phi}(s,a,s_{t+1})}-c_{1},$
+$\log\frac{p(s_{t+1}|s_{t},a_{t})}{m(s_{t+1}|s_{t},a_{t})}\approx\log\frac{C_{\phi}(s,a,s_{t+1})}{1-C_{\phi}(s,a,s_{t+1})}-c_{1},$
 where $c_{1}$ is a constant that does not influence the training of SAMBO.
-A.3.2 Action Classifier Derivation.
+A.2.2. Action Classifier Derivation.
 The derivation of the action classifier is similar to that of the transition classifier. Define two events $\mathbb{A}$ and $\mathbb{C}$, where
 […]
-={Given the state $\mathbb{C}=\left\{\text{Given the state }s_{t},\text{ the agent takes action }a_{t}\text{ according to the behavior policy }\pi_{c}\right\}.$
+[…]
 The action classifier is defined as $C_{\psi_{1}}(s_{t},a_{t})=P(\mathbb{A}|s_{t},a_{t})$, Then
@@
 $\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}=\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$
+A.3. The Proof of Theorem 3
+Theorem. 3.
+Under the assumption that $D_{\operatorname{KL}}(m(\cdot\mid s,a)\,\|\,p(\cdot\mid s,a))\geq\epsilon_{KL}>0$, there exist $\alpha>0$ and $\beta>0$ such that the penalty of SAR forms a $\xi$-uncertainty quantifier.
+Proof.
+$\quad\left|\widehat{\mathbb{B}}^{\pi}\widehat{V}_{t+1}(s,a)-\mathbb{B}^{\pi}\widehat{V}_{t+1}(s,a)\right|$
+$=\gamma\left|\mathbb{E}_{s^{\prime}\sim m}[\widehat{V}_{t+1}(s^{\prime})]-\mathbb{E}_{s^{\prime}\sim p}[\widehat{V}_{t+1}(s^{\prime})]\right|$
+$\leq\frac{\gamma r_{\max}}{1-\gamma}D_{\mathrm{TV}}(m(\cdot|s,a),p(\cdot|s,a))$
+$\leq\frac{\gamma r_{\max}}{1-\gamma}\sqrt{\frac{1}{2}D_{KL}(m(\cdot|s,a)||p(\cdot|s,a))}$
+$\leq\frac{\gamma r_{\max}}{1-\gamma}\frac{1}{\sqrt{2\epsilon_{KL}}}D_{KL}(m(\cdot|s,a)||p(\cdot|s,a))$
+[…]
+For data in the model dataset $\mathcal{D}_{m}$, the dynamics are governed by the model $m$ and the data collection policy can be seen as $\pi$, i.e.
+$\pi_{c}=\pi$.
+Then the penalty of shifts-aware reward is $D_{KL}(m(\cdot|s,a)||p(\cdot|s,a))$.
+Thus, if $\alpha>0$ is properly set, the penalty of SAR forms a $\xi$-uncertainty quantifier.
+∎
+A.4. The Proof of Theorem 4
+Before proving Theorem 4, we introduce the pessimistic value iteration (PEVI) algorithm Jin et al. (2021), a meta-algorithm for offline RL.
+Algorithm 2 Pessimistic Value Iteration (PEVI): General MDP
+1:Dataset $\mathcal{D}=\{(x_{\tau,h}^{r},a_{\tau,h}^{r},r_{\tau,h}^{r})\}_{\tau,h}^{K,H}$.
+2:Initialize $\hat{V}_{H+1}(\cdot)\leftarrow 0$.
+3:for $h=H,H-1,\ldots,1$ do
+4:  Construct $(\hat{\mathbb{B}}_{h}\hat{V}_{h+1})(\cdot,\cdot)$ and $\Gamma_{h}(\cdot,\cdot)$ based on $\mathcal{D}$.
+5:  Set $\hat{Q}_{h}(\cdot,\cdot)\leftarrow(\hat{\mathbb{B}}_{h}\hat{V}_{h+1})(\cdot,\cdot)-\Gamma_{h}(\cdot,\cdot)$.
+6:  Set $\hat{Q}_{h}(\cdot,\cdot)\leftarrow\min\{\hat{Q}_{h}(\cdot,\cdot),H-h+1\}^{+}$.
+7:  Set $\hat{\pi}_{h}(\cdot)\leftarrow\arg\max_{\pi_{h}(\cdot)}\langle\hat{Q}_{h}(\cdot,\pi_{h}(\cdot)),\pi_{h}(\cdot)\rangle_{\mathcal{A}}$.
+8:  Set $\hat{V}_{h}(\cdot)\leftarrow\langle\hat{Q}_{h}(\cdot,\cdot),\hat{\pi}_{h}(\cdot)\rangle_{\mathcal{A}}$.
+9:end for
+10:Output: $\mathrm{Pess}(\mathcal{D})=\{\hat{\pi}_{h}\}_{h=1}^{H}$.
+The $\xi$-uncertainty quantifier characterizes the error between the empirical Bellman update and the true Bellman update.
+With the $\xi$-uncertainty quantifier, the following suboptimality theorem about Algorithm 2 holds.
+Definition 2.
+($\xi$-Uncertainty Quantifier Jin et al. (2021)). The set of penalization $\{\Gamma_{h}\}_{h\in[H]}$ forms a $\xi$-uncertainty quantifier if it holds with probability at least $1-\xi$ that
+$\left|\widehat{\mathbb{B}}\widehat{V}_{t+1}(s,a)-\mathbb{B}\widehat{V}_{t+1}(s,a)\right|\leq\Gamma_{t}(s,a),\quad\forall(s,a)\in\mathcal{S}\times\mathcal{A}$
+where $\widehat{V}_{t+1}$ is an estimated value function at step $t+1$ and $\widehat{\mathbb{B}}$ is an empirical Bellman operator estimating the Bellman operator $\mathbb{B}$.
+Theorem. A (Suboptimality for PEVI Jin et al. (2021)).
+Suppose $\{\Gamma_{h}\}_{h=1}^{H}$ in Algorithm 2 is a $\xi$-uncertainty quantifier. For any $s\in\mathcal{S}$, $\widehat{\pi}\in\text{Pess}(\mathcal{D})$ in Algorithm 2 satisfies
+$\left|V^{\pi^{*}}(s_{1})-V^{\widehat{\pi}}(s_{1})\right|\leq 2\sum_{h=1}^{H}\mathbb{E}_{\pi^{*}}\big[\Gamma_{h}(s_{h},a_{h})\mid s_{1}=s\big].$
+Here $\mathbb{E}_{\pi^{*}}$ is taken with respect to the trajectory induced by $\pi^{*}$ in the underlying MDP given the fixed function $\Gamma_{h}$.
+Theorem. 4.
+Let $\pi^{*}$ be the optimal policy in the environment, $\pi^{*}_{G}$ be the optimal policy in MDP $G=(\mathcal{S},\mathcal{A},p,\log r,\mu_{0},\gamma)$, $V_{G}^{\pi}(s)$ the value function in $G$ and $\epsilon_{s}\geq D_{\operatorname{KL}}(m(\cdot\mid s,a)\,\|\,p(\cdot\mid s,a))\geq\epsilon_{KL}>0$.
+Then the derived policy $\widehat{\pi}$ according to SAR satisfies the following inequality
+$\left|V^{\pi^{*}}(s_{1})-V^{\widehat{\pi}}(s_{1})\right|\leq 2\alpha H\epsilon_{s}+\epsilon(G;s_{1}),$
+with probability at least $1-\xi$ for all $s\in\mathcal{S}$. Here,
+$\epsilon(G;s)=|V^{\widehat{\pi}}(s)+V_{G}^{\pi^{*}_{G}}(s)-V^{\pi^{*}}(s)-V_{G}^{\widehat{\pi}}(s)|,$
+which quantify the value error between different MDPs and $\epsilon(M;s)=0,\forall s\in\mathcal{S}$.
+Proof.
+Decomposing the value bound,
+$V^{\pi^{*}}(s)-V^{\widehat{\pi}}(s)=V^{\pi^{*}}(s)-V^{\widehat{\pi}}_{G}(s)+V^{\widehat{\pi}}_{G}(s)-V^{\pi^{*}_{G}}_{G}(s)+V^{\pi^{*}_{G}}_{G}(s)-V^{\widehat{\pi}}(s).$
+As established in the proof of Theorem 3, the penalty term employed by SAR serves as a $\xi$-uncertainty quantifier. Thus, according to Theorem A, we have
+$\left|V^{\widehat{\pi}}_{G}(s)-V^{\pi^{*}_{G}}_{G}(s)\right|\leq 2\alpha\sum_{h=1}^{H}\mathbb{E}_{\pi^{*}_{G}}\left[D_{KL}(m(\cdot|s_{h},a_{h})||p(\cdot|s_{h},a_{h}))\mid s_{1}=s\right].$
+Then
+$\quad\left|V^{\pi^{*}}(s_{1})-V^{\widehat{\pi}}(s_{1})\right|$
+$\leq|V^{\widehat{\pi}}_{G}(s_{1})-V^{\pi^{*}_{G}}_{G}(s_{1})|+|V^{\pi^{*}}(s_{1})-V^{\widehat{\pi}}_{G}(s_{1})+V^{\pi^{*}_{G}}_{G}(s_{1})-V^{\widehat{\pi}}(s_{1})|$
+$\leq 2\alpha\sum_{h=1}^{H}\mathbb{E}_{\pi^{*}_{G}}\left[D_{KL}(m(\cdot|s_{h},a_{h})||p(\cdot|s_{h},a_{h}))\mid s_{1}=s\right]+\epsilon(G;s_{1})$
+$\leq 2\alpha H\epsilon_{s}+\epsilon(G;s_{1}).$
+[…]
 Appendix B Implementation Details
-We present a detailed outline of the SAMBO algorithm in Algorithm 1.
+We present a detailed outline of SAMBO algorithm in Algorithm 3.
+B.1. Model Training
+In our approach, the model is represented by a probabilistic neural network that outputs a Gaussian distribution for both the next state and reward, given the current state and action:
+$m_{\theta}\left(s_{t+1},r_{t}\mid s_{t},a_{t}\right)=\mathcal{N}\left(\mu_{\theta}\left(s_{t},a_{t}\right),\Sigma_{\theta}\left(s_{t},a_{t}\right)\right).$
+Our model training approach is consistent with the methodology used in prior works Janner et al.
+(2019); Yu et al.
+(2020).
+We train an ensemble of seven dynamics models and select the best five based on their validation prediction error from a held-out set containing 1000 transitions in the offline dataset $\mathcal{D}_{env}$.
+Each model in the ensemble is a 4-layer feedforward neural network with 200 hidden units.
+During model rollouts, we randomly choose one model from the best five models.
 Algorithm 3 SAMBO-RL
@@
 4:  Initialize the policy dataset buffer $\mathcal{D}_{\pi}\leftarrow\varnothing$.
-5:     for $1,2,...,b$ (in parallel) do
+5:  for 1,2,,b1,2,...,b (in parallel) do
 6:   Sample state $s_{1}$ from $\mathcal{D}_{env}$ for the initialization of the rollout.
-7:        for $j=1,2,...,h$ do
+7:   for j=1,2,,hj=1,2,...,h do
 8:     Sample an action $a_{j}\sim\pi(s_{j})$.
@@
 16:end for
-B.1 Model Training
-In our approach, the model is represented by a probabilistic neural network that outputs a Gaussian distribution for both the next state and reward, given the current state and action:
-$m_{\theta}\left(s_{t+1},r_{t}\mid s_{t},a_{t}\right)=\mathcal{N}\left(\mu_{\theta}\left(s_{t},a_{t}\right),\Sigma_{\theta}\left(s_{t},a_{t}\right)\right).$
-Our model training approach is consistent with the methodology used in prior works (Janner et al., 2019; Yu et al., 2020).
-We train an ensemble of seven dynamics models and select the best five based on their validation prediction error from a held-out set containing 1000 transitions in the offline dataset $\mathcal{D}_{env}$.
-Each model in the ensemble is a 4-layer feedforward neural network with 200 hidden units.
-During model rollouts, we randomly choose one model from the best five models.
-B.2 Classifiers Training
+B.2. Classifiers Training
 The transition classifier $C_{\phi}(s,a,s^{\prime})\in[0,1]$ represents the probability that the transition originated from the environment.
@@
 The action classifier $C_{\psi}(s,a)\in[0,1]$ represents the probability that a given state-action pair $(s,a)$ was generated by the current policy $\pi$. This classifier is trained by minimizing:
-[…]
-For data in the model dataset $\mathcal{D}_{m}$, the dynamics are governed by the model $q$.
-Furthermore, we approximate the training data collection policy for $\mathcal{D}_{m}$ using the current policy $\pi$, i.e.
+$\begin{aligned} \mathcal{L}_{\psi_{1}}&=-\mathbb{E}_{\mathcal{D}_{\pi}}\left[\log C_{\psi_{1}}(s,a)\right]-\mathbb{E}_{\mathcal{D}_{env}}\left[\log(1-C_{\psi_{1}}(s,a))\right],\\ \mathcal{L}_{\psi_{2}}&=-\mathbb{E}_{\mathcal{D}_{\pi}}\left[\log C_{\psi_{2}}(s)\right]-\mathbb{E}_{\mathcal{D}_{env}}\left[\log(1-C_{\psi_{2}}(s))\right].\end{aligned}$ (22)
+For data in the model dataset $\mathcal{D}_{m}$, the dynamics are governed by the model $m$ and the data collection policy can be seen as $\pi$, i.e.
 $\pi_{c}=\pi$.
-Therefore, $\forall(s,a,r,s^{\prime})\in\mathcal{D}_{m}$, the shifts-aware reward simplifies to
-$\tilde{r}(s,a,s^{\prime})=\log r(s,a)+\alpha\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}.$
+Therefore, according to (7), $\forall(s,a,r,s^{\prime})\in\mathcal{D}_{m}$, the shifts-aware reward simplifies to
+$\tilde{r}(s,a,s^{\prime})=\log r(s,a)+\alpha\log\frac{p(s^{\prime}|s,a)}{m(s^{\prime}|s,a)}.$
 Then, for all $(s,a,s^{\prime})\in\mathcal{D}_{m}$, the shift-aware reward can be estimated using the following formulation,
 $\tilde{r}(s,a,s^{\prime})\approx\log r(s,a)+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})}.$ (23)
-For data in the offline dataset $\mathcal{D}_{env}$, the dynamics are governed by environment $p$, i.e.
-$q=p$, and the training data collection policy aligns with the behavioral policy, i.e.
+For data in the offline dataset $\mathcal{D}_{env}$, the dynamics are governed by environment $p$, and the training data collection policy aligns with the behavioral policy, i.e.
 $\pi_{c}=\pi_{b}$.
@@
 $\tilde{r}(s,a,s^{\prime})\approx\log r(s,a)+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$ (24)
-B.3 Policy Learning
-The policy optimization in our approach is based on the SAC algorithm, with hyperparameter configurations aligning with those employed in MOBILE (Sun et al., 2023).
+B.3. Policy Learning
+The policy optimization in our approach is based on the SAC algorithm, with hyperparameter configurations aligning with those employed in MOBILE Sun et al.
+(2023).
 For each update, we draw a batch of 256 transitions, with $5\%$ sourced from the offline dataset $\mathcal{D}_{env}$ and the remaining $95\%$ from the synthetic dataset $\mathcal{D}_{m}$.
 The detailed hyperparameter settings used for the D4RL benchmark are provided in Table 3.
-Table 3: Hyperparameters for policy optimization in SAMBO
+Table 3. Hyperparameters for policy optimization in SAMBO
 Hyperparameters Value Description
@@
 Appendix C Experimental Details
-C.1 Benchmark
-We conduct experiments on Gym tasks (“v2” version) within the D4RL (Fu et al., 2020) benchmark.
-We also evaluated SAMBO on the NeoRL (Qin et al., 2022) benchmark, a more challenging dataset generated using a more conservative policy to better reflect real-world data collection scenarios.
+C.1. Benchmark and Baselines
+We conduct experiments on Gym tasks (“v2” version) within the D4RL Fu et al.
+(2020) benchmark.
+We also evaluated SAMBO on the NeoRL Qin et al.
+(2022) benchmark, a more challenging dataset generated using a more conservative policy to better reflect real-world data collection scenarios.
 These datasets are characterized by their narrow scope and limited coverage.
-D4RL.
+D4RL.We evaluate SAMBO against several baseline methods, including model-free methods such as BC Bain and Sammut (1995); Ross et al.
+(2011), which learns the behavior policy, CQL Kumar et al.
+(2020), TD3+BC Fujimoto and Gu (2021), and EDAC An et al.
+(2021), as well as model-based methods such as MOPO Yu et al.
+(2020), COMBO Yu et al.
+(2021), TT Janner et al.
+(2021), RAMBO Rigter et al.
+(2022), and MOBILE Sun et al.
+(2023).
 We now provide the sources for the reported performance on this benchmark.
-For TD3+BC (Fujimoto & Gu, 2021), EDAC (An et al., 2021), TT (Janner et al., 2021), RAMBO (Rigter et al., 2022), and MOBILE (Sun et al., 2023), we cite the scores reported in the original papers, as these evaluations were conducted on the “v2” datasets in Gym.
-Since COMBO (Yu et al., 2021) does not provide source code, we include the results as reported in the original paper.
-The results of CQL (Kumar et al., 2020) are taken from the performance table in MOBILE, where CQL was retrained on the “v2” datasets in Gym.
+For TD3+BC Fujimoto and Gu (2021), EDAC An et al.
+(2021), TT Janner et al.
+(2021), RAMBO Rigter et al.
+(2022), and MOBILE Sun et al.
+(2023), we cite the scores reported in the original papers, as these evaluations were conducted on the “v2” datasets in Gym.
+Since COMBO Yu et al.
+(2021) does not provide source code, we include the results as reported in the original paper.
+The results of CQL Kumar et al.
+(2020) are taken from the performance table in MOBILE, where CQL was retrained on the “v2” datasets in Gym.
 The results of MOPO* are obtained from experiments using the OfflineRL-Kit library [^2] https://github.com/yihaosun1124/OfflineRL-Kit on the “v2” datasets.
 These scores are referenced in Table 1.
-NeoRL.
-We report the performance of BC, CQL, and MOPO based on the original NeoRL paper, while the performance of TD3+BC, EDAC, and MOBILE is derived from the scores reported in the MOBILE (Sun et al., 2023) paper.
-C.2 Hyperparameters
+NeoRL.We compare SAMBO with baseline methods that align with the d4rl experiments, excluding TT, COMBO, and RAMBO, as no results for these methods are reported in their original papers or the NeoRL paper, and determining suitable hyperparameters for them would require an excessive amount of computational time.
+We report the performance of BC, CQL, and MOPO based on the original NeoRL paper, while the performance of TD3+BC, EDAC, and MOBILE is derived from the scores in the MOBILE paper.
+C.2. Hyperparameters
 We list the hyperparameters that have been tuned as follows. The specific settings are detailed in Table 4 and Table 5.
-Table 4: Hyperparameters of SAMBO used in the D4RL datasets.
-Task Name $\boldsymbol{\alpha}$ $\boldsymbol{\beta}$ $\boldsymbol{h}$ $\boldsymbol{c}$
-halfcheetah-random 0.01 0.01 5 -0.2
-hopper-random 0.02 0.01 5 0
-walker2d-random 0.3 0.3 5 0
-halfcheetah-medium 0.03 0.03 5 0
-hopper-medium 0.02 0.01 5 0
-walker2d-medium 0.035 0.01 5 0
-halfcheetah-medium-replay 0.01 0.01 5 0
-hopper-medium-replay 0.05 0.05 5 0
-walker2d-medium-replay 1 1 1 0
-halfcheetah-medium-expert 0.3 0.3 5 0
-hopper-medium-expert 1 1 5 0
-walker2d-medium-expert 2 2 1 0
-Table 5: Hyperparameters of SAMBO used in the NeoRL datasets.
-Task Name $\boldsymbol{\alpha}$ $\boldsymbol{\beta}$ $\boldsymbol{h}$ $\boldsymbol{c}$
-Halfcheetah-L 0.01 0.01 5 0
-Hopper-L 0.3 0.3 5 0
-Walker2d-L 3 1 1 0
-Halfcheetah-M 0.1 0.1 5 0
-Hopper-M 0.3 0.3 5 0
-Walker2d-M 3 1 1 0
-Halfcheetah-H 5 1 5 0
-Hopper-H 0.3 0.3 5 0
-Walker2d-H 3 1 1 0
+Table 4. Hyperparameters of SAMBO used in the D4RL.
+Task Name $\boldsymbol{\alpha}$ $\boldsymbol{\beta}$ $\boldsymbol{h}$
+halfcheetah-random 0.01 0.01 5
+hopper-random 0.02 0.01 5
+walker2d-random 0.3 0.3 5
+halfcheetah-medium 0.03 0.03 5
+hopper-medium 0.02 0.01 5
+walker2d-medium 0.03 0.01 5
+halfcheetah-medium-replay 0.01 0.01 5
+hopper-medium-replay 0.05 0.05 5
+walker2d-medium-replay 1 1 1
+halfcheetah-medium-expert 0.3 0.3 5
+hopper-medium-expert 1 1 5
+walker2d-medium-expert 2 2 1
+Table 5. Hyperparameters of SAMBO used in the NeoRL.
+Task Name $\boldsymbol{\alpha}$ $\boldsymbol{\beta}$ $\boldsymbol{h}$
+Halfcheetah-L 0.1 0.1 5
+Hopper-L 0.3 0.3 5
+Walker2d-L 3 1 1
+Halfcheetah-M 0.1 0.1 5
+Hopper-M 0.3 0.3 5
+Walker2d-M 3 1 1
+Halfcheetah-H 5 1 5
+Hopper-H 0.3 0.3 5
+Walker2d-H 3 1 1
 Coefficient $\alpha$ and $\beta$.
@@
 Rollout Length $h$.
-Similar to MOPO, We perform short-horizon branch rollouts in SAMBO. we tune $h$ in the range of $\{1,5\}$ for Gym tasks in the D4RL benchmark.
-Reward Truncation Coefficient $c$.
-To compute the logarithm of the reward term in the shifts-aware reward, we apply a translation to the real reward, which does not affect policy optimization.
-Specifically, given the offline dataset $\mathcal{D}_{env}$, let $r_{max}$ denote the maximum reward and $r_{min}$ denote the minimum reward in this dataset.
-Then we perform a translation and truncation on the real reward as follows:
-$r^{\prime}=\min\{1\times 10^{-8},r-c(r_{max}-r_{min})+1\times 10^{-8}\}.$
-Then we calculate the shifts-aware reward by
-$\tilde{r}(s,a,s^{\prime})\approx\log r^{\prime}(s,a)+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})}+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$
-In most datasets, we set $c=0$. However, we observe that employing a negative value for $c$ yields favorable results in certain datasets.
+Similar to MOPO, We perform short-horizon branch rollouts in SAMBO.
+we tune $h$ in the range of $\{1,5\}$ for Gym tasks in the D4RL benchmark.
 We terminate training at 1M to prevent potential performance degradation due to the rapid convergence of the HalfCheetah environment in NeoRL.
@@
 Similarly, in the walker2d-medium dataset of the D4RL benchmark, we ensure training stability by fixing the transition classifier after 1 million iterations.
-C.3 Reward Function Discussion
+C.3. Reward Function Discussion
 Based on our derivation, if we start with $\mathbb{E}_{p^{\pi}(\tau)}e^{R(\tau)}$, we can obtain a new surrogate objective, which we refer to as SAMBO-r.
@@
     $\tilde{r}(s,a,s^{\prime})=r^{\prime}(s,a)+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})}+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$
-C.4 Computing Infrastructure
+    In this case, $\epsilon(G;s)=0,\forall s\in\mathcal{S}$ in Theorem 4, but this derived surrogate objective is not a lower bound of the original RL objective.
+C.4. Computing Infrastructure
 All experiments are conducted using a single GeForce GTX 3090 GPU and an Intel(R) Xeon(R) Gold 6330 CPU @ 2.00GHz.
@@
 Appendix D Omitted Experiments
-D.1 Visualization
-Figure 6: The $(s,a,s^{\prime})$ distributions generated by SAMBO result in the walker2d-medium-expert task in D4RL benchmark.
+D.1. Visualization
+Figure 6.
+The $(s,a,s^{\prime})$ distributions generated by SAMBO in the walker2d-medium-expert task in D4RL benchmark.
 Model Data represents the distribution of data generated by executing the final learned policy in the dynamic model, while Real Data represents the distribution of data generated by executing the final learned policy in the environment.
@@
 This highlights the effectiveness of SAMBO in mitigating distribution shift across both benchmarks.
-D.2 Detailed Ablation Discussion
-SAMBO is regulated by the model bias adjustment coefficient $\alpha$ and the policy shift modification coefficient $\beta$.
-The model bias adjustment coefficient $\alpha$ controls the degree of adjustment to model-generated data.
-A larger $\alpha$ leads to excessive conservatism, restricting exploration even in well-performing regions and potentially causing suboptimal performance.
-Conversely, a smaller $\alpha$ may encourage excessive optimism in exploring risky regions, impairing performance.
-To validate this, we set $\beta=0.01$ and varied $\alpha$ in the Halfcheetah task.
-The results, shown in Fig.
-7 (a), indicate that both very low ($\alpha=0$) and high ($\alpha=0.02$ or $0.05$) values lead to decreased performance.
-Specifically, as $\alpha$ increases from 0 to 0.05, the algorithm initially improves but later declines as it becomes increasingly conservative, which impedes exploration.
-This underscores that both excessively low and high values of $\alpha$ are detrimental to policy learning.
-The policy shift modification coefficient $\beta$ governs the extent of modification applied to the offline data, thereby influencing the degree of deviation from the behavior policy.
-A larger $\beta$ value promotes increased exploratory behavior by encouraging the algorithm to diverge from the behavior policy.
-However, this also introduces a heightened risk of taking unsafe actions.
-As shown in Fig.
-7 (b), the algorithm exhibits some performance fluctuations as $\beta$ increases from 0 to 0.1, though it generally maintains robustness.
-A low value of $\beta$ may lead to suboptimal solutions, whereas a high value of $\beta$ can induce excessive aggressiveness of the algorithm, resulting in performance instability.
-Thus, $\beta$ should also be set appropriately to balance exploration and stability.
-(a) Tuning for $\alpha$
-(b) Tuning for $\beta$
-Figure 7: Illustration of hyperparameter experiments in the HalfCheetah-random task in D4RL benchmark.
-D.3 Performance Compare
+Figure 7. Hopper-L
+Figure 8. Hopper-M
+Figure 9. Hopper-H
+Figure 10. Walker2d-L
+Figure 11. Walker2d-M
+Figure 12. Walker2d-H
+Figure 13. HalfCheetah-L
+Figure 14. HalfCheetah-M
+Figure 15. HalfCheetah-H
+D.2. Performance Comparison
 We also compared the performance curves of model-based approaches on other tasks in the NeoRL benchmark.
@@
 Since MOPO does not have an official implementation on NeoRL, we report its score based on the results presented in the original NeoRL paper.
-As shown in Figures 10 to 16, SAMBO outperforms in all Hopper and Walker2d tasks, both of which are more complex and challenging.
-Although it exhibits slower convergence initially in the Walker2d-M and Walker2d-H tasks, its performance improves rapidly as the experiment progresses, ultimately achieving state-of-the-art (SOTA) levels.
-On the other hand, SAMBO also demonstrates comparable performance in simpler HalfCheetah tasks.
-Figure 8: Hopper-L
-Figure 9: Hopper-M
-Figure 10: Hopper-H
-Figure 11: Walker2d-L
-Figure 12: Walker2d-M
-Figure 13: Walker2d-H
-Figure 14: HalfCheetah-L
-Figure 15: HalfCheetah-M
-Figure 16: HalfCheetah-H
+As shown in Figures 15 to 15, SAMBO outperforms in all Hopper and Walker2d tasks, both of which are more complex and challenging.
+Although it exhibits slower convergence initially in the Walker2d-M and Walker2d-H tasks, its performance improves rapidly as the experiment progresses, ultimately achieving SOTA levels.
+On the other hand, SAMBO does not achieve strong performance on the HalfCheetah tasks.
+We attribute this to the fact that the HalfCheetah score mainly depends on achieving high running speed rather than mere stability.
+Since SAMBO’s dual conservatism with respect to model bias and policy shift places greater emphasis on stability, it performs less competitively on HalfCheetah, while demonstrating strong results on Hopper and Walker2d, where stability is more critical.